### v1

# GlucOS: Security, correctness, and simplicity for automated insulin delivery

Abstract
Type 1 Diabetes (T1D) is a metabolic disorder where an individual’s pancreas stops producing insulin. To compensate, they inject synthetic insulin. Computer systems, called automated insulin delivery systems, exist that inject insulin automatically. However, insulin is a dangerous hormone, where too much insulin can kill people in a matter of hours and too little insulin can kill people in a matter of days.
In this paper, we take on the challenge of building a new trustworthy automated insulin delivery system, called GlucOS. In our design, we apply separation principles to keep our implementation simple, we use formal methods to prove correct the most critical parts of the system, and we design novel security mechanisms and policies to withstand malicious components and attacks on the system.
We report on real world use for one individual for 6 months using GlucOS. Our data shows that for this individual, our ML-based algorithm runs safely and manages their T1D effectively. We also run our system on 21 virtual humans using simulations and show that our security and safety mechanisms enable ML to improve their core T1D measures of metabolic health by 4.3% on average. Finally, we show that our security and safety mechanisms maintain recommended levels of control over T1D even in the face of active attacks that would have otherwise led to death.
GlucOS is open source and our code is available on GitHub.
1 Introduction
Type 1 diabetes (T1D) is an autoimmune disease in which an individual’s pancreas stops producing insulin. People living with T1D must therefore take synthetic insulin and play the role of their pancreas. However, replicating the pancreas’s native insulin release patterns is extremely difficult, and optimal T1D self-management often requires hourly adjustments in insulin dosing both in response to physiological changes and in prediction of upcoming carbohydrate intake and energy expenditure (e.g., exercise, stress) [51, 56]. Even when following this labor-intensive process of T1D self-management, affected individuals remain at risk for acute episodes of hypoglycemia (low blood sugar, which can cause loss of consciousness, seizure, or death) or ketoacidosis (which carries an associated risk for multi-organ injury), as well as for long term vascular complications (e.g., kidney damage, nerve damage, blindness, heart attack, stroke) that are a result of chronic hyperglycemia (high blood sugar) [46]. 8.4 million people are living with T1D worldwide [25].
(a) Devices for automated insulin delivery. This figure is from jdrf.org.
(b) Closed-loop architecture.
(c) Our system for automated insulin delivery.
Figure 1: Overview of automated insulin delivery devices, closed-loop architecture, and apps.
The last two decades have witnessed the development of therapeutic technology to assist T1D management, including continuous glucose monitoring (CGM) devices and insulin pumps. Most recently, the integration of CGM sensors and insulin pumps into “artificial pancreas” or automated insulin delivery systems has enabled a simultaneous improvement in health outcomes and reduction in the burden of care for T1D-affected individuals [38].
However, existing automated insulin delivery systems are complex, and there is no way to assure beyond any doubt that they are implemented correctly. For example, techniques from formal methods have not been applied to verify automated insulin delivery systems. With little reason to trust the systems and their core algorithms, users and developers are also reluctant to incorporate experimental algorithms into the prediction loop, such as advanced ML-driven insulin delivery. Researchers have shown that ML can improve automated insulin delivery systems [64, 33, 63, 59, 65, 58], yet the lack of security and safety for these ML predictions keeps people from using it. If ML is going to be part of medical decisions, we need the security community to lead the way on how to do this safely.
In this paper, we argue for applying traditional security and operating systems design principles to make automated insulin delivery systems more trustworthy. This includes going beyond just patient privacy or medical device security, to systems security for automated insulin delivery systems to ensure that they are implemented correctly and that the incorporation of ML does not jeopardize their safety.
Automated insulin delivery systems are different from traditional applications that the security community focuses on, but they are important. Research on operating systems or web browsers can list scores of CVEs from current systems and show how their system improves security by eliminating the impact of some or all of these vulnerabilities by virtue of their design. Automated insulin delivery systems do not have a long list of CVEs that we can reference. However, our position is that we need to make sure that these systems never get CVEs, because if they do, security vulnerabilities can kill people. Thus, we focus on building a new system, using security first principles and formal methods, to avoid security vulnerabilities and mitigate their damage from the start.
Designing and implementing trustworthy automated insulin delivery systems brings with it three main challenges. (1) How can automated insulin delivery design decisions improve rather than worsen the security posture of such software for end users? (2) How do we apply formal methods to a complex, distributed system that includes a human? (3) How can we cope with ML mispredictions, including those from malicious ML?
In this paper, we take on these challenges and design and implement a new trustworthy automated insulin delivery system, called GlucOS. Our design for GlucOS strives to achieve the following four primary goals: to use sound security principles to separate security and safety concerns into critical and non-critical components; to keep our implementation as simple as possible; to prove correct the most critical components; and to enable any predictive algorithms, including deep neural network ML, to control the insulin pump securely and safely.
GlucOS improves on the current state-of-the-art in automated insulin delivery systems by, for the first time, enabling deep neural network ML to control the insulin pump securely and safely. Our key insight for security is that over a long enough period of time, all correct algorithms will dose the same amount of insulin. For example, on a given day, a person will need to dose a fixed amount of insulin to absorb all of the glucose from the food they eat. Since synthetic insulin action is slow, injecting insulin early by predicting future metabolic states is critical for long-term health. GlucOS’s support for modern ML-based algorithms provides users with this predictive power to dose insulin. For security, GlucOS pairs a predictive ML model with a conservative and safe model to provide the ML with enough flexibility to control the insulin pump proactively while staying within the bounds dictated by the safe model.
Our architecture and system are a stark contrast to current systems. Current systems are complex, their architecture provides little or no isolation boundaries, they are unable to withstand attacks on their system, and they are unable to use deep neural network ML securely and safely. Given that mistakes with insulin delivery can kill people, our goal is to have a fully-featured automated insulin delivery system with security and safety as the primary design considerations. Even if parts of the system are malicious, GlucOS continues to operate and ensure that the individual using the system can manage their T1D safely.
We report our experiences running GlucOS on one individual for six months to manage their T1D. Running our software on a real human forces us to design a practical and real system. Our evaluation using virtual humans in a simulator shows that our security and safety mechanisms generalize beyond the specific individual.
Our novel contributions include:
  • The clean-slate design and implementation of a simple yet effective automated insulin delivery system.
  • Security mechanisms and policies that protect individuals from malicious ML, vulnerable pump drivers, and drastic changes in the human’s glycemic response.
  • The application of formal methods to prove correct the most critical parts of GlucOS.
  • A case study describing our experiences from a real-world deployment, and results from simulation to show that our techniques generalize.
We give top priority to the health, safety, and ethics of the human using GlucOS to manage their T1D. Section 11.1 summarizes our safety protocol and Appendix A provides a detailed description.
2 Background
Type 1 Diabetes. T1D is an autoimmune condition in which the immune system attacks and destroys the insulin-making cells in the pancreas. Insulin is a hormone that enables the absorption of glucose (sugar) from the bloodstream into the brain, the muscles, and other organs and tissues where it is used as an energy source. When people eat, the digestive system converts food into glucose, which makes its way into the bloodstream, but it cannot then be used by the body without the action of insulin.
Since people with T1D are unable to produce their own insulin, they must inject synthetic insulin to replace this key function, and the appropriate dosing of this insulin is both complicated and labor-intensive. For example, insulin doses must be closely correlated with food intake, but the absorption of exogenous insulin is slower than food digestion, so people with T1D must precisely predict their future food intake in order to inject the right amount of insulin prior to each meal. Once injected, typical insulin formulations will affect glucose metabolism for up to three hours, and during that time will interact with other factors like exercise, sleep, stress, caffeine, and alcohol. However, the exact effects of these factors on an individual’s glucose levels are even harder to predict than those of food alone, resulting in the need for continuous adjustments by the individual in order to maintain a balanced glycemic state. One solution that has emerged to address the complications of daily T1D management is automated insulin delivery systems, which connect a CGM sensor and insulin pump to enable automatic, individualized adjustments in insulin delivery based on glucose trends [14, 18, 40, 48, 17, 7, 4].
The goal of T1D management is to maintain glucose levels that are in a safe, target range (70-180 mg/dl) for the majority of the time – an outcome that has been linked to positive long-term health outcomes. Glucose levels 54-69 mg/dl (level 1 hypoglycemia) or less than 54 mg/dl (level 2 hypoglycemia) put individuals at risk for acute complications like impaired consciousness and seizure, while levels 181-250 mg/dl (level 1 hyperglycemia) or greater than 250 mg/dl (level 2 hyperglycemia) increase risks for long-term vascular damage. The main physiological concept in this paper is a goal of maintaining glucose levels that are in the target range the majority of the time.
Automated insulin delivery systems. Automated insulin delivery systems connect a CGM device – which measures interstitial glucose concentration every 1-5 minutes – with an insulin pump, which delivers a continuous subcutaneous insulin infusion. Using a variety of algorithms, these systems vary insulin delivery in response to CGM data, reducing the amount of human input and cognitive load required to achieve glycemic stability (Figure 1). Several commercial [17, 7, 4] and open source [40, 48] automated insulin delivery systems exist today.
3 Overview
This paper describes our design for GlucOS, a system for trustworthy automated insulin delivery. In this section, we first provide an overview of GlucOS’s system architecture to describe the key components involved in automated insulin delivery. While we provide an overview of all components, we specifically focus on those components that are part of our trusted computing base. We then define our threat model in terms of our assumptions and goals with GlucOS.
3.1 Architecture
From a system design perspective, our primary goals are to keep our implementation simple and enable our use of formal methods to prove correct the most critical parts. For our software system architecture, we use separation principles from the OS and microkernel areas [10, 30] applied to the application layer to push complexity away from the most critical parts of the system. This architecture is similar to secure web browsers [26, 61, 57, 53], which also apply OS separation principles at the application layer. By decomposing the overall automated insulin delivery system, we strive to keep our trusted computing base simple. Also, we want flexibility so that people can run whatever closed-loop algorithms they want to and use whatever pump hardware they want – we focus on providing mechanisms, policies, and systems for security and safety.
Figure 2: Overall architecture for GlucOS.
Figure 2 shows our overall system architecture. At the core of our overall architecture is the BioKernel. The BioKernel is the component that interacts with the CGM and insulin pump hardware, executes the closed-loop dosing algorithm, runs security and safety checks, and produces event logs that other components use to learn the state of the system.
The BioKernel runs the closed-loop algorithm every five minutes as new CGM readings come in. Once the BioKernel receives new CGM readings, it collects the current metabolic state of the individual and sends it to a predictive ML model. The predictive ML model calculates the amount of insulin to inject based on its assessment of the individual’s current metabolic state and predictions of future metabolic states. The BioKernel then issues commands to the pump to deliver the calculated amount of insulin. A pump driver converts these commands into low-level I/O to program the pump and adjust insulin delivery.
To withstand attacks, the BioKernel serves as a reference monitor [12] to enforce security policies on the predictive ML model’s pump commands and the pump driver’s low-level I/O. This reference monitor architecture ensures that we can withstand attacks from malicious predictive ML as well as malicious pump drivers without having to know the internals of how they operate.
For predictive ML, our reference monitor enforces algorithmic security, where we bound the amount of insulin calculated by ML to stay within dynamic safety bounds computed by a reactive safe physiological model. For the pump driver, we interpose on the I/O path to ensure that only the pump commands issued by the BioKernel are programmed into the pump. These checks constitute our pump invariants, which, when violated switch the pump to operate in a safe mode.
Our use of formal methods focuses on the software related to delivering insulin, which is the most risky operation in any automated insulin delivery system. We use Hoare logic to prove key functions correct, define system states and transitions to handle runtime verification check failures safely, and we introduce the notion of a novel biological invariant. Our biological invariant is an end-to-end check of our theoretical calculations of how the individual’s metabolism should behave compared to what we observe in practice, where violations of this invariant suggest a potentially catastrophic failure of the system.
3.2 Threat Model
We consider attacks from malicious predictive ML algorithms as well as malicious pump drivers. We assume that malicious ML can intentionally miscalculate the amount of insulin to inject and malicious pump drivers can issue arbitrary pump I/O operations to add, drop, or modify commands sent to the pump.
We also consider real-world safety concerns as part of our threat model. ML mispredictions are part of these safety concerns since even well-designed ML will sometimes make poor decisions. Safety concerns also include pump communication errors emerging from bugs in the pump driver, pump mechanical failures [55], and pump site skin inflammation or infection [36], all of which result in a mismatch between the amount of insulin calculated in closed loop and the actual amount of insulin active in the human.
Our goal with GlucOS is to not merely detect attacks and safety violations. Our goal is to build a secure system for automated insulin delivery that can withstand active attacks and real-world safety concerns.
We assume that CGM sensors are correct and operate within their error bounds. If the CGM sensors stop producing measurements, our system will be unable to handle it. We also assume that the Swift type system is correct and that the underlying operating system, the CGM driver and hardware, and the pump hardware we use are free of attacks and their device communications maintain integrity. Thus, we do not consider attacks emerging from tampered CGM sensors or pump hardware (such as adversarial attacks against predictive ML) in our threat model. These are important but complementary problems to our work, which focuses on the system architecture and algorithms for automated insulin delivery.
4 Algorithmic security: Withstanding malicious ML
None of the current automated insulin delivery systems use deep neural network ML. Although they do use ML in general, they use statistical methods, like linear regression or model predictive control [19, 60, 42, 52, 48], to learn physiological parameters. The reason for using this more traditional form of ML is sound: it provides determinism and has a physiological basis that people can use to vet its decisions. However, all of the recent advances in ML, like image classifiers and large language models, have come from deep neural networks, which researchers have shown also work well for predicting metabolic states for use in automated insulin delivery systems [64, 33, 63, 59, 65, 58]. Despite this promise, none of the current automated insulin delivery systems use more advanced deep neural networks due to the black-box nature of deep neural networks and the potentially dire consequences of mispredictions – an automated insulin delivery system’s unchecked “hallucination” [1] could be lethal if it delivers an inappropriate insulin dose [20, 13].
Modern ML models make mistakes. High-profile examples include Google Photos classifying black people as gorillas [2], LinkedIn recruiting ML learning gender bias [5], Microsoft’s chatbot, Tay, trained on Twitter data (what could go wrong) becoming racist [3], and ChatGPT large language model hallucinations [1]. In all of these examples, professional ML engineers with plenty of resources produced well-designed ML systems that hit a gray area in their training data or decision space and made bad decisions.
All automated insulin delivery systems that use ML need to have a mechanism to protect against malicious ML or benign ML that makes inevitable bad decisions. In this section, we discuss the design and implementation of GlucOS’s algorithmic security and safety mechanisms.
4.1 Mechanism fundamentals
Our key insight that underpins our security mechanism design is that all correct insulin delivery algorithms will deliver the same amount of insulin over a long enough period of time. However, the timing of this insulin delivery matters for maintaining balanced glucose levels (Section 10). Thus, our algorithmic security mechanism pairs a predictive model that anticipates future metabolic states and doses insulin accordingly with a reactive safe model that measures the current metabolic state and doses insulin based only on what it can measure currently.
Conceptually, the predictive model is the primary model in our system and should control the pump most of the time, and we use the reactive safe model for accounting so that we can track how far the predictive model deviates from a known-to-be-safe baseline. Then, we bound the size of this deviation. By combining these two models, we strive to get both the safety of a reactive safe model that operates only on facts (Section 4.2) and the tight control we expect from an ML-based predictive model (Section 4.3).
Figure 3: Algorithmic security mechanism for GlucOS.
Figure 3 shows how our algorithmic security mechanism works. The whole process starts when a new CGM reading arrives. Once the BioKernel receives a new CGM reading, it collects a recent history of CGM readings and insulin dosing and sends these as inputs to both the reactive safe model and the predictive ML model. These models independently produce pump commands, called a temporary basal command. The temporary basal command sets the insulin delivery rate for the insulin pump for a 30-minute duration. The BioKernel will collect both temporary basal commands and look back at the differences between the reactive safe model and predictive ML model for the past three hours to see how much insulin delivery it should attribute to the predictive ML model. If the predictive ML model is still within its insulin delivery limits, it uses the predictive ML model’s temporary basal command to program the pump. If it has exhausted its limits, we clamp it to fit within the insulin bounds we set relative to the reactive safe model. For more details about how this algorithm works, please see our implementation in GitHub (link omitted).
4.2 Reactive safe model
Since we use our reactive safe model to anchor our security policy, and it runs within our trusted computing base, our goals for this model are to have it be safe, easy to understand, and have a simple implementation.
To make this model safe, we base its calculations solely on facts observed automatically from the CGM and the insulin pump. Given these high-quality sources of data, and a simple insulin absorption physiological model, we can calculate how much insulin an individual needs and how much insulin they have injected but have not yet activated. From this calculation, we know how much insulin to inject or withhold.
These calculations are standard calculations that people who inject insulin manually use to determine dosing, making them easy for people who manage T1D or their medical care team to understand. The difference in our system is that we run these calculations automatically and every five minutes to adjust to the latest sensor readings. This basic formulation is a standard formulation used in most automated insulin delivery systems.
One wrinkle we add to this otherwise standard calculation is that we learn the physiological parameters we use in these calculations using a data-driven approach for safety. Our learning system consists of a weekly learning task that evaluates three weeks worth of data to infer broader physiological trends. These learned settings become the baseline for our parameters. Then, we use a PID controller from feedback control [22] to refine these values given recent observations. Being able to adapt is important for handling pump malfunction and pump insertion site safety violations (Section 3.2).
Calculating insulin is a simple calculation to implement, consisting of only 25 lines of formally verified code in the BioKernel (Section 8). However, despite this simplicity, it is effective (Sections 10 and 11). The main drawback of this model is that is it unable to detect underdosing insulin during food digestion quickly enough to avoid high glucose.
4.3 Predictive model
To predict insulin underdosing during digestion, we use a food prediction deep-learning model from the literature [45]. Our version of this model is a generalization where we model any source of glucose, not just food, but the details of it are beyond the scope of this paper. In our evaluation our model outperforms the reactive safe model both in simulation (Section 10) and for the individual that uses our system for their real-world insulin dosing (Section 11), showing why it is important for automated insulin delivery systems to have a practical way to adopt ML.
Figure 1 shows an example of when the predictive ML model correctly detects digestion and doses insulin to maintain balance. This figure shows a screenshot from the BioKernel’s main user interface. The individual ate breakfast at 9 am and their glucose levels rose. The predictive ML model correctly detects that the individual is going to absorb additional glucose from digestion, resulting in a predicted glucose level rise over the next hour. Accordingly, the predictive ML model injects enough insulin to cover the rising glucose from the meal and keep the individual’s glucose levels in range.
5 Driver security: Withstanding malicious pump drivers
Pump drivers provide abstractions for insulin delivery applications to communicate with insulin pumps. Much like traditional device drivers, they have complex implementations since they help accomplish a wide range of tasks including pump setup/teardown, managing pump UX, handling insulin delivery/suspension, managing pump errors, accurately tracking insulin doses etc. Different pump drivers use different implementations to carry out these tasks.
While CVEs do not currently exist for pump drivers, there is strong possibility for bugs or vulnerabilities in their implementation given their complexity and the fact that they not been formally verified. Since drivers control insulin pumps, these vulnerabilities can even lead to the death of individuals. For example, consider a scenario where the user decides to inject 3 U of insulin using the BioKernel. The BioKernel would relay this as a command to the pump through abstractions provided by the pump driver. If the pump driver inadvertently or intentionally issues this command twice, 6 U of insulin would be delivered to the user, which would put the user in danger by drastically reducing the amount of glucose in their body.
With GlucOS, our goal is to support their complexity by allowing users to employ pump drivers of their choice while upholding security. Our security goals are to not just protect users from inadvertent bugs in pump drivers, but to also protect them from malicious pump drivers which could intentionally add, modify or drop pump commands.
Figure 4: We interpose on communications between the pump driver and the pump to handle untrusted pump drivers to deliver insulin.
5.1 Reference monitor overview
We introduce a reference monitor as part of the BioKernel to handle untrusted pump drivers, which interposes on communications between the pump driver and the pump. We choose to vet these communications instead of implementing our own pump driver to give users the ability to use any pump driver of their choice. This also helps keep our trusted computing base small thereby easing formal verification.
Figure 4 provides an overview of our reference monitor. We only validate commands pertaining to insulin delivery since they pose the most dangerous threats to the user. From a high-level, the pump registers insulin commands with the reference monitor before issuing them to the pump driver. In the reference monitor, we check to ensure that the command issued by the pump driver is consistent with the command registered by the BioKernel. At the reference monitor, we also intercept all records of delivered insulin doses to ensure their integrity for closed loop calculations. On detecting manipulation from the pump driver, we transition insulin delivery to a default safe state (described in Section 7) where there is no immediate threat to the user. We do not validate communication pertaining to pump setup / UX since users can directly detect the respective manipulations. Figure 5 lists the different types of messages pertaining to insulin delivery exchanged between a pump driver and an insulin pump. We list other message types in Appendix B.
5.2 Design Principles
Avoid programming the pump. We avoid issuing insulin delivery commands to keep our implementation simple.
Ensure that the user always stays safe. Delay in delivering insulin does not pose any immediate threat to the user. However, delay in cancelling a wrongly issued insulin delivery command (such as one with a high insulin dose) can potentially kill the user. Thus, in the interest of keeping users safe, we violate the principle of not issuing commands to the pump in two specific cases of cancelling insulin delivery.
Disregard non-critical functionality in the interest of security. Some pump drivers incorporate an automatic retry mechanism when they’re unable to communicate with the pump. Such retry mechanisms complicate validation, since it becomes difficult to distinguish between duplicate commands issued as part of a retry versus duplicate commands issued as the result of a bug or an attack. We, thus, disregard such functionality by treating all duplicate commands as bugs or attacks.
Message Message
name description
Request insulin delivery
getStatus status from the pump
Insulin dose delivery
statusResponse information from the pump
Insulin dose delivery
errorResponse error from the pump
Set default rate of
setBasalSchedule insulin delivery
Set temporary rate
setTempBasal of insulin delivery
bolus Deliver instant insulin dose
cancelDelivery Cancel insulin delivery
Figure 5: Types of insulin related messages exchanged between between a pump driver and an insulin pump.
5.3 Reference monitor design and implementation
The BioKernel invokes methods exposed by the reference monitor to register commands to be sent to the pump before invoking the relevant methods in the pump driver. The BioKernel also registers two callbacks with the reference monitor, one to be invoked on successfully receiving a response from the pump and the other in case of error.
The pump driver then updates its internal states, encodes the command and its relevant parameters and sends it to the reference monitor. The reference monitor then checks that the pump driver has not added new commands or modified the parameters before encrypting and sending them to the pump. If the reference monitor detects added or modified commands, it transitions the pump to operate in a default safe state and rejects all further commands from the driver.
We employ a bounded delay of one second to detect dropped commands. Concretely, the reference monitor detects dropped commands if it does not receive anything from the pump driver within one second of registering a command from the BioKernel. It treats dropped commands the same way as added or modified commands by transitioning the pump to a default safe state and rejecting all further commands. We transition the user out of this safe state and start accepting commands only after the user manually configures the BioKernel to trust the driver again.
The reference monitor operates differently on detecting dropped commands that deal with cancelling insulin delivery. Users will seek to cancel insulin delivery if they accidentally triggered an insulin delivery command or incorrectly set its parameters. Since excess insulin injection is dangerous, the reference monitor would automatically send these commands to the pump before transitioning the pump to the default state and rejecting all further commands. There are two commands that deal with cancelling insulin delivery: cancelDelivery and setting the suspend option on setTempBasal. cancelDelivery seeks to cancel a previous insulin delivery command and (if the delivery is still ongoing) and the suspend option on setTempBasal seeks to suspend all insulin delivery until the user resumes delivery.
We focus our implementation on Loop’s OmniBLE [6] pump driver. OmniBLE allows iOS applications to communicate with OmniPod pumps over Bluetooth. We include OmniBLE’s cryptography and Bluetooth communication modules as part of the BioKernel and incorporate the reference monitor just before encrypting messages sent to the pump/ decrypting messages received from the pump.
OmniBLE does not incorporate automatic retry in its implementation. However, OmniBLE occasionally sends additional commands as part of its operation. For example, when sending a setTempBasal command, if OmniBLE’s internal states indicate that delivery is still in progress as part of a previous setTempBasal command, it first issues a cancelDelivery command to cancel that delivery. To remove ambiguity in terms of whether the cancelDelivery command was issued as part of setTempBasal, we enforce OmniBLE to always issue the cancelDelivery command. We evaluate GlucOS in terms of driver security in Section 10.4.
Overall, we employ straight-forward checks to detect manipulations from buggy/malicious pump drivers. However, the reference monitor’s design is more nuanced to handle the full range on insulin commands, missing out on any of which could potentially kill the user.
6 End-to-end security: Biological invariant
A key novelty in our overall approach is our introduction of a biological invariant that provides end-to-end verification. End-to-end verification in this context means that we verify the BioKernel’s glucose calculations against observations of the user’s metabolic state. If there is a large mismatch between our calculations and what we observe, then it means that our calculations are wrong. These deviations can happen from drastic changes in the user’s glucose absorption, like during exceptionally vigorous exercise or if the pump insertion site hits a vein and delivers insulin directly into the blood stream. Regardless of the cause, the biological invariant detects these conditions and prevents harm to the user.
The biological invariant closes the gap between verified software and human metabolism by incorporating a direct check against the user’s physiological data into the formal verification process. The biological invariant provides a runtime mechanism to detect when GlucOS’s calculations have become unsound. Specifically, the biological invariant compares the BioKernel’s theoretical calculation of the user’s blood glucose level against the actual observed blood glucose readings. Theoretical glucose levels provide an estimate of how the system thinks a given insulin dose will influence glucose levels. GlucOS uses its estimate of the human’s insulin sensitivity when calculating these doses. A significant divergence between the theoretical and observed values indicate a potential catastrophic failure mode, where GlucOS’s estimates of the user’s insulin sensitivity and calculations have diverged from the user’s real physiological state. The biological invariant allows the system to safely transition into a default safe state (Section 7). If the divergence between theoretical and observed glucose exceeds a threshold, indicating a loss of confidence in the model’s calculations, GlucOS discontinues automated insulin delivery and enters the safe state.
The biological invariant primarily focuses on resolving the undesirable situation where GlucOS believes that the person is less sensitive to insulin than they really are. In this case, the delivered insulin will remove more glucose than required, leading to potentially fatal low glucose levels. Consider the example of microbolusing when the user is above recommended glucose levels. If the system assumes a state where it believes the user is less sensitive to insulin than they really are and the user is 2x more sensitive, it will inject 2U of insulin when only 1U is required. While the system believes this will restore the user to safe levels, the actual glucose suppression will now push the user towards dangerously low glucose levels. Due to the higher clinical risk, we emphasize triggering the invariant in the undercalculation of insulin sensitivity case to transition to the default safe state. We set a threshold of 30 mg/dL for the maximum allowable divergence as shown in Figure 6 based on clinical evaluations showing detectability of significant glycemic deviations above this level[37, 32].
Figure 6: This figure shows an example of diverging differences between theoretical glucose and the observed glucose levels for a user. GlucOS will lose confidence in its calculationsand tranisition to a default safe state once it goes beyond a fixed difference threshold, where the biological invariant fails.
We formally model and verify the biological invariant using the Dafny language. The core aspects we specify include:
(a) A ghost variable representing the ground-truth physiological blood glucose level of the user. This models the actual, unobserved glucose state inside the human body that the system is trying to estimate and control.
(b) The theoretical calculation of the blood glucose level according to the system’s model. This captures the glucose prediction made by the system based on its assumptions about insulin dynamics, metabolism etc.
(c) The observed value of the blood glucose level from glucose readings from the user.
(d) The invariant property itself, which asserts that the absolute difference between the theoretical model glucose and the ground-truth physiological glucose cannot exceed a threshold of 30 mg/dL.
Our primary goal of verification in this segment is to make sure that there are no execution sequences that can violate the invariant without reaching a state where automated insulin delivery is disabled. This verification also ensures the biological invariant correctly captures our intention - if the system’s theoretical model deviates too far from the real physiological state, automated dosing is stopped to prevent potential over/under delivery of insulin. We also verify framing properties, showing the invariant’s state updates are precise and do not inadvertently modify other system variables. The verified Dafny specifications are then translated to Swift to integrate the biological invariant into the runtime system implementation for execution, with high assurance from the Dafny verification that they correctly implement the safe-guarding behavior in correspondence with the proven formal model.
Incorporating the biological invariant enables fundamental safety assurances not possible with traditional software verification alone. First, it constrains the set of possible erroneous system behaviors by restricting manual divergence between the theoretical model and actual physiological state. Section 10.5 shows that the biological invariant improves the system’s time-in-range glycemic control significantly compared to baseline when the system’s model and knowledge of user parameters becomes unsound. Second, it provides higher confidence of overall system reliability by directly incorporating physiological cross-checks into the verification process. We consider the biological invariant an important conceptual contribution that can extend the formal verification process to dynamically check and respond to deviations at runtime forming the complete bridge between the software and human physiology.
Traditional software verification techniques like model checking [35] and deductive verification [24] have been explored in the past for hybrid system verification [23] that interact with the human body, but they do not completely account for physiological variability between individuals or model inaccuracies during the insulin delivery process.
7 Runtime failure management
Figure 7: GlucOS’s state machine to transition users to the safest manual dosing state when losing trust in insulin delivery (i.e., with a buggy or malicious pump driver) or when the biological invariant is not satisfied.
In this section, we describe GlucOS’s state machine that combines its security components, i.e., algorithmic security (Section 4), driver security (Section 5) and end-to-end security (Section 6) together to protect users in light of their corresponding failures.
GlucOS’s state machine consists of three operating states: (1) ML with safety operating in closed loop, (2) reactive safe model operating in closed loop and (3) manual dosing in closed loop. ML provides the tightest level of control in managing glucose but poses the largest immediate threat from excess insulin since it preemptively injects insulin by anticipating rise in glucose. The reactive safe model provides slightly lesser control, but poses milder immediate threat since it only injects insulin to account for measured excess glucose. Manual dosing offers the least amount of control but poses no threat since it only injects the minimum insulin needed for sustenance.
GlucOS predominantly operates in the state running ML (with safety) in closed loop. Certain real-world failures (such as missing CGM readings) transition GlucOS to operate in the state running the reactive safe model in closed loop. However, such failures are simple and we discuss ways to handle them in Appendix C. In this section, we focus on transitions pertaining to the manual dosing state.
We transition to the safest manual dosing state when we do not trust insulin delivery (i.e., with a buggy or malicious pump driver). We also transition to the safest manual operation when the biological invariant is not satisfied to ensure that we do not risk the possibility of removing too much glucose from our insulin doses. If we end up in the manual dosing state, we enforce that only manual user intervention can transition back to other states. We visualize these state transitions in Figure 7.
8 Formal verification
A single error or vulnerability in an automated insulin delivery system can have life-threatening consequences for the user. This criticality motivates our use of formal verification in our development and testing process. Through formal verification, we aim to confirm two critical aspects of our insulin delivery system. First, we validate the goals and claims behind our critical insulin delivery functions. Second, we verify that GlucOS applies the correct state transitions in precarious scenarios to protect the user when the system loses confidence in one or more of its components due to miscalculation or malicious behavior.
Systematically identifying the invariants and postconditions in alignment with our claims that we verify for a critical function in our system is important. Our first step is to clearly define what the function is supposed to do. This helps list all the inputs that the function takes, including their types and ranges. Then, we define the expected outputs, including their types and valid ranges. Determining safety and domain-specific constraints for each function plays a major role in ensuring that the function never produces an unexpected output that puts a user at risk. For example, a general constraint would be that a basal rate should never be negative as it is invalid. A domain-specific constraint is defining a maximum basal rate threshold that is in accordance with medical guidelines. For a full list of the invariants that we define and prove, see Appendix D.
These steps help define the preconditions, postconditions and invariants that help prove the properties of our verified functions. We run formal verification in Dafny to verify our claims on how a function works. After we verify a function, we port the code to Swift to make it part of the BioKernel.
We verify invariants and postconditions for the system’s critical functions in the aforementioned approach. For instance, we verify the microbolusing function to ensure a minimum time interval between consecutive micro-boluses to prevent frequent dosing in short intervals (within 4 minutes), confirm that the user glucose level is significantly above the target(atleast 20 mg/dL) before administering a bolus and that the bolus amount is within safe and predefined limits. These are claims about our function that we formally verify to assure that they are always correct.
In another example, we prove correct our insulin delivery calculations to ensure that we maintain accurate accounting of insulin delivery. For this verification task, we verify our insulin delivery over a duration of time function to guarantee non-negative results, correctly handle zero duration or non-overlapping time segments, making sure that the dosage does not exceed recommended limits, and accurately calculate insulin delivery based on time segment overlap.
We perform this type of verification on all the critical functions involved in the insulin delivery process. This high-level of assurance helps us trust the the decisions that GlucOS makes.
We also formally verify the state machine we use for defining system transitions under undesirable scenarios (Section 7). We discover several potential issues in the state transition logic and function design during verification. For example, there were cases where the system could transition to an unsafe state under certain conditions, potentially leading to excessive insulin delivery. We also identified bugs related to violating valid bounds and domain-specific constraints. A more detailed discussion of the bugs have been included as part of subsection  10.6 in our evaluation.
9 Implementation
In our implementation, the BioKernel is an iOS app on iPhone or iPad hardware. It communicates with a Dexcom G7 CGM and Omnipod DASH insulin pump via Bluetooth low energy and runs the core closed-loop algorithm. For the BioKernel, we use the Swift programming language, the Actor abstraction for data-race-free concurrency, and JSON files stored on disk to capture persistent state.
Although the BioKernel is a fully featured and stand alone automated insulin delivery system, GlucOS uses an event logging system that enables separate apps to implement complementary features while remaining isolated from the BioKernel. We implement our event logging system using a Google Cloud App Engine app and the BioKernel produces logs that other apps consume. The other apps include an app for alerting when the individual is likely to experience hypoglycemia or hyperglycemia, an app to replay closed-loop runs to verify the results, and an app for running simple ML on data to calculate therapeutic settings.
To keep predictive ML and the pump driver isolated, we use the Swift type system and object boundaries to maintain isolation. Our architecture also supports the use of process boundaries for even strong isolation and to implement OS-level sandboxing. iOS lacks support for proper process APIs (fork, exec etc), so we use the Swift type system instead.
For CGM and insulin pump drivers, we use LoopKit [41] and isolate untrusted portions using the Swift type system.
10 Evaluation
Given the risks associated with running ML for automated insulin delivery, we answer most questions on 21 virtual humans from an FDA-approved simulator [9]. The simulator generates CGM readings across different age groups including children, adolescents and adults. We also evaluate security mechanisms on virtual patients to see if they can generalize to individuals with different physiologies. We present experiences of running GlucOS with a real individual in the next section.
10.1 Does ML lead to better outcomes for managing T1D?
We confirm findings from prior research that ML leads to better outcomes [47, 66, 34] by comparing the average proportion of time spent by virtual patients in range when using our reactive safe model (Section 4.2) against our predictive ML model (Section 4.3). As mentioned in Section 2, the American Diabetes Association (ADA) recommends that an individual with T1D should spend more than 70% of a day (also called TIR) [8, 21] in order to remain healthy.
We train separate predictive ML models (feed-forward neural networks trained to minimize MSE loss) for each virtual patient and similarly tune the physiological parameters of the reactive safe model to suit each patient. We ensure that we run the reactive safe model and the predictive ML model on identical conditions within the simulator.
Patient TIR with TIR with
type ML model reactive safe model
Adults 93.49% 88.25%
Adolescents 85.23% 81.98%
Children 86.6% 77.31%
Figure 8: Average proportion of time spent in range by 21 virtual patients in identical scenarios with a our predictive ML model and our reactive safe model.
We show the average proportion of time spent in range by virtual patients from each age group (adults, adolescents and children) in Figure 8. We can see that ML helps virtual patients across all age groups spend a larger amount of time in range, thereby showing that ML leads to better outcomes for managing T1D.
10.2 What are the risks associated with untrusted ML algorithms for managing T1D?
We assess the impact of predictions made by intentionally malicious ML algorithms on managing T1D. Since GlucOS is designed to be flexible with any predictive algorithm, it could be subject to attacks from malicious ML algorithms. We intentionally introduce an order of magnitude difference to the amount of insulin predicted by the predictive ML model in Section 10.1. In one experiment, we dose one-tenth of the amount of insulin and in another experiment, we dose ten times the amount of insulin. We run these malicious ML algorithms within the simulator on each virtual patient on the same scenarios as the previous experiment.
Figure 9 and Figure 10 show the CGM readings on 3 randomly chosen virtual patients on a particular scenario for the experiments where we dose less and dose more respectively. When dosing less, we see that all patients spend most of their time above Level 1 hyperglycemia which would have long-term effects on their well-being. When dosing more, all patients fully go below Level 2 hypoglycemia, which would have killed them.
Figure 9: Variation in CGM readings for 3 virtual patients when running a predictive ML algorithm that intentionally doses one-tenth the amount of the required insulin. The CGM readings staying mostly above the Level 1 hyperglycemic range would have drastic effects on well-being.
Figure 10: Variation in CGM readings for 3 virtual patients when running a predictive ML algorithm that intentionally doses ten times the amount of the required insulin. The CGM readings staying below the Level 2 hypoglycemic range would have likely been fatal to all 3 patients.
Across all virtual patients, we see an average drop in TIR from 88.44% with correct dosing to 23.32% when dosing less and an average drop in TIR from 88.44% to 0% when dosing more. Correspondingly, when dosing less, we observe an average increase in the timein hyperglycemia from 3.77% with correct dosing to 65.05%. Similarly, when dosing more, we observe an average increase in the time in hypoglycemia from 0% to 100%.
These results further motivate the need for security mechanisms when providing individuals with the flexibility to use any algorithm to manage T1D.
10.3 What is the impact of GlucOS’s algorithmic security mechanism on managing T1D?
GlucOS’s algorithmic security mechanisms should keep the blood glucose of individuals in range even in presence of malicious ML algorithms. The mechanisms should also ensure that they do not significantly dampen the level of control provided by an efficient algorithm.
Impact on malicious algorithms. We repeat the experiments from the previous section with the same scenarios. We run the same malicious algorithms, but also run the reactive safe model to set bounds on the amount of insulin the malicious algorithms can inject. Figure 11 and Figure 12 show CGM readings on the same 3 random virtual patients on the same scenarios from the experiment in the previous section. When running either malicious model, we see that all patients spend over 70% of their time in range. These results show that GlucOS protects individuals even in presence of intentionally malicious algorithms. On average, across all virtual patients, we still see a drop in TIR from 88.44% with correct dosing to 73.89% when employing safety while dosing less and a drop in TIR from 88.44% to 78.22% when dosing more. While the times spent in range is less than correct dosing, the impact of incorrect dosing is significantly dampened with patients spending the recommended amount of time in range.
Figure 11: Variation in CGM readings for 3 virtual patients when clamping an algorithm that intentionally tries to dose one-tenth the amount of the required insulin. On average, the CGM readings stay in range over 70% of the time.
Figure 12: Variation in CGM readings for 3 virtual patients when clamping an algorithm that intentionally tries to dose ten times the amount of the required insulin. On average, CGM readings stay in range over 70% of the time.
Patient TIR with clamped
type Predictive ML model
Adults 92.88%
Adolescents 82.38%
Children 85.42%
Figure 13: Average proportion of time spent in range by 21 virtual patients in the same scenarios in Section 10.1 when clamping the predictive ML with the reactive safe model.
Impact on effective benign algorithms. We evaluate the impact of GlucOS’s algorithmic security mechanisms on benign algorithms such as the predictive ML model used in Section 10.1. We repeat the experiment of running the predictive ML model on all virtual patients in the same random scenarios while maintaining the same parameters for the clamps and reactive model used for the malicious algorithms evaluation.
Comparing Figure 13 against Figure 8 we see that there is only a slight drop in the average proportion of time spent in range by virtual patients across all age groups when employing GlucOS’s algorithmic security mechanisms. The average proportion of time spent in range being higher than the reactive safe model shows that GlucOS’s security does not come at the cost of the benefits from ML.
Figure 14: We attached an Omnipod DASH insulin pump to a stuffed animal and modified the OmniBLE pump driver to intentionally add, drop and modify insulin commands. Our driver security mechanism detected all manipulations.
10.4 Does GlucOS ensure pump operation while overcoming buggy/malicious pump drivers?
We first evaluate GlucOS’s driver security mechanism in terms of its impact on benign pump drivers that accurately relay pump communications. We then evaluate the impact on detecting additional, modified or dropped communication from a malicious pump driver. Our mechanism should ensure that all communications go through with a benign pump driver and should catch all manipulated communication with a malicious pump driver. We do not use the simulator for this evaluation since it does not incorporate a pump driver.
Impact on malicious pump drivers. As shown in Figure 14, we attached an Omnipod DASH insulin pump to a stuffed animal. We then modified the source code of the OmniBLE pump driver to intentionally add, drop and modify insulin commands. We detected all manipulated pump commands showing that our driver security mechanism has 0 false negatives in detecting manipulated commands.
Impact on benign pump drivers. For one week, we logged messages from the BioKernel to unmodified OmniBLE as well as messages sent from OmniBLE to the pump on an individual running GlucOS (See Section 11). We report 100% consistency between the commands sent to OmniBLE and commands issued from OmniBLE showing that GlucOS’s driver security mechanisms do not incur any false positives.
These results show that GlucOS can ensure pump operation while withstanding attacks from malicious drivers.
Patient TIR without TIR with
type biological invariant biological invariant
Adults 49.93% 88.92%
Adolescents 55.83% 75.26%
Children 67.41% 77.45%
Figure 15: Average proportion of time spent in range by 21 virtual patients when fluctuating insulin sensitivities. We compare cases with and without our end-to-end security mechanism using the biological invariant.
10.5 Is the biological invariant effective?
We run experiments on the simulator to evaluate if GlucOS’s use of the biological invariant to transition to manual insulin dosing leads to better T1D management. Since the simulator does not provide support to fluctuate insulin sensitivity, we intentionally scale up or scale down values for insulin sensitivity when calculating insulin doses in closed loop. We trigger cases where we scale down the insulin sensitivity to simulate scenarios where the system assumes a lower sensitivity value than the user’s actual sensitivity more frequently since this are more relevant to evaluating the biological invariant.
Across all virtual patients, we record an average TIR of 57.56% when fluctuating insulin sensitivity without incorporating the biological invariant. In contrast, we record an average TIR of 80.54% when incorporating the biological invariant to transition to manual dosing. Figure 15 shows the average TIR across age groups. We clarify that we use the correct insulin sensitivity values to compute bolus doses when dosing manually, since users would know their true insulin sensitivity when taking in manual doses.
10.6 Did formal verification help with correctness?
During the formal verification process of our secure insulin delivery system, we uncovered various bugs and inconsistencies that could have compromised the safety and reliability of the system in niche scenarios. We identified a total of 9 bugs across different categories. 3 of these bugs were related to invalid ranges for variables such as negative value returns, domain-specific out of range values, potential division by zero errors etc. in the insulin dosing calculations which were resolved by adding appropriate checks and error handling mechanisms to prevent such operations and ensure graceful handling of these cases. Additionally, we found 4 instances where the implemented functions violated their intended postconditions along the insulin delivery path, which were addressed by thoroughly reviewing and modifying the implementations to correctly satisfy the postconditions, preserving the intended behavior and meeting our claims for critical funcrions. Moreover, our analysis uncovered 2 situations where certain functions lacked proper error handling mechanisms, potentially leading to undefined behavior or system failures. By addressing these bugs and incorporating formal verification throughout the development lifecycle, we significantly enhanced the robustness, reliability, and safety of our secure insulin delivery system.
11 Using GlucOS in the real world
In this section, we report on our experiences working with one individual, Bob, since November 2023. We show our results for using GlucOS with our reactive safe model closed-loop algorithm as well as with our predictive ML model closed-loop algorithm paired with our security mechanism. We also measure system performance, report on errors, and quantify our implementation simplification efforts.
Overall, our study is a correlation study, where we evaluate one individual using our system and compare segments of time with each other.
11.1 Ethical considerations
In this section, we outline the steps we take to ensure that we uphold high ethical standards and ensure that the human using our system, who we will call Bob, is safe.
First, one of the co-authors of this paper is a board-certified Endocrinologist, who specializes in T1D. They are both a clinician with an active practice seeing patients living with T1D and a researcher who focuses on research on T1D. They designed the safety protocol that we detail in Appendix A, which defines the testing we conduct to ensure that the individual using our system is safe, and the specific criteria we use to stop the study, if needed.
Second, we submitted our study to our university’s IRB, which determined that our “self evaluation” is exempt from IRB (case number omitted, but available).
Third, Bob increased the frequency of his visits to his Endocrinologist, which includes lab work to confirm that he remained in good overall metabolic health.
Overall, GlucOS system had a large positive impact on the individual who used it. Bob’s lab-based A1C test (a blood test that measures average glucose levels over time), which includes his time using GlucOS, was 5.8%. That is nearly non-diabetic levels of control – healthy individuals will have an A1C of 5.6% or lower. This result is a profound outcome for Bob, and GlucOS had a massive positive impact on Bob’s health and quality of life.
11.2 Is an ML-based closed-loop system safe and effective in practice?
Bob has been using GlucOS in a closed loop since November 23rd, 2023. From November 23 to January 28th, 2024 we used our reactive safe model to make control decisions. Starting on January 28th Bob started to use our predictive ML model for periods of time to get used to the new system, and then starting on January 31st, 2024 starting using our predictive ML model exclusively. We report on one week’s worth of data from the predictive ML model starting on January 31st ending on February 6th, 2024.
Glucose Range Reactive Reactive ML
range (mg/dl) label 67 days last wk. 1 wk
<54 L2 Hypo $0$%
54-69 L1 Hypo 0.59% 0.48% 0%
70-180 In range 96.35% 91.90% 97.22%
181-250 L1 Hyper 2.79% 7.62% 2.78%
>250 L2 Hyper 0.18% 0% 0%
Figure 16: The percent of time spent in different glucose ranges for the reactive safe model and the predictive ML model. Bold values represent the best results for each range.
Figure 16 shows the amount of time Bob spent in different glucose level ranges over the duration of our study. Data from the reactive safe model covers 67 days, followed by one week of data from our predictive ML model. We also show data for the last week of using the reactive safe model to show trends over time. Overall, Bob had the best results when using ML.
Next, we measure how often the predictive ML model was responsible for programming the pump vs our reactive safe model. For the one week when Bob was using the predictive ML model, 33.0% of the time both the predictive ML model and reactive safe model issued the same commands, 39.9% of the time the BioKernel used the predictive ML model’s commands, and 27.1% of the time the BioKernel used the reactive safe model’s commands. We believe that this distribution represents a suitable balance for providing the ML with enough flexibility to improve outcomes, while having tight enough bounds to limit potential damage from bad or malicious ML predictions.
11.3 How does the BioKernel perform?
We look at two aspects of system performance when evaluating our real-world use of GlucOS. First, we measure how long our closed-loop algorithm takes to run and second, we count how often the closed-loop algorithm runs successfully end-to-end. We use Bob’s iPhone 14 to collect all of the data we show here and these numbers are averages across the entire six month period.
On average, the BioKernel spends 4.3s running the closed loop algorithm, where almost half of that time is dominated by CGM and pump communications. Of the components within the closed-loop algorithm, the ML predictions take 2.2s on average and the safety logic takes 19ms (0.019s). Since the system runs in the background, we believe that iOS runs the device with reduced computational capacity to save power. But overall, 4.3s of wall-clock time for every five minutes our closed-loop algorithm runs with only a tiny fraction of that coming from our safety logic represents practical system performance.
For successful closed-loop end-to-end execution, over the evaluation period the closed-loop algorithm runs successfully 96.5% of the time, with 2.1% of the runs failing due to pump communication failures, which is our largest class of failure.
11.4 How simple is our BioKernel?
To evaluate the complexity of our BioKernel, we count the lines of code in our implementation and compare against Loop, another open-source automated insulin delivery system.
Using the “cloc” utility the BioKernel app has 4.0k lines of code compared to 39.8k lines of code in the Loop app, an order of magnitude reduction. We omit the lines of code coming from LoopKit, which both the BioKernel and Loop use, because our pump security measures remove most of it from our trusted computing base. However, LoopKit and the drivers have a substantial amount of code, weighing in at 70.1k lines of code, showing the importance of our security measures.
To explain why we have such a large reduction in source code, we outline the differences between the two systems. First, Loop has several features that we move outside of the BioKernel. These features include a remote interface for insulin dosing, a Watch app, Siri command interfaces, third party libraries, tutorials, complex meal announcements, and physiological simulations. Second, we simplify our implementation of features that are shared between the BioKernel and Loop. These shared features include a simplified local storage implementation, simplified concurrency support, minimal therapeutic settings, and a simplified closed-loop algorithm. As our evaluation shows, we can still provide tight control over T1D with this simplified core system and still support most or all these additional features architecturally.
12 Related work
Previous research has looked at the security of implanted medical devices in general [29, 16, 54], in addition to looking at insulin pumps in particular [39, 50], with more recent work looking at providing improved security [44, 11]. Also, recent work has looked at applying formal methods to insulin pumps for high assurance [49]. These works focus on the device and their communication channel. In contrast, with GlucOS we assume that these devices are correct and secure and focus our efforts on the software we use to run the automated insulin delivery system.
Our work on coping with malicious pump drivers builds on previous work on malicious device drivers in operating systems [62], [15], [27], [28], [43], [31]. Our work applies these principles to a new domain: automated insulin delivery systems.
13 Conclusion
Our audacious long-term goal is to turn a T1D diagnosis from a death sentence into an indicator of longevity, where people living with T1D will be expected to live longer than their healthy peers. This longevity will come by virtue of the tight control that they maintain over their metabolism through advanced computer systems. Our first step towards this goal is to ensure that people can use trustworthy computer systems to manage their glucose.
Appendix A Protocol for real world use
One individual, who we will call Bob, used GlucOS for 2.5 months. We put in place a protocol where we would stop our study and ask Bob to cease using our system and revert back to his previous insulin regime. In our protocol, we have formal in-person meetings with Bob every week to check his data and make sure that he is not incurring risk as a result of using GlucOS. We would have stopped our study if any of the following conditions occurred:
  • Excessive hypoglycemia. We define excessive hypoglycemia as having spent 5% or more of time during a week with CGM readings below 70 mg/dl or more than 1% below 54 mg/dl. These amounts of time spent in hypoglycemia would have represented a significant increase in time spent in hypoglycemia for Bob.
  • Insufficient CGM data. If Bob would have had less than 70% of the time during a week with CGM readings, then the lack of CGM data would represent a decrease for Bob.
  • Insufficient closed-loop runs. We designed GlucOS to run closed-loop algorithms every five minutes, or 288 times per day. If Bob had a day where our closed-loop algorithm ran successfully less than 200 times, it would represent a fundamental flaw in the system.
  • Serious complications due to diabetes. Serious complications include any hospitalization for diabetes related issues, severe hypoglycemia where Bob was unable to recover from hypoglycemia himself and had to get help from someone else, diabetic ketoacidosis, or seizure.
  • New medical diagnosis requiring significant attention. If Bob had received a new medical diagnosis during the study, the study would have had the potential to be a distraction for Bob.
  • New mental health diagnosis. If Bob were struggling with mental health during the study, the study would have the potential to be a distraction.
Appendix B List of pump setup and UX communication messages
Message Message Message
name type description
setupPod Setup Setup a new pump
assignAddress Setup Pair pump with phone
deactivatePod Setup Teardown pump
acknowledgeAlert UX Acknowledge alert from pump
configureAlerts UX Configure alerts from pump
beepConfig UX Configure pump beeps
Figure 17: Types of messages exchanged between a pump driver and pump pertaining to setup/UX.
We list insulin message types pertaining to pump setup and UX in Figure 17.
Figure 18: GlucOS’s CGM state machine to handle missing readings from CGMs.
Appendix C CGM state machine
Inability to retrieve readings from a CGM also count as runtime failures in addition to those discussed in Section 7. However, these failures are easier to handle. In this section, we describe the CGM state machine that handles such failures. We note that transitions from the state machine described in Section 7 supersede state transitions from the CGM state machine.
Since ML needs a stream of CGM readings as input, we transition from ML to the reactive safe model if we don’t receive an input from the CGM. As long as the number of missing CGM readings is low (say, within 30 minutes), we can interpolate to account for the missing readings to operate the reactive safe model. Once we have too many missing CGM readings that make us lose confidence in interpolation, we transition to manual dosing. From manual dosing, we transition back to the reactive safe model on receiving a CGM reading. Lastly, from the reactive safe model, we transition back to ML once we have enough CGM readings that are required by the ML model. Figure 18 summarizes these transitions.
Bug Type Description
Division by Zero Error Potential division by zero error when calculating insulin dosage if the user entered the same start and end time for a basal rate, resulting in a duration of zero.
Negative Value Handling The insulin dosing calculation for microbolusing and basal rate erroneously allowed negative values.
Out-of-Range Value Handling The microbolusing function failed to handle cases where one of its calculation components was out of the expected range or resulted in a negative value.
Violation of Intended Postcondition The function responsible for setting the insulin delivery duration violated its intended postcondition by allowing the end date to be earlier than the start date.
Violation of Intended Postcondition The function responsible for calculating the total insulin dose violated its intended postcondition by returning an incorrect upper bound value under certain conditions.
Violation of Intended Postcondition The function responsible for updating the user’s insulin dosage violated its intended postcondition by failing to validate the input parameters correctly.
Violation of Intended Postcondition The function responsible for calculation of ‘unitsDelivered‘ needed additional postconditions accurately reflect the amount of insulin delivered over the time gap.
Lack of Error Handling Additional error handling was added to insulin delivered over a time segment function to gracefully catch undesirable state.
Lack of Error Handling Additional error handling was added to microbolusing function to gracefully catch undesirable states.
Figure 19: Report of bugs fixed by formally verifying GlucOS.
Appendix D Formal verification: invariants and bugs
In this section, we list the invariants used to formally verify critical functions of GlucOS. Figure 19 lists the bugs we found and fixed during formal verification.
D.1 Invariants
The MicroBolusing function ensures that micro boluses are always administered at safe time intervals while making sure that they are actually needed (the glucose level of the person is in fact high).
  • Ensure no micro bolus in the last 4.2 minutes.
  • Ensure glucose is at least 20 mg/dl above the target glucose level.
  • Ensure predicted glucose is greater than the current glucose level minus 2.
  • Ensure micro bolus amount is within the allowed range.
  • Ensure bolus amount is rounded correctly if pumpManager is available.
We assume that the BioKernel is the only software issuing bolus commands.
The GuardRails function is responsible for ensuring a safe basal rate value, mindful of the user’s glucose levels. It enforces several invariants:
  • Ensure the result of rounding newBasalRateRaw to the supported basal rate is within the allowable range.
  • Ensure the new basal rate does not exceed the maximum basal rate in the settings.
  • Ensure the new basal rate is not negative.
  • Ensure that if either the current glucose level or the predicted glucose level falls below or equals the shut-off glucose threshold, the new basal rate is set to 0.0.
The function operates under the assumption that all paths must go through the guard rails, with a specific evaluation comparing with the pump invariant. Additionally, it assumes the integrity of settings, trusting the underlying OS and file system and storage stack.
The insulinDeliveredForSegment function ensures that intersectionStart is not greater than intersectionEnd when computing the intersection between self.startDate and self.endDate with startDate and endDate.
  • Verifies that intersectionDuration is non-negative and does not exceed doseDuration, ensuring it accurately reflects the overlapping time duration.
  • Checks that the units of insulin (units) used for calculation are either deliveredUnits or programmedUnits, ensuring consistency in how the insulin amount is derived.
  • Verifies that the computation (units * intersectionDuration / doseDuration) accurately represents the amount of insulin delivered during the intersection period.
The function assumes that the input parameters (self.startDate, self.endDate, startDate, endDate, units, and doseDuration) are valid and within the expected ranges. It is also assumed that the underlying data structures and calculations are accurate and consistent throughout the system.
The CreateBasalDose function requires a time gap greater than 1 second to proceed and return a valid DoseEntry.
  • Ensures that the calculation of basalRatePerSecond from basalRate accurately represents the rate of insulin delivery per second.
  • Ensures that the calculation of unitsDelivered accurately reflects the amount of insulin delivered over the time gap.
  • The DoseEntry constructed should have consistent attributes and adhere to the specified type, units (unitsPerHour), and mutability (false).
The function assumes that the input parameters (basalRate, startDate, and endDate) are valid and within the expected ranges. It also assumes that the underlying time calculations and conversions between units are accurate and consistent.
The inferBasalDoses function ensures that basalDoses contains only DoseEntry instances where type is .tempBasal, .resume, or .suspend, sorted in ascending order by startDate.
  • InsulinType should be determined correctly based on the last .tempBasal or .bolus type dose in doses, defaulting to .humalog if none are found.
  • Each inferred basal dose added to inferredBasalDoses must be created using the createBasalDose function with valid parameters and added in the correct chronological order.
  • If the last dose in basalDoses is not of type .suspend, a valid basal dose must be inferred from its endDate to at.
The function assumes that the input parameters (doses, basalDoses, at, and pumpRecordsBasalProfileStartEvents) are valid and within the expected ranges. It also assumes the correctness of the underlying data structures and the createBasalDose function.
The insulinOnBoard function ensures that doses contains unique and valid DoseEntry instances filtered up to at.
  • The total iob should accurately represent the cumulative insulin on board from all valid dose entries in doses.
  • basalDoses should contain inferred basal doses that accurately reflect insulin on board contributions when pumpRecordsBasalProfileStartEvents is false.
  • The final return value of insulinOnBoard should be the sum of iob and basalIob, representing the total insulin on board at at.
The function assumes that the input parameters (doses, at, and pumpRecordsBasalProfileStartEvents) are valid and within the expected ranges. It also assumes the correctness of the underlying data structures and the inferBasalDoses function.
The ActiveBolus function ensures that data is read consistently from disk before proceeding with any operations.
  • Verifies that DeduplicatedDoses(events: eventLog, at: at) correctly filters out duplicates and retains only relevant dose entries up to at.
  • Ensures that doses.filter accurately selects bolus entries that were active at the specified at.
  • Verifies that doses.last correctly returns the last active bolus entry, ensuring it is non-null if a bolus exists within the specified criteria.
The function assumes that the input parameters (eventLog and at) are valid and within the expected ranges. It also assumes the correctness of the underlying data structures and the DeduplicatedDoses function, as well as the consistency and reliability of disk read operations.
The TempBasal (Safety Clamps) function ensures that the events array includes only events from safetyStates that fall within the time range [start - duration, at).
  • Verifies that the historicalMlInsulin value represents the total units of insulin delivered by the machine learning system over the specified time horizon.
  • Ensures that the machine learning and safety temp basal rates are correctly converted to units of insulin based on the given duration.
  • Ensures that the upper and lower bounds for insulin units are correctly adjusted based on historicalMlInsulin.
  • Ensures that the difference between machine learning and safety temp basal units is correctly clamped within the calculated bounds.
  • The clamped delta units should be converted back to a temp basal rate and added to the safety temp basal rate.
  • The returned SafetyTempBasal object should correctly encapsulate the adjusted temp basal rate and historical machine learning insulin.
The function assumes that the input parameters (safetyStates, start, duration, at, mlTempBasal, safetyTempBasal, and lastHistoricalMlInsulin) are valid and within the expected ranges. It also assumes the correctness of the underlying data structures and the conversion functions between units, as well as the proper configuration and appropriateness of the safety clamp parameters and thresholds for the user.

### v2

@@
 Abstract
+We present GlucOS, a novel system for trustworthy automated insulin delivery.
+Fundamentally, this paper is about a system we designed, implemented, and deployed on real humans and the lessons learned from our experiences.
+GlucOS combines algorithmic security, driver security, and end-to-end verification to protect against malicious ML models, vulnerable pump drivers, and drastic changes in human physiology.
+We use formal methods to prove correctness of critical components and incorporate humans as part of our defensive strategy.
+Our evaluation includes both a real-world deployment with seven individuals and results from simulation to show that our techniques generalize.
+Our results show that GlucOS maintains safety and improves glucose control even under attack conditions.
+This work demonstrates the potential for secure, personalized, automated healthcare systems.
+1 Introduction
 Type 1 Diabetes (T1D) is a metabolic disorder where an individual’s pancreas stops producing insulin.
 To compensate, they inject synthetic insulin.
-Computer systems, called automated insulin delivery systems, exist that inject insulin automatically.
-However, insulin is a dangerous hormone, where too much insulin can kill people in a matter of hours and too little insulin can kill people in a matter of days.
-In this paper, we take on the challenge of building a new trustworthy automated insulin delivery system, called GlucOS.
-In our design, we apply separation principles to keep our implementation simple, we use formal methods to prove correct the most critical parts of the system, and we design novel security mechanisms and policies to withstand malicious components and attacks on the system.
-We report on real world use for one individual for 6 months using GlucOS.
-Our data shows that for this individual, our ML-based algorithm runs safely and manages their T1D effectively.
-We also run our system on 21 virtual humans using simulations and show that our security and safety mechanisms enable ML to improve their core T1D measures of metabolic health by 4.3% on average.
-Finally, we show that our security and safety mechanisms maintain recommended levels of control over T1D even in the face of active attacks that would have otherwise led to death.
-GlucOS is open source and our code is available on GitHub.
-1 Introduction
-Type 1 diabetes (T1D) is an autoimmune disease in which an individual’s pancreas stops producing insulin.
-People living with T1D must therefore take synthetic insulin and play the role of their pancreas.
-However, replicating the pancreas’s native insulin release patterns is extremely difficult, and optimal T1D self-management often requires hourly adjustments in insulin dosing both in response to physiological changes and in prediction of upcoming carbohydrate intake and energy expenditure (e.g., exercise, stress) [51, 56].
-Even when following this labor-intensive process of T1D self-management, affected individuals remain at risk for acute episodes of hypoglycemia (low blood sugar, which can cause loss of consciousness, seizure, or death) or ketoacidosis (which carries an associated risk for multi-organ injury), as well as for long term vascular complications (e.g., kidney damage, nerve damage, blindness, heart attack, stroke) that are a result of chronic hyperglycemia (high blood sugar) [46].
-8.4 million people are living with T1D worldwide [25].
-(a) Devices for automated insulin delivery. This figure is from jdrf.org.
+Computer systems, called automated insulin delivery systems, exist that use subcutaneous sensors to continuously monitor glucose concentrations (glucose is the body’s primary energy source) and inject insulin automatically using an insulin pump device (Figure 1).
+However, insulin is a dangerous hormone where too much insulin can kill people in a matter of hours [22] and too little insulin can kill people in a matter of days [13].
+For automated insulin delivery systems, the key challenge is to maintain a balance: provide enough insulin to prevent dangerously high glucose levels while avoiding excessive insulin that could lead to life-threateningly low glucose levels.
+In this context, security means ensuring integrity for insulin dosing against malicious attacks or inadvertant errors.
+Several commercial [17, 5, 3] and open-source [42, 54] automated insulin delivery systems exist today, but none of them consider security as a primary design constraint.
+(a) Devices for automated insulin delivery. This figure is from breakthrought1d.org.
 (b) Closed-loop architecture.
 (c) Our system for automated insulin delivery.
-Figure 1: Overview of automated insulin delivery devices, closed-loop architecture, and apps.
-The last two decades have witnessed the development of therapeutic technology to assist T1D management, including continuous glucose monitoring (CGM) devices and insulin pumps.
-Most recently, the integration of CGM sensors and insulin pumps into “artificial pancreas” or automated insulin delivery systems has enabled a simultaneous improvement in health outcomes and reduction in the burden of care for T1D-affected individuals [38].
-However, existing automated insulin delivery systems are complex, and there is no way to assure beyond any doubt that they are implemented correctly.
-For example, techniques from formal methods have not been applied to verify automated insulin delivery systems.
-With little reason to trust the systems and their core algorithms, users and developers are also reluctant to incorporate experimental algorithms into the prediction loop, such as advanced ML-driven insulin delivery.
-Researchers have shown that ML can improve automated insulin delivery systems [64, 33, 63, 59, 65, 58], yet the lack of security and safety for these ML predictions keeps people from using it.
-If ML is going to be part of medical decisions, we need the security community to lead the way on how to do this safely.
-In this paper, we argue for applying traditional security and operating systems design principles to make automated insulin delivery systems more trustworthy.
-This includes going beyond just patient privacy or medical device security, to systems security for automated insulin delivery systems to ensure that they are implemented correctly and that the incorporation of ML does not jeopardize their safety.
+Figure 1: Overview of automated insulin delivery devices, closed-loop architecture, and a screenshot of GlucOS.
+The complexity of automated insulin delivery systems and the fact that they make real-time healthcare decisions, makes their security paramount.
+Current research on security for these systems focuses primarily on evaluating algorithms using simulations [71, 72].
+We argue that this approach falls short of addressing the full spectrum of challenges inherent in building trustworthy systems that operate within real-world constraints.
+Real-world systems must cope with runtime platform limitations, changes in human physiology, and the necessity of human interaction with the system.
+Plus, algorithms and simulations model attacks abstractly through their glucose response, missing out on details like which components of the system are likely to be compromised or how the system architecture can prevent damage from these potentially vulnerable components.
+Our experience shows that when faced with these practical challenges, algorithms designed in simulations often prove insufficient or inapplicable.
+These real-world complexities require a fundamentally different approach to designing security for automated insulin delivery systems.
+To address these challenges, we need practical systems that move beyond simulation-based designs and tackle the challenges of real-world systems.
+Current open-source automated insulin delivery systems are complex.
+For example, the OpenAPS [54] core function for calculating insulin doses consists of 1192 lines of Javascript code, 63 input and configuration parameters, and 90 branch statements.
+All of this logic influences the ultimate dosing decision, making reasoning about its correctness, both manually and automatically, challenging.
+Automated insulin delivery systems that use the OpenAPS algorithm embed it into native apps using a WebView for Javascript interpretation, adding to the overall complexity and increasing the attack surface.
+In contrast, our implementation of this same function is 25 lines of native and formally verified Swift code.
+None of the open-source automated insulin delivery systems use formal methods, can withstand attacks on components, or have mechanisms or policies for security.
+In this paper, we take on the challenge of building the first trustworthy automated insulin delivery system, called GlucOS.
+Our design includes: (1) an architecture where we apply separation principles for isolated and simple components, (2) a novel security mechanism and policy to enable secure ML-driven insulin delivery, (3) the ability to withstand attacks on insulin pump driver software, and (4) the application of formal methods to prove correct our insulin delivery path, the most sensitive part of the system.
+Despite these efforts, we identify scenarios in automated insulin delivery that require users to take action for security.
+Thus, we incorporate humans as part of our defensive strategy, providing them with agency and trust to take the required, corrective actions.
+Our approach is distinguished by its holistic consideration of the entire problem of automated insulin delivery.
+We provide clear security boundaries, mechanisms and policies to mitigate attacks on vulnerable components, and formal methods to provide assurances that our implementation is correct.
 Automated insulin delivery systems are different from traditional applications that the security community focuses on, but they are important.
 Research on operating systems or web browsers can list scores of CVEs from current systems and show how their system improves security by eliminating the impact of some or all of these vulnerabilities by virtue of their design.
-Automated insulin delivery systems do not have a long list of CVEs that we can reference.
+Automated insulin delivery systems do not have a long list of CVEs that we can reference, although there are a few [46, 47, 48, 49].
 However, our position is that we need to make sure that these systems never get CVEs, because if they do, security vulnerabilities can kill people.
 Thus, we focus on building a new system, using security first principles and formal methods, to avoid security vulnerabilities and mitigate their damage from the start.
-Designing and implementing trustworthy automated insulin delivery systems brings with it three main challenges.
-(1) How can automated insulin delivery design decisions improve rather than worsen the security posture of such software for end users?
-(2) How do we apply formal methods to a complex, distributed system that includes a human?
-(3) How can we cope with ML mispredictions, including those from malicious ML?
-In this paper, we take on these challenges and design and implement a new trustworthy automated insulin delivery system, called GlucOS.
-Our design for GlucOS strives to achieve the following four primary goals: to use sound security principles to separate security and safety concerns into critical and non-critical components; to keep our implementation as simple as possible; to prove correct the most critical components; and to enable any predictive algorithms, including deep neural network ML, to control the insulin pump securely and safely.
-GlucOS improves on the current state-of-the-art in automated insulin delivery systems by, for the first time, enabling deep neural network ML to control the insulin pump securely and safely.
-Our key insight for security is that over a long enough period of time, all correct algorithms will dose the same amount of insulin.
-For example, on a given day, a person will need to dose a fixed amount of insulin to absorb all of the glucose from the food they eat.
-Since synthetic insulin action is slow, injecting insulin early by predicting future metabolic states is critical for long-term health.
-GlucOS’s support for modern ML-based algorithms provides users with this predictive power to dose insulin.
-For security, GlucOS pairs a predictive ML model with a conservative and safe model to provide the ML with enough flexibility to control the insulin pump proactively while staying within the bounds dictated by the safe model.
-Our architecture and system are a stark contrast to current systems.
-Current systems are complex, their architecture provides little or no isolation boundaries, they are unable to withstand attacks on their system, and they are unable to use deep neural network ML securely and safely.
-Given that mistakes with insulin delivery can kill people, our goal is to have a fully-featured automated insulin delivery system with security and safety as the primary design considerations.
-Even if parts of the system are malicious, GlucOS continues to operate and ensure that the individual using the system can manage their T1D safely.
-We report our experiences running GlucOS on one individual for six months to manage their T1D.
-Running our software on a real human forces us to design a practical and real system.
-Our evaluation using virtual humans in a simulator shows that our security and safety mechanisms generalize beyond the specific individual.
+Our main contribution lies in a new end-to-end system that we built and the lessons learned from our experience designing, implementing, and deploying GlucOS to humans to manage their T1D.
 Our novel contributions include:
-  • The clean-slate design and implementation of a simple yet effective automated insulin delivery system.
   • Security mechanisms and policies that protect individuals from malicious ML, vulnerable pump drivers, and drastic changes in the human’s glycemic response.
   • The application of formal methods to prove correct the most critical parts of GlucOS.
-  • A case study describing our experiences from a real-world deployment, and results from simulation to show that our techniques generalize.
-We give top priority to the health, safety, and ethics of the human using GlucOS to manage their T1D. Section 11.1 summarizes our safety protocol and Appendix A provides a detailed description.
-2 Background
-Type 1 Diabetes.
-T1D is an autoimmune condition in which the immune system attacks and destroys the insulin-making cells in the pancreas.
-Insulin is a hormone that enables the absorption of glucose (sugar) from the bloodstream into the brain, the muscles, and other organs and tissues where it is used as an energy source.
-When people eat, the digestive system converts food into glucose, which makes its way into the bloodstream, but it cannot then be used by the body without the action of insulin.
-Since people with T1D are unable to produce their own insulin, they must inject synthetic insulin to replace this key function, and the appropriate dosing of this insulin is both complicated and labor-intensive.
-For example, insulin doses must be closely correlated with food intake, but the absorption of exogenous insulin is slower than food digestion, so people with T1D must precisely predict their future food intake in order to inject the right amount of insulin prior to each meal.
-Once injected, typical insulin formulations will affect glucose metabolism for up to three hours, and during that time will interact with other factors like exercise, sleep, stress, caffeine, and alcohol.
-However, the exact effects of these factors on an individual’s glucose levels are even harder to predict than those of food alone, resulting in the need for continuous adjustments by the individual in order to maintain a balanced glycemic state.
-One solution that has emerged to address the complications of daily T1D management is automated insulin delivery systems, which connect a CGM sensor and insulin pump to enable automatic, individualized adjustments in insulin delivery based on glucose trends [14, 18, 40, 48, 17, 7, 4].
-The goal of T1D management is to maintain glucose levels that are in a safe, target range (70-180 mg/dl) for the majority of the time – an outcome that has been linked to positive long-term health outcomes.
-Glucose levels 54-69 mg/dl (level 1 hypoglycemia) or less than 54 mg/dl (level 2 hypoglycemia) put individuals at risk for acute complications like impaired consciousness and seizure, while levels 181-250 mg/dl (level 1 hyperglycemia) or greater than 250 mg/dl (level 2 hyperglycemia) increase risks for long-term vascular damage.
-The main physiological concept in this paper is a goal of maintaining glucose levels that are in the target range the majority of the time.
-Automated insulin delivery systems.
-Automated insulin delivery systems connect a CGM device – which measures interstitial glucose concentration every 1-5 minutes – with an insulin pump, which delivers a continuous subcutaneous insulin infusion.
-Using a variety of algorithms, these systems vary insulin delivery in response to CGM data, reducing the amount of human input and cognitive load required to achieve glycemic stability (Figure 1).
-Several commercial [17, 7, 4] and open source [40, 48] automated insulin delivery systems exist today.
+  • A biological invariant that combines end-to-end physiological security checks with formal methods.
+  • A case study describing our experiences from a real-world deployment with seven individuals, and results from simulation to show that our techniques generalize.
+  • A human-centric design that considers the user as an integral part of the system, influencing both the overall architecture and including them in our defensive strategy.
+We give top priority to the health, safety, and ethics of the humans using GlucOS to manage their T1D. Section 9.1 summarizes our safety protocol and Appendix A provides a detailed description.
+There are 8.4 million people living with T1D worldwide [30], underscoring the need for providing trustworthy systems for automated insulin delivery.
+Our work, carried out with rigorous medical and ethical oversight, is a first step in addressing this need.
 2 Overview
-This paper describes our design for GlucOS, a system for trustworthy automated insulin delivery.
-In this section, we first provide an overview of GlucOS’s system architecture to describe the key components involved in automated insulin delivery.
-While we provide an overview of all components, we specifically focus on those components that are part of our trusted computing base.
-We then define our threat model in terms of our assumptions and goals with GlucOS.
+This paper describes our design for GlucOS, a system for trustworthy automated insulin delivery; we have two primary goals.
+First, we want the software to be simple and correct.
+Second, we want to identify the most vulnerable parts of the system and be able to prevent attacks on these components or withstand successful attacks.
+This section describes our overall architecture for GlucOS that strives to achieve these goals, and our threat model.
+2.1 Locking down the insulin delivery path
+We focus our system architecture on providing security mechanisms and policies on the insulin delivery path.
+The insulin delivery path includes models for calculating the amount of insulin still active in the individual (insulin can take up to six hours to fully absorb), closed-loop algorithms for calculating insulin dosing, deep neural network ML to predict future metabolic states, and drivers for programming the insulin pump.
+Together, these components are both the most complex, and the most sensitive from a security perspective because they are responsible for insulin delivery.
 2.2 Architecture
-From a system design perspective, our primary goals are to keep our implementation simple and enable our use of formal methods to prove correct the most critical parts.
+From a system design perspective, we strive to keep our implementation simple and enable our use of formal methods to prove correct the most critical parts.
 For our software system architecture, we use separation principles from the OS and microkernel areas [8, 33] applied to the application layer to push complexity away from the most critical parts of the system.
@@
 By decomposing the overall automated insulin delivery system, we strive to keep our trusted computing base simple.
-Also, we want flexibility so that people can run whatever closed-loop algorithms they want to and use whatever pump hardware they want – we focus on providing mechanisms, policies, and systems for security and safety.
 Figure 2: Overall architecture for GlucOS.
@@
 The BioKernel is the component that interacts with the CGM and insulin pump hardware, executes the closed-loop dosing algorithm, runs security and safety checks, and produces event logs that other components use to learn the state of the system.
-The BioKernel runs the closed-loop algorithm every five minutes as new CGM readings come in.
+The BioKernel runs the closed-loop algorithm every five minutes as the new CGM readings come in.
 Once the BioKernel receives new CGM readings, it collects the current metabolic state of the individual and sends it to a predictive ML model.
@@
 This reference monitor architecture ensures that we can withstand attacks from malicious predictive ML as well as malicious pump drivers without having to know the internals of how they operate.
-For predictive ML, our reference monitor enforces algorithmic security, where we bound the amount of insulin calculated by ML to stay within dynamic safety bounds computed by a reactive safe physiological model.
+For predictive ML, our reference monitor enforces algorithmic security, where we bound the amount of insulin calculated by ML to stay within dynamic safety bounds computed by a simple reactive safe model.
 For the pump driver, we interpose on the I/O path to ensure that only the pump commands issued by the BioKernel are programmed into the pump.
@@
 2.3 Threat model
-We consider attacks from malicious predictive ML algorithms as well as malicious pump drivers.
-We assume that malicious ML can intentionally miscalculate the amount of insulin to inject and malicious pump drivers can issue arbitrary pump I/O operations to add, drop, or modify commands sent to the pump.
-We also consider real-world safety concerns as part of our threat model.
-ML mispredictions are part of these safety concerns since even well-designed ML will sometimes make poor decisions.
-Safety concerns also include pump communication errors emerging from bugs in the pump driver, pump mechanical failures [55], and pump site skin inflammation or infection [36], all of which result in a mismatch between the amount of insulin calculated in closed loop and the actual amount of insulin active in the human.
-Our goal with GlucOS is to not merely detect attacks and safety violations.
-Our goal is to build a secure system for automated insulin delivery that can withstand active attacks and real-world safety concerns.
-We assume that CGM sensors are correct and operate within their error bounds.
-If the CGM sensors stop producing measurements, our system will be unable to handle it.
-We also assume that the Swift type system is correct and that the underlying operating system, the CGM driver and hardware, and the pump hardware we use are free of attacks and their device communications maintain integrity.
-Thus, we do not consider attacks emerging from tampered CGM sensors or pump hardware (such as adversarial attacks against predictive ML) in our threat model.
+The most serious health risk for people who use synthetic insulin is low glucose, or hypoglycemia.
+Hypoglycemia results from an overdose of insulin.
+Once an automated insulin delivery system has delivered insulin, there is no way to remove it from the body.
+If untreated, severe hypoglycemia can lead to loss of consciousness, seizure, or death in a matter of hours.
+Thus, our primary security objective is to reduce or eliminate hypoglycemia that results from an attack on the system.
+Quantitatively, our goal is to maintain the American Diabetes Association’s recommended time in hypoglycemia of less than 4% [12].
+As our design philosophy centers around humans, providing flexibility for people to use whatever ML they want by loading personalized models or using any pump hardware they want is important.
+This flexibility helps guide our focus on providing novel security mechanisms and policies for these components.
+Our goal is to make sure that they stay safe even if the model they pick is malicious or the pump driver they choose has security vulnerabilities.
+Attackers may target automated insulin delivery systems because they can directly harm the human using the system.
+We assume that the Swift type system is correct and that the underlying operating system, the CGM driver and hardware, and the pump hardware we use are free of attacks and their device communications maintain integrity.
+Thus, we do not consider attacks emerging from tampered CGM sensors or pump hardware in our threat model.
 These are important but complementary problems to our work, which focuses on the system architecture and algorithms for automated insulin delivery.
+We consider attacks from malicious predictive ML algorithms.
+These attacks can come from trojan attack logic embedded directly in a model, or from an adversarial attack on predictive ML.
+For adversarial attacks, although we assume that the CGM and insulin pump remain attack free, an adversary can manipulate these entities by serving an individual who is using our system with calculated amounts of carbohydrates in their food (e.g., at a restaurant), thus influencing the resulting CGM readings and corresponding insulin commands.
+In our threat model, malicious predictive ML can issue arbitrary insulin doses.
+We also consider malicious pump drivers. In our threat model, malicious pump drivers can issue arbitrary pump I/O operations to add, drop, or modify commands sent to the pump.
+Successful attacks on predictive ML or pump drivers, if unchecked, can be devastating because they directly affect insulin delivery and can be used by an attacker to induce potentially life-threatening hypoglycemia.
 3 Algorithmic security: Withstanding malicious ML
-None of the current automated insulin delivery systems use deep neural network ML.
-Although they do use ML in general, they use statistical methods, like linear regression or model predictive control [19, 60, 42, 52, 48], to learn physiological parameters.
-The reason for using this more traditional form of ML is sound: it provides determinism and has a physiological basis that people can use to vet its decisions.
+Currently, no automated insulin delivery systems use deep neural network ML.
+Some current systems use statistical methods, like linear regression or model predictive control [21, 66, 43, 58, 54].
+The reason for using this more traditional form of ML is sound.
+It provides consistent, explainable results and has a physiological basis that people can use to vet its decisions.
 However, all of the recent advances in ML, like image classifiers and large language models, have come from deep neural networks, which researchers have shown also work well for predicting metabolic states for use in automated insulin delivery systems [73, 35, 70, 65, 74, 64].
-Despite this promise, none of the current automated insulin delivery systems use more advanced deep neural networks due to the black-box nature of deep neural networks and the potentially dire consequences of mispredictions – an automated insulin delivery system’s unchecked “hallucination” [1] could be lethal if it delivers an inappropriate insulin dose [20, 13].
-Modern ML models make mistakes.
-High-profile examples include Google Photos classifying black people as gorillas [2], LinkedIn recruiting ML learning gender bias [5], Microsoft’s chatbot, Tay, trained on Twitter data (what could go wrong) becoming racist [3], and ChatGPT large language model hallucinations [1].
-In all of these examples, professional ML engineers with plenty of resources produced well-designed ML systems that hit a gray area in their training data or decision space and made bad decisions.
-All automated insulin delivery systems that use ML need to have a mechanism to protect against malicious ML or benign ML that makes inevitable bad decisions.
-In this section, we discuss the design and implementation of GlucOS’s algorithmic security and safety mechanisms.
+Despite this promise, none of the current automated insulin delivery systems incorporate more advanced deep neural networks due to the black-box nature of deep neural networks and the potentially dire consequences of mispredictions – an automated insulin delivery system’s unchecked “hallucination” [1] could have lethal consequences if it delivers an inappropriate insulin dose [22, 13].
+In this section, we discuss the design and implementation of GlucOS’s algorithmic security mechanisms to enable predictive models, including deep neural network ML, to run safely in automated insulin delivery systems.
 3.1 Mechanism fundamentals
-Our key insight that underpins our security mechanism design is that all correct insulin delivery algorithms will deliver the same amount of insulin over a long enough period of time.
+Our key novel insight underpinning our security mechanism design is that all correct insulin delivery algorithms will deliver the same amount of insulin over a sufficiently long period.
+People inject insulin to facilitate glucose absorption after eating.
+Digestion converts food to glucose, and insulin facilitates the absorption of glucose to tissue (e.g., the brain) where the body can use it as an energy source.
+Since insulin dosing is driven by food consumption, all correct algorithms should dose the same amount of insulin to cover glucose from digestion.
 However, the timing of this insulin delivery matters for maintaining balanced glucose levels (Section 8).
-Thus, our algorithmic security mechanism pairs a predictive model that anticipates future metabolic states and doses insulin accordingly with a reactive safe model that measures the current metabolic state and doses insulin based only on what it can measure currently.
-Conceptually, the predictive model is the primary model in our system and should control the pump most of the time, and we use the reactive safe model for accounting so that we can track how far the predictive model deviates from a known-to-be-safe baseline.
+Insulin activation is a slow process, with injected insulin taking up to six hours to fully activate.
+To meet this challenge, our algorithmic security mechanism pairs two models.
+Our first model is a predictive model that anticipates future metabolic states and doses insulin proactively.
+Our second model is a reactive safe model that measures the current metabolic state and doses insulin based only on what it can measure currently.
+This pairing allows us to balance proactive insulin dosing for faster responses with a more conservative approach for safety.
+Architecturally, in a novel approach for automated insulin delivery systems, the predictive model runs outside of our trusted computing base while the reactive safe model is part of the trusted computing base.
+This separation enables individuals to use whatever predictive models they want, while maintaining security.
+It also defends against malicious predictive models, as all insulin delivery decisions are ultimately vetted by the reactive safe model.
+This design also enables people to update their predictive models without violating core safety guarentees.
+Conceptually, the predictive model is the primary model in our system and should control the pump most of the time.
+We use the reactive safe model for accounting so that we can track how far the predictive model deviates from a known-to-be-safe baseline.
 Then, we bound the size of this deviation.
-By combining these two models, we strive to get both the safety of a reactive safe model that operates only on facts (Section 4.2) and the tight control we expect from an ML-based predictive model (Section 4.3).
+By combining these two models, we strive to get beneficial properties from both.
+Our approach combines the safety of a reactive safe model and the performance from a predictive model.
+Our security mechanism draws inspiration from previous work on combining simple and complex models for safety [61, 51].
+However, our approach extends these ideas by providing a novel security mechanism specifically designed for insulin delivery models.
+Our key contributions lie in our insight into the security properties of insulin delivery algorithms and our principles for how to apply our mechanism to automated insulin delivery systems (Section 3.2).
 Figure 3: Algorithmic security mechanism for GlucOS.
@@
 The BioKernel will collect both temporary basal commands and look back at the differences between the reactive safe model and predictive ML model for the past three hours to see how much insulin delivery it should attribute to the predictive ML model.
-If the predictive ML model is still within its insulin delivery limits, it uses the predictive ML model’s temporary basal command to program the pump.
+If the predictive ML model is still within its safe insulin delivery limits, it uses the predictive ML model’s temporary basal command to program the pump.
 If it has exhausted its limits, we clamp it to fit within the insulin bounds we set relative to the reactive safe model.
-For more details about how this algorithm works, please see our implementation in GitHub (link omitted).
-4.2 Reactive safe model
-Since we use our reactive safe model to anchor our security policy, and it runs within our trusted computing base, our goals for this model are to have it be safe, easy to understand, and have a simple implementation.
-To make this model safe, we base its calculations solely on facts observed automatically from the CGM and the insulin pump.
+To set bounds, we base our calculations on human physiology.
+People produce a background level of glucose for energy.
+People living with T1D need to inject insulin to facilitate the absorption of this background glucose.
+Our method for setting bounds allows the predictive model to use up to three hours worth of background insulin in advance.
+If the prediction is correct and glucose levels rise (e.g., from eating), the proactive insulin will manage it.
+If the prediction is wrong and no extra glucose appears, the body’s ongoing glucose production will help use the extra insulin over the next three hours.
+In the worst case, an incorrect prediction might lead to low glucose.
+However, with our mechanism in place, we limit the impact by our three-hour bound and offset by the body’s continuous glucose production.
+3.2 Model principles
+Our contribution is not in the specific algorithms we use, but rather in the insight around insulin delivery being equal for all correct algorithms and how we use this insight for security.
+In this section, we outline the algorithms we use in our current implementation and define the principles for how to apply security to any predictive insulin dosing algorithms.
+Algorithmically, we use a PID controller [25] from feedback control for our reactive safe model, and we use a deep neural-network from the literature for meal predictions for our predictive ML model [52].
+Although we use this model in our current implementation, our system supports any predictive model.
+PID controllers are used in automated insulin delivery systems [68, 38] and research predictive ML models for insulin delivery have been well covered in the literature [73, 35, 70, 65, 74, 64].
+The design of the reactive safe model is crucial for algorithmic security, as it serves as the anchor for the security mechanism.
+The ideal reactive safe model should be safe, easy to understand, and have a formally verified implementation (Section 7).
+To ensure the safety of this model, we base its calculations solely on facts observed automatically from the CGM and the insulin pump.
 Given these high-quality sources of data, and a simple insulin absorption physiological model, we can calculate how much insulin an individual needs and how much insulin they have injected but have not yet activated.
@@
 This basic formulation is a standard formulation used in most automated insulin delivery systems.
-One wrinkle we add to this otherwise standard calculation is that we learn the physiological parameters we use in these calculations using a data-driven approach for safety.
-Our learning system consists of a weekly learning task that evaluates three weeks worth of data to infer broader physiological trends.
-These learned settings become the baseline for our parameters.
-Then, we use a PID controller from feedback control [22] to refine these values given recent observations.
-Being able to adapt is important for handling pump malfunction and pump insertion site safety violations (Section 3.2).
-Calculating insulin is a simple calculation to implement, consisting of only 25 lines of formally verified code in the BioKernel (Section 8).
-However, despite this simplicity, it is effective (Sections 10 and 11).
-The main drawback of this model is that is it unable to detect underdosing insulin during food digestion quickly enough to avoid high glucose.
-4.3 Predictive model
-To predict insulin underdosing during digestion, we use a food prediction deep-learning model from the literature [45].
-Our version of this model is a generalization where we model any source of glucose, not just food, but the details of it are beyond the scope of this paper.
-In our evaluation our model outperforms the reactive safe model both in simulation (Section 10) and for the individual that uses our system for their real-world insulin dosing (Section 11), showing why it is important for automated insulin delivery systems to have a practical way to adopt ML.
-Figure 1 shows an example of when the predictive ML model correctly detects digestion and doses insulin to maintain balance.
-This figure shows a screenshot from the BioKernel’s main user interface.
-The individual ate breakfast at 9 am and their glucose levels rose.
-The predictive ML model correctly detects that the individual is going to absorb additional glucose from digestion, resulting in a predicted glucose level rise over the next hour.
-Accordingly, the predictive ML model injects enough insulin to cover the rising glucose from the meal and keep the individual’s glucose levels in range.
+In contrast, predictive models have no restrictions since their outputs are constrained by the reactive safe model and our security mechanism.
+The ideal predictive model anticipates future metabolic states and adjusts insulin proactively to avoid glycemic imbalance.
+In our evaluation, our predictive model outperforms the reactive safe model both in simulation (Section 8) and for the individual that uses our system for their real-world insulin dosing (Section 9), showing why it is important for automated insulin delivery systems to have a practical way to adopt ML.
 4 Driver security: Withstanding malicious pump drivers
 Pump drivers provide abstractions for insulin delivery applications to communicate with insulin pumps.
-Much like traditional device drivers, they have complex implementations since they help accomplish a wide range of tasks including pump setup/teardown, managing pump UX, handling insulin delivery/suspension, managing pump errors, accurately tracking insulin doses etc.
-Different pump drivers use different implementations to carry out these tasks.
+Much like traditional device drivers, they have complex implementations since they help accomplish a wide range of tasks including pump setup/teardown, managing pump UX, handling insulin delivery/suspension, managing pump errors, accurately tracking insulin doses, and so on.
+Different pump drivers use different implementations with varying amounts of customizations and features to carry out these tasks.
 While CVEs do not currently exist for pump drivers, there is strong possibility for bugs or vulnerabilities in their implementation given their complexity and the fact that they not been formally verified.
-Since drivers control insulin pumps, these vulnerabilities can even lead to the death of individuals.
-For example, consider a scenario where the user decides to inject 3 U of insulin using the BioKernel.
+Since drivers control insulin pumps, these vulnerabilities can harm people using them, even lead to the death.
+For example, consider a scenario where the user decides to inject 3 units (3U) of insulin using the BioKernel.
 The BioKernel would relay this as a command to the pump through abstractions provided by the pump driver.
-If the pump driver inadvertently or intentionally issues this command twice, 6 U of insulin would be delivered to the user, which would put the user in danger by drastically reducing the amount of glucose in their body.
-With GlucOS, our goal is to support their complexity by allowing users to employ pump drivers of their choice while upholding security.
-Our security goals are to not just protect users from inadvertent bugs in pump drivers, but to also protect them from malicious pump drivers which could intentionally add, modify or drop pump commands.
+If the pump driver inadvertently or intentionally issues this command twice, 6U of insulin would be delivered to the user, which would put the user in danger by dosing too much insulin.
+We need a mechanism to verify that the pump driver is executing correctly.
+Our security goals for GlucOS’s driver security are to protect users from inadvertent bugs in pump drivers and safeguard against potentially malicious pump drivers.
+To achieve these goals, we (1) introduce driver security logic in the BioKernel that interposes on pump driver-pump communications, (2) validate all commands against known high-level pump commands, and (3) detect and handle anomalies by stopping any added/modified commands and proactively handling dropping critical commands.
+This approach allows people to choose between various pump drivers while maintaining security, balancing user flexibility and safety with automated insulin delivery.
+It also keeps our trusted computing base small, simplifying formal verification.
+4.1 Driver security logic design
 Figure 4: We interpose on communications between the pump driver and the pump to handle untrusted pump drivers to deliver insulin.
-5.1 Reference monitor overview
-We introduce a reference monitor as part of the BioKernel to handle untrusted pump drivers, which interposes on communications between the pump driver and the pump.
-We choose to vet these communications instead of implementing our own pump driver to give users the ability to use any pump driver of their choice.
-This also helps keep our trusted computing base small thereby easing formal verification.
-Figure 4 provides an overview of our reference monitor.
-We only validate commands pertaining to insulin delivery since they pose the most dangerous threats to the user.
-From a high-level, the pump registers insulin commands with the reference monitor before issuing them to the pump driver.
-In the reference monitor, we check to ensure that the command issued by the pump driver is consistent with the command registered by the BioKernel.
-At the reference monitor, we also intercept all records of delivered insulin doses to ensure their integrity for closed loop calculations.
-On detecting manipulation from the pump driver, we transition insulin delivery to a default safe state (described in Section 7) where there is no immediate threat to the user.
-We do not validate communication pertaining to pump setup / UX since users can directly detect the respective manipulations.
-Figure 5 lists the different types of messages pertaining to insulin delivery exchanged between a pump driver and an insulin pump.
-We list other message types in Appendix B.
-5.2 Design Principles
-Avoid programming the pump. We avoid issuing insulin delivery commands to keep our implementation simple.
-Ensure that the user always stays safe.
-Delay in delivering insulin does not pose any immediate threat to the user.
-However, delay in cancelling a wrongly issued insulin delivery command (such as one with a high insulin dose) can potentially kill the user.
-Thus, in the interest of keeping users safe, we violate the principle of not issuing commands to the pump in two specific cases of cancelling insulin delivery.
-Disregard non-critical functionality in the interest of security.
-Some pump drivers incorporate an automatic retry mechanism when they’re unable to communicate with the pump.
-Such retry mechanisms complicate validation, since it becomes difficult to distinguish between duplicate commands issued as part of a retry versus duplicate commands issued as the result of a bug or an attack.
-We, thus, disregard such functionality by treating all duplicate commands as bugs or attacks.
 Message Message
@@
 Figure 5: Types of insulin related messages exchanged between between a pump driver and an insulin pump.
-5.3 Reference monitor design and implementation
-The BioKernel invokes methods exposed by the reference monitor to register commands to be sent to the pump before invoking the relevant methods in the pump driver.
-The BioKernel also registers two callbacks with the reference monitor, one to be invoked on successfully receiving a response from the pump and the other in case of error.
-The pump driver then updates its internal states, encodes the command and its relevant parameters and sends it to the reference monitor.
-The reference monitor then checks that the pump driver has not added new commands or modified the parameters before encrypting and sending them to the pump.
-If the reference monitor detects added or modified commands, it transitions the pump to operate in a default safe state and rejects all further commands from the driver.
-We employ a bounded delay of one second to detect dropped commands.
-Concretely, the reference monitor detects dropped commands if it does not receive anything from the pump driver within one second of registering a command from the BioKernel.
-It treats dropped commands the same way as added or modified commands by transitioning the pump to a default safe state and rejecting all further commands.
-We transition the user out of this safe state and start accepting commands only after the user manually configures the BioKernel to trust the driver again.
-The reference monitor operates differently on detecting dropped commands that deal with cancelling insulin delivery.
-Users will seek to cancel insulin delivery if they accidentally triggered an insulin delivery command or incorrectly set its parameters.
-Since excess insulin injection is dangerous, the reference monitor would automatically send these commands to the pump before transitioning the pump to the default state and rejecting all further commands.
-There are two commands that deal with cancelling insulin delivery: cancelDelivery and setting the suspend option on setTempBasal.
-cancelDelivery seeks to cancel a previous insulin delivery command and (if the delivery is still ongoing) and the suspend option on setTempBasal seeks to suspend all insulin delivery until the user resumes delivery.
+Figure 4 provides an overview of our driver security logic.
+The BioKernel registers high-level commands, like “deliver 2U of insulin” with our driver security logic, and sends them to the pump driver.
+The pump driver converts these commands into I/O messages to program the pump.
+However, these I/O messages come back to the driver security logic, which checks to make sure that they are consistent with the previously registered high-level pump command, before forwarding them to the pump hardware.
+We only validate commands for insulin delivery since they pose the biggest threats to the user.
+We do not validate communication pertaining to pump setup / UX since users can directly detect the respective manipulations.
+Figure 5 lists the different types of messages pertaining to insulin delivery exchanged between a pump driver and the OmniBLE insulin pump, which we use in our current implementation.
+We list other message types in Appendix B.
+By interposing on the pump I/O message path, the driver security logic can ensure that the driver has not added, modified, or dropped any pump commands.
+Detecting added or modified commands is relatively straightforward.
+Our driver security logic in the BioKernel knows all high-level pump commands that it sends to the driver.
+Thus, we have a simple logic to confirm the corresponding I/O messages from the driver.
+The driver security logic stops any anomalous I/O messages before sending them to the pump.
+Detecting dropped commands is more nuanced when the command that the driver drops is a command to cancel an ongoing insulin delivery.
+Users can cancel insulin delivery if they accidentally triggered an insulin delivery command or incorrectly set its parameters.
+Since excess insulin delivery is dangerous, the driver security logic automatically sends these cancel commands to the pump when it detects a dropped insulin cancellation command.
+There are two commands that deal with canceling insulin delivery: cancelDelivery and setting the suspend option on setTempBasal.
+The cancelDelivery command cancels a previous insulin delivery command (if the delivery is still ongoing) and the suspend option on the setTempBasal command suspends all insulin delivery until the user resumes delivery.
+The most interesting lesson learned from our experience is in coping with dropped pump commands.
+In our original design, our goal was to ensure that our driver security logic only understood I/O commands and prevented any malicious I/O commands from going through.
+However, our driver security logic ended up needing to issue its own I/O commands to deal with dropped insulin delivery cancellation commands.
+This more proactive approach to driver security is different than previous approaches that also use specifications for runtime driver verification [69, 63, 44], which simply infer I/O semantics and do not issue their own I/O commands.
+In general, we avoid programming the pump directly from the driver security logic.
+But in this case, we decided to program the pump directly from the driver security logic for human safety.
+When we detect a malicious driver, we transition the system into a safe manual mode where the individual needs to take over control of their own insulin delivery.
+They can still use the BioKernel to manage their insulin, but they would be unable to use the automated closed-loop algorithms.
+In our evaluation, our system had no false positives and detected all attacks (Section 8.2).
+See Appendix C for more details on how we maintain safety even when detecting an attack.
+4.2 Driver security logic implementation
 We focus our implementation on Loop’s OmniBLE [4] pump driver.
@@
 To remove ambiguity in terms of whether the cancelDelivery command was issued as part of setTempBasal, we enforce OmniBLE to always issue the cancelDelivery command.
-We evaluate GlucOS in terms of driver security in Section 10.4.
-Overall, we employ straight-forward checks to detect manipulations from buggy/malicious pump drivers.
-However, the reference monitor’s design is more nuanced to handle the full range on insulin commands, missing out on any of which could potentially kill the user.
 5 End-to-end security: Biological invariant
-A key novelty in our overall approach is our introduction of a biological invariant that provides end-to-end verification.
-End-to-end verification in this context means that we verify the BioKernel’s glucose calculations against observations of the user’s metabolic state.
+A key novelty in our overall approach is our introduction of an end-to-end check that ensures the BioKernel’s glucose calculations against observations of the user’s metabolic state.
+This raises an interesting challenge for formal verification as it requires modeling the true biological state of the user – a parameter that is both unknown at runtime and distinct from measured parameters which are part of the software implementation.
+We refer to the true safety invariant we enforce on the user’s state as the biological invariant, and proving it is both a key part of our implementation and one of the key conclusions of our formal verification efforts.
 If there is a large mismatch between our calculations and what we observe, then it means that our calculations are wrong.
-These deviations can happen from drastic changes in the user’s glucose absorption, like during exceptionally vigorous exercise or if the pump insertion site hits a vein and delivers insulin directly into the blood stream.
-Regardless of the cause, the biological invariant detects these conditions and prevents harm to the user.
-The biological invariant closes the gap between verified software and human metabolism by incorporating a direct check against the user’s physiological data into the formal verification process.
-The biological invariant provides a runtime mechanism to detect when GlucOS’s calculations have become unsound.
-Specifically, the biological invariant compares the BioKernel’s theoretical calculation of the user’s blood glucose level against the actual observed blood glucose readings.
-Theoretical glucose levels provide an estimate of how the system thinks a given insulin dose will influence glucose levels.
-GlucOS uses its estimate of the human’s insulin sensitivity when calculating these doses.
-A significant divergence between the theoretical and observed values indicate a potential catastrophic failure mode, where GlucOS’s estimates of the user’s insulin sensitivity and calculations have diverged from the user’s real physiological state.
-The biological invariant allows the system to safely transition into a default safe state (Section 7).
-If the divergence between theoretical and observed glucose exceeds a threshold, indicating a loss of confidence in the model’s calculations, GlucOS discontinues automated insulin delivery and enters the safe state.
-The biological invariant primarily focuses on resolving the undesirable situation where GlucOS believes that the person is less sensitive to insulin than they really are.
-In this case, the delivered insulin will remove more glucose than required, leading to potentially fatal low glucose levels.
-Consider the example of microbolusing when the user is above recommended glucose levels.
-If the system assumes a state where it believes the user is less sensitive to insulin than they really are and the user is 2x more sensitive, it will inject 2U of insulin when only 1U is required.
-While the system believes this will restore the user to safe levels, the actual glucose suppression will now push the user towards dangerously low glucose levels.
-Due to the higher clinical risk, we emphasize triggering the invariant in the undercalculation of insulin sensitivity case to transition to the default safe state.
-We set a threshold of 30 mg/dL for the maximum allowable divergence as shown in Figure 6 based on clinical evaluations showing detectability of significant glycemic deviations above this level[37, 32].
-Figure 6: This figure shows an example of diverging differences between theoretical glucose and the observed glucose levels for a user.
-GlucOS will lose confidence in its calculationsand tranisition to a default safe state once it goes beyond a fixed difference threshold, where the biological invariant fails.
-We formally model and verify the biological invariant using the Dafny language. The core aspects we specify include:
-(a) A ghost variable representing the ground-truth physiological blood glucose level of the user.
-This models the actual, unobserved glucose state inside the human body that the system is trying to estimate and control.
-(b) The theoretical calculation of the blood glucose level according to the system’s model.
-This captures the glucose prediction made by the system based on its assumptions about insulin dynamics, metabolism etc.
-(c) The observed value of the blood glucose level from glucose readings from the user.
-(d) The invariant property itself, which asserts that the absolute difference between the theoretical model glucose and the ground-truth physiological glucose cannot exceed a threshold of 30 mg/dL.
-Our primary goal of verification in this segment is to make sure that there are no execution sequences that can violate the invariant without reaching a state where automated insulin delivery is disabled.
-This verification also ensures the biological invariant correctly captures our intention - if the system’s theoretical model deviates too far from the real physiological state, automated dosing is stopped to prevent potential over/under delivery of insulin.
-We also verify framing properties, showing the invariant’s state updates are precise and do not inadvertently modify other system variables.
-The verified Dafny specifications are then translated to Swift to integrate the biological invariant into the runtime system implementation for execution, with high assurance from the Dafny verification that they correctly implement the safe-guarding behavior in correspondence with the proven formal model.
-Incorporating the biological invariant enables fundamental safety assurances not possible with traditional software verification alone.
-First, it constrains the set of possible erroneous system behaviors by restricting manual divergence between the theoretical model and actual physiological state.
-Section 10.5 shows that the biological invariant improves the system’s time-in-range glycemic control significantly compared to baseline when the system’s model and knowledge of user parameters becomes unsound.
-Second, it provides higher confidence of overall system reliability by directly incorporating physiological cross-checks into the verification process.
-We consider the biological invariant an important conceptual contribution that can extend the formal verification process to dynamically check and respond to deviations at runtime forming the complete bridge between the software and human physiology.
+These deviations can happen from drastic changes in the user’s glucose absorption, like during exercise when the body uses glucose without needing insulin [14], or if the pump insertion site hits a vein and delivers insulin directly into the blood stream [62].
+Regardless of the cause, the end-to-end check detects these conditions and prevents harm to the user.
+For the purposes of the following example, suppose the end-to-end check is that the measured glucose levels are always within 30mg/dl of the predicted level; we refer to this as the implementation invariant that is enforced in an end-to-end way across the system:
+$|g_{\mathsf{measured}}-g_{\mathsf{predicted}}|<30\mathsf{mg/dl}\quad(1)$
+Naively, one might look at the $30\mathsf{mg/dl}$ on the right-hand side and think that the user’s true glucose levels are within $30\mathsf{mg}$ of the actual glucose levels.
+But this would be wrong!
+The problem is that the measured glucose level is subject to measurement error; the state-of-the-art Dexcom G7 CGM has a mean absolute relative difference of 8.2% from the true glucose level in the body [29].
+For the sake of example, suppose that the measurement error on the CGM reading is around 5%; we refer to this as the CGM invariant:
+$|g_{\mathsf{measured}}-g_{\mathsf{actual}}|<5\mathsf{mg/dl}\quad(2)$
+and from this we can derive the true conclusion on the user’s glucose levels, which is what we refer to as the biological invariant.
+We stress that this invariant involves the parameter $g_{\mathsf{actual}}$ that is not available in software; it instead ensures a property (which we are able to formally prove) about the real level of glucose in the biological system.
+That statement we prove in the software is actually the following, not (1):
+$|g_{\mathsf{measured}}-g_{\mathsf{predicted}}|<35\mathsf{mg/dl}\quad(3).$
+When we run BioKernel on a real human, we don’t want to assume safety with respect to the measured glucose levels (invariant (1)); we actually want to ensure safety with respect to the real glucose levels in the human that is using the system (invariant (3)).
+Failing to distinguish these cases could be a severe software bug, and would jeopardize the end-to-end safety that we aim to provide by failing to account for CGM error.
+In the rest of the section, we describe the invariants that we enforce in the implementation (including invariant (1) as a simple case) and what additional invariants we prove in our formal verification development (including invariant (3)) below.
+Implementation invariant.
+In the implementation, we check that changes in glucose levels that we observe from CGM readings and the individual’s insulin delivery history are consistent with our theoretical calculations of how much the person’s glucose should change given these same inputs.
+If their glucose is dropping by more than our theoretical calculations show, then they are more sensitive to insulin then our system believes they should be.
+We set a threshold of 30 mg/dl per hour, for the maximum allowable divergence based on clinical evaluations showing detectability of significant glycemic deviations above this level [37, 34].
+We take a three-tiered approach to dealing with deviations in our glucose change calculations.
+First, for small differences of less than 30 mg/dl, we accumulate errors and adjust the target glucose level to compensate, clamping the adjustment to 10 mg/dl.
+For example, if the individual’s insulin sensitivity is elevated and their target glucose level is 90 mg/dl, GlucOS will update their target to 100 mg/dl dynamically to adjust.
+Second, when the differences are 30 mg/dl or greater, we shut off insulin delivery.
+Our rationale is that the individual is more sensitive to insulin than our models believe they should be, thus we want to avoid overdosing.
+If their difference goes back to within the expected range, normal execution resumes.
+Third, if they remain at 30 mg/dl or more for more than two hours, we alert the individual and transition the system into a manual mode that requires the user to dose insulin manually.
+Verification.
+In the verification development (described in Section 7), we prove that the biological invariant holds, given error bounds set by the CGM manufacturer.
+First, in the ported Dafny code, we locate the portions which enforce implementation invariants described above, including invariant (1).
+The code includes the two parameters $g_{\mathsf{measured}}$ and $g_{\mathsf{actual}}$, as these are available to the software implementation.
+Second, we model actual glucose by putting error bounds on our CGM readings, in the CGM invariant (2).
+This CGM invariant is inserted as a precondition on all functions that operate on CGM values.
+Third, we track the individual’s actual glucose level using a ghost variable in Dafny.
+Ghost variables are variables that are only available to the prover, these do not exist in the real software system.
+Fourth and lastly, we express as a postcondition the biological invariant (4), which asserts that the absolute difference between the theoretical model glucose and the ground-truth actual glucose cannot exceed a threshold of 35 mg/dl per hour.
+When the code is compiled, the Dafny verifier proves that the biological invariant holds on all inputs to the verified function.
+Evaluation results.
+Our evaluation shows that the biological invariant is effective at keeping humans safe when they face drastic changes in insulin sensitivity (Section 8.3), and that our thresholds of 30 mg/dl per hour and two hours were effective for the human that used GlucOS to manage their T1D without any false positives (Section 9.3).
+Novelty.
+We beleive that having a parameter real biological state of the system is an important modeling decision; and it diverges, for example, from previous work which only models the biological data in simulation [71, 72].
 Traditional software verification techniques like model checking [36] and deductive verification [27] have been explored in the past for hybrid system verification [26] that interact with the human body, but they do not completely account for physiological variability between individuals or model inaccuracies during the insulin delivery process.
-7 Runtime failure management
-Figure 7: GlucOS’s state machine to transition users to the safest manual dosing state when losing trust in insulin delivery (i.e., with a buggy or malicious pump driver) or when the biological invariant is not satisfied.
-In this section, we describe GlucOS’s state machine that combines its security components, i.e., algorithmic security (Section 4), driver security (Section 5) and end-to-end security (Section 6) together to protect users in light of their corresponding failures.
-GlucOS’s state machine consists of three operating states: (1) ML with safety operating in closed loop, (2) reactive safe model operating in closed loop and (3) manual dosing in closed loop.
-ML provides the tightest level of control in managing glucose but poses the largest immediate threat from excess insulin since it preemptively injects insulin by anticipating rise in glucose.
-The reactive safe model provides slightly lesser control, but poses milder immediate threat since it only injects insulin to account for measured excess glucose.
-Manual dosing offers the least amount of control but poses no threat since it only injects the minimum insulin needed for sustenance.
-GlucOS predominantly operates in the state running ML (with safety) in closed loop.
-Certain real-world failures (such as missing CGM readings) transition GlucOS to operate in the state running the reactive safe model in closed loop.
-However, such failures are simple and we discuss ways to handle them in Appendix C.
-In this section, we focus on transitions pertaining to the manual dosing state.
-We transition to the safest manual dosing state when we do not trust insulin delivery (i.e., with a buggy or malicious pump driver).
-We also transition to the safest manual operation when the biological invariant is not satisfied to ensure that we do not risk the possibility of removing too much glucose from our insulin doses.
-If we end up in the manual dosing state, we enforce that only manual user intervention can transition back to other states.
-We visualize these state transitions in Figure 7.
+Also, researchers have used real-world invariants for security in the context of autonomous vehicles [18], but these were based on quad-copter physics not human physiology and did not include formal verification of their implementation.
+6 Last line of defense: Humans
+In this section, we describe our design and implementation for incorporating humans as a critical component of our multi-layered security system.
+While our algorithmic, system-level, and biological invariant defenses form the primary barriers against attacks, humans serve as the last line of defense, capable of mitigating attacks that manifest as glycemic imbalance.
+When an individual’s glucose levels are out-of-range, they need to intervene to bring them back in range.
+Low glucose, or hypoglycemia, occurs from too much insulin, and there is no way to remove insulin that the automated insulin delivery system has already delivered.
+To recover from hypoglycemia, people need to eat sugar.
+Given that people with T1D are already active participants in managing glycemic imbalance, we focus on how we can use this fact to our advantage to create a more comprehensive security system.
+Figure 6: GlucOS predicts hypoglycemia and notifies the individual before it happens so they can proactively prevent it.
+We use predictive yet deterministic mobile notifications to alert people when they are likely to experience hypoglycemia before they reach it (Figure 6), as hypoglycemia presents the most immediate health risk and could be an indicator of an attack.
+Alerts for hypoglycemia are not new.
+All CGM software comes with non-predictive hypoglycemia alerts by default [2] and glucose prediction algorithms have been well covered in the literature [65, 70, 11].
+Our contribution is using predictive alerts before someone reaches hypoglycemia as a security defense.
+In a user study (Section 9.4), we reduced the amount of time six people spent in hypoglycemia as a result of using GlucOS, despite them already having CGM alerts set up prior to our study.
+These results demonstrate a reduction in hypoglycemia for the non-attack case, but the most dangerous successful attacks will result in hypoglycemia.
+Thus, these results suggest that predictive alerting can be an effective security measure.
+Using a simple and deterministic algorithm is important since it runs as a part of our trusted computing base.
+Algorithmically, we predict hypoglycemia by doing a least squares fit for the last 20 minutes of CGM readings and extrapolate forward 15 minutes.
+We trigger alerts if our prediction shows hypoglycemia is likely in the next 15 minutes.
+The most interesting finding from our study was that personalization was important.
+We provided people with the ability to set the prediction threshold and specify how often they wanted the alert to repeat while they remained out-of-range.
+Five of the six participants customized their alerts and the participant noted that the default values we picked coincided with their alerting preferences.
+In our first implementation we did not provide the ability to customize their security settings and we tried to make the usability vs security tradeoff algorithmically and automatically.
+But in the end, simplifying the system and allowing users to personalize their security policy was important for our participants.
+Our finding has the potential to have broader implications for security in healthcare systems.
+Often, security systems favor uniformity and consistency [23, 28], but when managing people’s health we found that allowing people to balance alert fatigue and security increased engagement and potentially increased protection, based on each individual’s needs.
 7 Formal verification
-A single error or vulnerability in an automated insulin delivery system can have life-threatening consequences for the user.
-This criticality motivates our use of formal verification in our development and testing process.
-Through formal verification, we aim to confirm two critical aspects of our insulin delivery system.
-First, we validate the goals and claims behind our critical insulin delivery functions.
-Second, we verify that GlucOS applies the correct state transitions in precarious scenarios to protect the user when the system loses confidence in one or more of its components due to miscalculation or malicious behavior.
-Systematically identifying the invariants and postconditions in alignment with our claims that we verify for a critical function in our system is important.
-Our first step is to clearly define what the function is supposed to do.
-This helps list all the inputs that the function takes, including their types and ranges.
-Then, we define the expected outputs, including their types and valid ranges.
-Determining safety and domain-specific constraints for each function plays a major role in ensuring that the function never produces an unexpected output that puts a user at risk.
-For example, a general constraint would be that a basal rate should never be negative as it is invalid.
-A domain-specific constraint is defining a maximum basal rate threshold that is in accordance with medical guidelines.
-For a full list of the invariants that we define and prove, see Appendix D.
-These steps help define the preconditions, postconditions and invariants that help prove the properties of our verified functions.
-We run formal verification in Dafny to verify our claims on how a function works.
-After we verify a function, we port the code to Swift to make it part of the BioKernel.
-We verify invariants and postconditions for the system’s critical functions in the aforementioned approach.
-For instance, we verify the microbolusing function to ensure a minimum time interval between consecutive micro-boluses to prevent frequent dosing in short intervals (within 4 minutes), confirm that the user glucose level is significantly above the target(atleast 20 mg/dL) before administering a bolus and that the bolus amount is within safe and predefined limits.
-These are claims about our function that we formally verify to assure that they are always correct.
-In another example, we prove correct our insulin delivery calculations to ensure that we maintain accurate accounting of insulin delivery.
-For this verification task, we verify our insulin delivery over a duration of time function to guarantee non-negative results, correctly handle zero duration or non-overlapping time segments, making sure that the dosage does not exceed recommended limits, and accurately calculate insulin delivery based on time segment overlap.
-We perform this type of verification on all the critical functions involved in the insulin delivery process. This high-level of assurance helps us trust the the decisions that GlucOS makes.
-We also formally verify the state machine we use for defining system transitions under undesirable scenarios (Section 7).
-We discover several potential issues in the state transition logic and function design during verification.
-For example, there were cases where the system could transition to an unsafe state under certain conditions, potentially leading to excessive insulin delivery.
-We also identified bugs related to violating valid bounds and domain-specific constraints.
-A more detailed discussion of the bugs have been included as part of subsection  10.6 in our evaluation.
-9 Implementation
-In our implementation, the BioKernel is an iOS app on iPhone or iPad hardware.
-It communicates with a Dexcom G7 CGM and Omnipod DASH insulin pump via Bluetooth low energy and runs the core closed-loop algorithm.
-For the BioKernel, we use the Swift programming language, the Actor abstraction for data-race-free concurrency, and JSON files stored on disk to capture persistent state.
-Although the BioKernel is a fully featured and stand alone automated insulin delivery system, GlucOS uses an event logging system that enables separate apps to implement complementary features while remaining isolated from the BioKernel.
-We implement our event logging system using a Google Cloud App Engine app and the BioKernel produces logs that other apps consume.
-The other apps include an app for alerting when the individual is likely to experience hypoglycemia or hyperglycemia, an app to replay closed-loop runs to verify the results, and an app for running simple ML on data to calculate therapeutic settings.
-To keep predictive ML and the pump driver isolated, we use the Swift type system and object boundaries to maintain isolation.
-Our architecture also supports the use of process boundaries for even strong isolation and to implement OS-level sandboxing.
-iOS lacks support for proper process APIs (fork, exec etc), so we use the Swift type system instead.
-For CGM and insulin pump drivers, we use LoopKit [41] and isolate untrusted portions using the Swift type system.
+To provide additional assurance that our implementaiton is correct, we port our Swift code to Dafny and use Hoare logic.
+As the system we are verifying is complex, our goal is to locate the pieces and target verification in places where formal methods can have the greatest impact; this is a common reality of applying formal methods at scale, see for example [19].
+Since we are porting the Dafny code to Swift, to reduce the chance of human error, we use a line-by-line port as much as possible and add unit tests for each function to provide confidence that the two implementations are equivalent.
+We port updated code back to Swift manually because Dafny does not include code generation for Swift.
+These efforts paid off: we found and fixed nine bugs in our original implementation (Appendix D); in addition to proving various invariants in the code for our insulin delivery path and security mechanisms.
+However, the most interesting part was where these bugs were: all of our core algorithms were correct, but we found bugs in seemingly less critical parts of the system, like user settings, which had subtle interactions with the core algorithms.
+For a full list of the invariants that we define and prove, see Appendix E.
 8 Evaluation
-Given the risks associated with running ML for automated insulin delivery, we answer most questions on 21 virtual humans from an FDA-approved simulator [9].
-The simulator generates CGM readings across different age groups including children, adolescents and adults.
-We also evaluate security mechanisms on virtual patients to see if they can generalize to individuals with different physiologies.
-We present experiences of running GlucOS with a real individual in the next section.
-10.1 Does ML lead to better outcomes for managing T1D?
-We confirm findings from prior research that ML leads to better outcomes [47, 66, 34] by comparing the average proportion of time spent by virtual patients in range when using our reactive safe model (Section 4.2) against our predictive ML model (Section 4.3).
-As mentioned in Section 2, the American Diabetes Association (ADA) recommends that an individual with T1D should spend more than 70% of a day (also called TIR) [8, 21] in order to remain healthy.
-We train separate predictive ML models (feed-forward neural networks trained to minimize MSE loss) for each virtual patient and similarly tune the physiological parameters of the reactive safe model to suit each patient.
-We ensure that we run the reactive safe model and the predictive ML model on identical conditions within the simulator.
-Patient TIR with TIR with
-type ML model reactive safe model
-Adults 93.49% 88.25%
-Adolescents 85.23% 81.98%
-Children 86.6% 77.31%
-Figure 8: Average proportion of time spent in range by 21 virtual patients in identical scenarios with a our predictive ML model and our reactive safe model.
-We show the average proportion of time spent in range by virtual patients from each age group (adults, adolescents and children) in Figure 8.
-We can see that ML helps virtual patients across all age groups spend a larger amount of time in range, thereby showing that ML leads to better outcomes for managing T1D.
-10.2 What are the risks associated with untrusted ML algorithms for managing T1D?
-We assess the impact of predictions made by intentionally malicious ML algorithms on managing T1D.
-Since GlucOS is designed to be flexible with any predictive algorithm, it could be subject to attacks from malicious ML algorithms.
-We intentionally introduce an order of magnitude difference to the amount of insulin predicted by the predictive ML model in Section 10.1.
-In one experiment, we dose one-tenth of the amount of insulin and in another experiment, we dose ten times the amount of insulin.
-We run these malicious ML algorithms within the simulator on each virtual patient on the same scenarios as the previous experiment.
-Figure 9 and Figure 10 show the CGM readings on 3 randomly chosen virtual patients on a particular scenario for the experiments where we dose less and dose more respectively.
-When dosing less, we see that all patients spend most of their time above Level 1 hyperglycemia which would have long-term effects on their well-being.
-When dosing more, all patients fully go below Level 2 hypoglycemia, which would have killed them.
-Figure 9: Variation in CGM readings for 3 virtual patients when running a predictive ML algorithm that intentionally doses one-tenth the amount of the required insulin.
-The CGM readings staying mostly above the Level 1 hyperglycemic range would have drastic effects on well-being.
-Figure 10: Variation in CGM readings for 3 virtual patients when running a predictive ML algorithm that intentionally doses ten times the amount of the required insulin.
-The CGM readings staying below the Level 2 hypoglycemic range would have likely been fatal to all 3 patients.
-Across all virtual patients, we see an average drop in TIR from 88.44% with correct dosing to 23.32% when dosing less and an average drop in TIR from 88.44% to 0% when dosing more.
+8.1 How effective is GlucOS’s algorithmic security mechanism?
+We evaluate GlucOS’s algorithmic security mechanism on 21 virtual humans from an FDA-approved simulator [7]. We use simulation since this evaluation involves injecting dangerous amounts of insulin.
+To be effective, GlucOS’s algorithmic security mechanism should protect users against untrusted ML algorithms that administer incorrect insulin doses (overdose or underdose), while minimally impeding algorithms that are efficient at managing T1D.
+Accordingly, we first run simulations with a predictive ML model inspired from prior research [52], without incorporating our security mechanism to set a baseline of an algorithm that is efficient at managing T1D.
+Then, in the same simulation scenarios, we introduce an order of magnitude difference (increase and decrease) to the insulin doses computed by the baseline algorithm, without incorporating our security mechanism to quantify the risks of untrusted algorithms.
+We then repeat both sets of simulations while incorporating our security mechanism to quantify the impact of the mechanism on correct and incorrect dosing.
+The goal of T1D management is to maintain glucose levels that are in a safe, target range (70-180 mg/dl) for the majority of the time – an outcome that has been linked to positive long-term health outcomes [53].
+Glucose levels 54-69 mg/dl (level 1 hypoglycemia) or less than 54 mg/dl (level 2 hypoglycemia) put individuals at risk for acute complications like impaired consciousness, seizure, and death, while levels 181-250 mg/dl (level 1 hyperglycemia) or greater than 250 mg/dl (level 2 hyperglycemia) increase risks for long-term vascular damage.
+The American Diabetes Association (ADA) recommends that individuals spend at least 70% of their time in range [6, 24] in order to remain healthy.
+8.1.1 Without algorithmic security mechanism
+Predictive ML model baseline With the predictive ML model, we followed all ML best practices in carefully tuning hyperparameters and evaluating the model on simulation scenarios that were different from those used in training.
+Figure 7 shows the average proportion of time spent in range (TIR) by virtual humans from different age groups (adults, adolescents and children) when employing the predictive ML model.
+We see that individuals across all groups spend over 85% of their time in range compared to 77% when using only the reactive safe model.
+These results provide a baseline for an algorithm that is effective in managing T1D.
+Reactive TIR with TIR with
+safe predictive clamped
+Age group model ML model ML model
+Adults 88.25% 93.49% 92.88%
+Adolescents 81.98% 85.23% 82.38%
+Children 77.31% 86.6% 85.42%
+Figure 7: Average proportion of time spent in range (TIR) across 21 virtual humans with a reactive safe model, a predictive ML model, and clamping the predictive ML with the reactive safe model for security.
+Incorrect ML dosing Since GlucOS is designed to be flexible with any predictive algorithm, it could be subject to attacks from malicious ML algorithms.
+To simulate attacks, we intentionally introduce an order of magnitude difference to the amount of insulin predicted by the baseline predictive ML model.
+In one experiment, we dose one-tenth of the amount of insulin and in another experiment, we dose ten times the amount of insulin computed by the predictive ML model.
+We run these malicious ML algorithms within the simulator on each virtual human on the same scenarios as the previous experiment.
+Figure 8 and Figure 9 show the CGM readings on three randomly chosen virtual humans on a particular scenario for the experiments where we dose less and dose more respectively.
+When dosing less, we see that all of them spend most of their time above Level 1 hyperglycemia.
+When dosing more, all individuals fully go below Level 2 hypoglycemia, which would have killed them.
+Figure 8: Variation in CGM readings for three virtual humans when running a predictive ML algorithm that intentionally doses one-tenth the amount of the required insulin.
+The CGM readings staying mostly above the Level 1 hyperglycemic range.
+Figure 9: Variation in CGM readings for three virtual humans when running a predictive ML algorithm that intentionally doses ten times the amount of the required insulin.
+The CGM readings staying below the Level 2 hypoglycemic range would have likely been fatal to all three individuals.
+Across all virtual humans, we see an average drop in TIR from 88.44% with correct dosing to 23.32% when dosing less and an average drop in TIR from 88.44% to 0% when dosing more.
 Correspondingly, when dosing less, we observe an average increase in the time in hyperglycemia from 3.77% with correct dosing to 65.05%.
@@
 These results further motivate the need for security mechanisms when providing individuals with the flexibility to use any algorithm to manage T1D.
-10.3 What is the impact of GlucOS’s algorithmic security mechanism on managing T1D?
-GlucOS’s algorithmic security mechanisms should keep the blood glucose of individuals in range even in presence of malicious ML algorithms.
-The mechanisms should also ensure that they do not significantly dampen the level of control provided by an efficient algorithm.
-Impact on malicious algorithms.
+8.1.2 With algorithmic security mechanism
+Impact on incorrect dosing.
 We repeat the experiments from the previous section with the same scenarios.
-We run the same malicious algorithms, but also run the reactive safe model to set bounds on the amount of insulin the malicious algorithms can inject.
-Figure 11 and Figure 12 show CGM readings on the same 3 random virtual patients on the same scenarios from the experiment in the previous section.
-When running either malicious model, we see that all patients spend over 70% of their time in range.
-These results show that GlucOS protects individuals even in presence of intentionally malicious algorithms.
-On average, across all virtual patients, we still see a drop in TIR from 88.44% with correct dosing to 73.89% when employing safety while dosing less and a drop in TIR from 88.44% to 78.22% when dosing more.
-While the times spent in range is less than correct dosing, the impact of incorrect dosing is significantly dampened with patients spending the recommended amount of time in range.
-Figure 11: Variation in CGM readings for 3 virtual patients when clamping an algorithm that intentionally tries to dose one-tenth the amount of the required insulin.
+We run the same models for incorrect dosing, but also run our reactive safe model (Section 3) to set bounds on the amount of insulin the malicious algorithms can inject.
+Figure 10 and Figure 11 show CGM readings on the same three virtual humans on the same scenarios from the experiment in the previous section.
+With both types of incorrect dosing, we see that all of them spend over 70% of their time in range.
+These results show that GlucOS protects individuals even in presence of malicious incorrect dosing.
+On average, across all virtual humans, we only see a drop in TIR from 88.44% with correct dosing to 73.89% when employing safety while dosing less and a drop in TIR from 88.44% to 78.22% when dosing more.
+While the times spent in range is less than that of correct dosing, the impact of incorrect dosing is significantly dampened with individuals spending the recommended amount of time in range, despite facing an active attack.
+Figure 10: Variation in CGM readings for three virtual individuals when clamping an algorithm that intentionally tries to dose one-tenth the amount of the required insulin.
 On average, the CGM readings stay in range over 70% of the time.
-Figure 12: Variation in CGM readings for 3 virtual patients when clamping an algorithm that intentionally tries to dose ten times the amount of the required insulin.
+Figure 11: Variation in CGM readings for three virtual humans when clamping an algorithm that intentionally tries to dose ten times the amount of the required insulin.
 On average, CGM readings stay in range over 70% of the time.
-Patient TIR with clamped
-type Predictive ML model
-Adults 92.88%
-Adolescents 82.38%
-Children 85.42%
-Figure 13: Average proportion of time spent in range by 21 virtual patients in the same scenarios in Section 10.1 when clamping the predictive ML with the reactive safe model.
 Impact on effective benign algorithms.
-We evaluate the impact of GlucOS’s algorithmic security mechanisms on benign algorithms such as the predictive ML model used in Section 10.1.
-We repeat the experiment of running the predictive ML model on all virtual patients in the same random scenarios while maintaining the same parameters for the clamps and reactive model used for the malicious algorithms evaluation.
-Comparing Figure 13 against Figure 8 we see that there is only a slight drop in the average proportion of time spent in range by virtual patients across all age groups when employing GlucOS’s algorithmic security mechanisms.
-The average proportion of time spent in range being higher than the reactive safe model shows that GlucOS’s security does not come at the cost of the benefits from ML.
+We evaluate the impact of GlucOS’s algorithmic security mechanism on benign algorithms that are effective in managing T1D such as the baseline predictive ML model.
+We repeat the experiment of running the predictive ML model on all virtual humans in the same random scenarios while maintaining the same parameters for the clamps and reactive model used to evaluate incorrect dosing.
+Comparing data in Figure 7 we see that there is only a slight drop in the average proportion of time spent in range by virtual humans across all age groups when employing GlucOS’s algorithmic security mechanism.
+When employing the reactive safe model as is to calculate insulin doses, we report an average TIR of 82.48% which is lower than the average TIR of 86.84% observed when running the predictive ML model with GlucOS’s security mechanism.
+This shows that GlucOS’s algorithmic security guarantees do not come at the cost of effective T1D management from predictive ML.
 Figure 12: We attached an Omnipod DASH insulin pump to a stuffed animal and modified the OmniBLE pump driver to intentionally add, drop and modify insulin commands.
 Our driver security mechanism detected all manipulations.
-10.4 Does GlucOS ensure pump operation while overcoming buggy/malicious pump drivers?
+8.2 Does GlucOS ensure pump operation while overcoming malicious pump drivers?
 We first evaluate GlucOS’s driver security mechanism in terms of its impact on benign pump drivers that accurately relay pump communications.
-We then evaluate the impact on detecting additional, modified or dropped communication from a malicious pump driver.
+We then evaluate the impact on detecting additional, modified, or dropped communication from a malicious pump driver.
 Our mechanism should ensure that all communications go through with a benign pump driver and should catch all manipulated communication with a malicious pump driver.
@@
 These results show that GlucOS can ensure pump operation while withstanding attacks from malicious drivers.
-Patient TIR without TIR with
-type biological invariant biological invariant
-Adults 49.93% 88.92%
-Adolescents 55.83% 75.26%
-Children 67.41% 77.45%
-Figure 15: Average proportion of time spent in range by 21 virtual patients when fluctuating insulin sensitivities.
-We compare cases with and without our end-to-end security mechanism using the biological invariant.
-10.5 Is the biological invariant effective?
-We run experiments on the simulator to evaluate if GlucOS’s use of the biological invariant to transition to manual insulin dosing leads to better T1D management.
-Since the simulator does not provide support to fluctuate insulin sensitivity, we intentionally scale up or scale down values for insulin sensitivity when calculating insulin doses in closed loop.
-We trigger cases where we scale down the insulin sensitivity to simulate scenarios where the system assumes a lower sensitivity value than the user’s actual sensitivity more frequently since this are more relevant to evaluating the biological invariant.
-Across all virtual patients, we record an average TIR of 57.56% when fluctuating insulin sensitivity without incorporating the biological invariant.
-In contrast, we record an average TIR of 80.54% when incorporating the biological invariant to transition to manual dosing.
+Age Group TIR without
+biological invariant
+Adults 49.93%
+Adolescents 55.83%
+Children 67.41%
+Figure 13: Average proportion of time spent in range across 21 virtual humans when fluctuating their insulin sensitivity.
+The low values for TIR stress the importance of an end-to-end security mechanism, i.e., the biological invariant.
+8.3 Is the biological invariant important?
+We run experiments on the simulator to assess the risks of not considering our end-to-end security mechanism, i.e., the biological invariant.
 Figure 13 shows the average TIR across age groups.
-We clarify that we use the correct insulin sensitivity values to compute bolus doses when dosing manually, since users would know their true insulin sensitivity when taking in manual doses.
-10.6 Did formal verification help with correctness?
-During the formal verification process of our secure insulin delivery system, we uncovered various bugs and inconsistencies that could have compromised the safety and reliability of the system in niche scenarios.
-We identified a total of 9 bugs across different categories.
-3 of these bugs were related to invalid ranges for variables such as negative value returns, domain-specific out of range values, potential division by zero errors etc.
-in the insulin dosing calculations which were resolved by adding appropriate checks and error handling mechanisms to prevent such operations and ensure graceful handling of these cases.
-Additionally, we found 4 instances where the implemented functions violated their intended postconditions along the insulin delivery path, which were addressed by thoroughly reviewing and modifying the implementations to correctly satisfy the postconditions, preserving the intended behavior and meeting our claims for critical funcrions.
-Moreover, our analysis uncovered 2 situations where certain functions lacked proper error handling mechanisms, potentially leading to undefined behavior or system failures.
-By addressing these bugs and incorporating formal verification throughout the development lifecycle, we significantly enhanced the robustness, reliability, and safety of our secure insulin delivery system.
+Across all virtual humans, we record an average TIR of 57.56%, even with our algorithmic security mechanism, when fluctuating insulin sensitivity.
+This corresponds to an average drop of 29.28% as a result of not considering the biological invariant.
+These results emphasize the importance of incorporating an end-to-end invariant and undermines results obtained from pure simulation.
+We present case-studies involving the biological invariant with a real user using GlucOS in Section 9.3.
 9 Using GlucOS in the real world
-In this section, we report on our experiences working with one individual, Bob, since November 2023.
-We show our results for using GlucOS with our reactive safe model closed-loop algorithm as well as with our predictive ML model closed-loop algorithm paired with our security mechanism.
-We also measure system performance, report on errors, and quantify our implementation simplification efforts.
-Overall, our study is a correlation study, where we evaluate one individual using our system and compare segments of time with each other.
+In this section, we report on our experiences working with seven individuals using GlucOS to help manage their T1D.
+Six people are students who only used the predictive alerting feature of the BioKernel (Section 6) for two months, from June 2024 - August 2024.
+One individual, who we will call Bob, has been using the full GlucOS system since November 2023 for insulin delivery and predictive alerting.
+Our user study for the six students was approved by our university’s IRB, and Bob’s use GlucOS was determined by our university’s IRB to be “self evaluation” and exempt from IRB.
+In our implementation, the BioKernel is an iOS app that consists of 5.1k lines of code in our trusted computing base.
+It communicates with a Dexcom G6 or G7 CGM and Omnipod DASH insulin pump via Bluetooth low energy, and runs the core closed-loop algorithm.
+For the BioKernel, we use the Swift programming language, the Actor abstraction for data-race-free concurrency, and JSON files stored on disk to capture persistent state.
+Although the BioKernel is a fully featured and stand alone automated insulin delivery system, GlucOS uses an event logging system that enables separate apps to implement complementary features while remaining isolated from the BioKernel.
 9.1 Ethical considerations
-In this section, we outline the steps we take to ensure that we uphold high ethical standards and ensure that the human using our system, who we will call Bob, is safe.
-First, one of the co-authors of this paper is a board-certified Endocrinologist, who specializes in T1D.
+In this section, we outline the steps we take to ensure that we uphold high ethical standards and ensure that the people using our system, are safe.
+One of the co-authors of this paper is a board-certified Endocrinologist, who specializes in T1D.
 They are both a clinician with an active practice seeing patients living with T1D and a researcher who focuses on research on T1D.
-They designed the safety protocol that we detail in Appendix A, which defines the testing we conduct to ensure that the individual using our system is safe, and the specific criteria we use to stop the study, if needed.
-Second, we submitted our study to our university’s IRB, which determined that our “self evaluation” is exempt from IRB (case number omitted, but available).
-Third, Bob increased the frequency of his visits to his Endocrinologist, which includes lab work to confirm that he remained in good overall metabolic health.
-Overall, GlucOS system had a large positive impact on the individual who used it.
-Bob’s lab-based A1C test (a blood test that measures average glucose levels over time), which includes his time using GlucOS, was 5.8%.
-That is nearly non-diabetic levels of control – healthy individuals will have an A1C of 5.6% or lower.
-This result is a profound outcome for Bob, and GlucOS had a massive positive impact on Bob’s health and quality of life.
+They designed the safety protocol, which defines the testing we conduct to ensure that the individuals using our system are safe, and the specific criteria we use to stop the study, if needed.
+Appendix A includes a more detailed discussion of our ethical considerations and safety protocol.
 9.2 Is an ML-based closed-loop system safe and effective in practice?
@@
 We believe that this distribution represents a suitable balance for providing the ML with enough flexibility to improve outcomes, while having tight enough bounds to limit potential damage from bad or malicious ML predictions.
-11.3 How does the BioKernel perform?
-We look at two aspects of system performance when evaluating our real-world use of GlucOS.
-First, we measure how long our closed-loop algorithm takes to run and second, we count how often the closed-loop algorithm runs successfully end-to-end.
-We use Bob’s iPhone 14 to collect all of the data we show here and these numbers are averages across the entire six month period.
-On average, the BioKernel spends 4.3s running the closed loop algorithm, where almost half of that time is dominated by CGM and pump communications.
-Of the components within the closed-loop algorithm, the ML predictions take 2.2s on average and the safety logic takes 19ms (0.019s).
-Since the system runs in the background, we believe that iOS runs the device with reduced computational capacity to save power.
-But overall, 4.3s of wall-clock time for every five minutes our closed-loop algorithm runs with only a tiny fraction of that coming from our safety logic represents practical system performance.
-For successful closed-loop end-to-end execution, over the evaluation period the closed-loop algorithm runs successfully 96.5% of the time, with 2.1% of the runs failing due to pump communication failures, which is our largest class of failure.
-11.4 How simple is our BioKernel?
-To evaluate the complexity of our BioKernel, we count the lines of code in our implementation and compare against Loop, another open-source automated insulin delivery system.
-Using the “cloc” utility the BioKernel app has 4.0k lines of code compared to 39.8k lines of code in the Loop app, an order of magnitude reduction.
-We omit the lines of code coming from LoopKit, which both the BioKernel and Loop use, because our pump security measures remove most of it from our trusted computing base.
-However, LoopKit and the drivers have a substantial amount of code, weighing in at 70.1k lines of code, showing the importance of our security measures.
-To explain why we have such a large reduction in source code, we outline the differences between the two systems.
-First, Loop has several features that we move outside of the BioKernel.
-These features include a remote interface for insulin dosing, a Watch app, Siri command interfaces, third party libraries, tutorials, complex meal announcements, and physiological simulations.
-Second, we simplify our implementation of features that are shared between the BioKernel and Loop.
-These shared features include a simplified local storage implementation, simplified concurrency support, minimal therapeutic settings, and a simplified closed-loop algorithm.
-As our evaluation shows, we can still provide tight control over T1D with this simplified core system and still support most or all these additional features architecturally.
+9.3 Do biological invariant violations trigger in practice?
+To determine if biological invariant violations trigger in practice, we review 90 days of data for Bob, from May 29th, 2024 to August 27th, 2024.
+During this period of time, biological invariant violations occurred 1.6k times, or 9.0% of the time.
+Of these violations there were 49 sequences of 30 minutes or longer with the longest sequence being 65 minutes.
+These results suggest that although natural variations in glucose metabolism do lead to biological invariant violations, none of them were long enough to trigger our most severe action of transitioning to manual mode after two hours.
+We reviewed one weeks worth of violations with Bob manually to determine the cause. During the most recent week of data, Bob had five sequences of 30 minutes or longer and all were caused by exercise.
+9.4 Are predictive alerts useful for security?
+To measure predictive alerts as a security mechanism (Section 6), we recruited six students who live with T1D and had them use the GlucOS predictive alerts to complement their default CGM alerts.
+Our primary metric is time spent below 70 mg/dl, where a decrease in the amount of time spent in hypoglycemia suggests that predictive alerts are effective to empower people to mitigate security attacks.
+For the two month study, our participants decreased their average time in hypoglycemia from 2.7% down to 1.6%, and all individuals experienced a decrease in the amount of time spent in hypoglycemia.
+For context, participants in a state-of-the-art automated insulin delivery clinical trial experienced a decrease of time in hypoglycemia of 0.9% [15], showing that our 1.1% decrease is substantial.
+Overall, our results suggest that predictive alerts can be a useful last line of defense for trustworthy automated insulin delivery systems.
 10 Related work
+In addition to the work already discussed in this paper, there are several other related works.
 Previous research has looked at the security of implanted medical devices in general [32, 16, 60], in addition to looking at insulin pumps in particular [39, 56], with more recent work looking at providing improved security [45, 9].
 Also, recent work has looked at applying formal methods to insulin pumps for high assurance [55].
-These works focus on the device and their communication channel.
+These works focus on devices and their communication channels.
 In contrast, with GlucOS we assume that these devices are correct and secure and focus our efforts on the software we use to run the automated insulin delivery system.
-Our work on coping with malicious pump drivers builds on previous work on malicious device drivers in operating systems [62], [15], [27], [28], [43], [31].
-Our work applies these principles to a new domain: automated insulin delivery systems.
+Researchers have started to look at combining untrustworthy ML with more traditional forms of automated decision making, like Markov Decision Processes, from a theoretical perspective [41, 50, 57, 20, 40].
+These theoretical works focus on providing algorithms to prove desirable properties about a combined ML / Markov system, like robustness and stability.
+In contrast, we take advantage of domain knowledge and formulate a simple and practical system for one specific problem: automated insulin delivery.
 11 Conclusion
-Our audacious long-term goal is to turn a T1D diagnosis from a death sentence into an indicator of longevity, where people living with T1D will be expected to live longer than their healthy peers.
-This longevity will come by virtue of the tight control that they maintain over their metabolism through advanced computer systems.
-Our first step towards this goal is to ensure that people can use trustworthy computer systems to manage their glucose.
-Appendix A Protocol for real world use
-One individual, who we will call Bob, used GlucOS for 2.5 months.
-We put in place a protocol where we would stop our study and ask Bob to cease using our system and revert back to his previous insulin regime.
-In our protocol, we have formal in-person meetings with Bob every week to check his data and make sure that he is not incurring risk as a result of using GlucOS.
+The design of GlucOS, while grounded in established OS and security principles, incorporates several non-obvious elements that emerged from our deep engagement with the problem domain and real-world testing.
+First, our approach to algorithmic security used classic security techniques, like reference monitors and the Simplex architecture.
+However, it was our novel insight around all correct algorithms dosing the same amount of insulin that led us to these classic techniques.
+Plus, from our exploration, the most interesting part was not the ML, it was around the principles of reactive safe models and in particular our principle that these safety-critical algorithms should be reactive.
+Second, our pump driver security went beyond traditional specification and checking.
+We needed to actively program the pump to handle dropped cancel commands, which was an unexpected but crucial insight.
+Third, the biological invariant emerged as a key check in our system, addressing a blind spot in typical simulations that ignore changing human dynamics.
+This forced us to develop novel formal methods to handle scenarios where our glucose metabolism models could be drastically wrong.
+Fourth, our approach to human interaction in the system was nuanced.
+We found that simply alerting humans or suspending insulin delivery were not always appropriate.
+For security alerts, personalization and customization were important.
+Stopping insulin delivery could be dangerous if the user has high glucose and insufficient insulin on board, which required more context-aware responses.
+These non-obvious design decisions underscore the complexity of creating a truly trustworthy automated insulin delivery system.
+They reflect the value of combining theoretical security principles with practical, real-world testing and domain-specific knowledge.
+This interdisciplinary research drew on deep expertise from computer security, medicine, and ethics to create a practical system.
+Our work improves the security of automated insulin delivery systems and also provides a framework for thinking about security in other critical healthcare technologies.
+Our exploration is an extreme example of personalized, distributed, and automated healthcare using trustworthy computer systems.
+Appendix A Ethical and safety considerations
+One individual, who we will call Bob, used GlucOS for 9 months.
+Six university students also used the predictive alerts from GlucOS for 2 months.
+We put in place a protocol where we would stop our study and ask people to stop using our system and revert back to their previous management regime.
+In our protocol, we have formal in-person meetings with Bob every week to check his data and make sure that he is not incurring risk as a result of using GlucOS, and weekly surveys for the six students, plus monthly meetings.
 We would have stopped our study if any of the following conditions occurred:
   • Excessive hypoglycemia.
-We define excessive hypoglycemia as having spent 5% or more of time during a week with CGM readings below 70 mg/dl or more than 1% below 54 mg/dl.
-These amounts of time spent in hypoglycemia would have represented a significant increase in time spent in hypoglycemia for Bob.
-  • Insufficient CGM data. If Bob would have had less than 70% of the time during a week with CGM readings, then the lack of CGM data would represent a decrease for Bob.
+We define excessive hypoglycemia as having spent 8% or more of time during a week with CGM readings below 70 mg/dl or more than 1% below 54 mg/dl.
+These amounts of time spent in hypoglycemia would have represented a significant increase in time spent in hypoglycemia for participants.
+  • Insufficient CGM data. If they had less than 70% of the time during a week with CGM readings, then the lack of CGM data would represent a decrease for the participants.
   • Insufficient closed-loop runs.
@@
   • Serious complications due to diabetes.
-Serious complications include any hospitalization for diabetes related issues, severe hypoglycemia where Bob was unable to recover from hypoglycemia himself and had to get help from someone else, diabetic ketoacidosis, or seizure.
-  • New medical diagnosis requiring significant attention. If Bob had received a new medical diagnosis during the study, the study would have had the potential to be a distraction for Bob.
-  • New mental health diagnosis. If Bob were struggling with mental health during the study, the study would have the potential to be a distraction.
+Serious complications include any hospitalization for diabetes related issues, severe hypoglycemia where a participant was unable to recover from hypoglycemia themselves and had to get help from someone else, diabetic ketoacidosis, or seizure.
+  • New medical diagnosis requiring significant attention. If anyone had received a new medical diagnosis during the study, the study would have had the potential to be a distraction for participants.
+  • New mental health diagnosis. If anyone were struggling with mental health during the study, the study would have the potential to be a distraction.
+In addition to our safety protocol, we followed best practices for ethical research to ensure that we conducted our study ethically.
+  • IRB.
+Our study went through our university’s IRB process, where the study including six university students who used GlucOS’s alerting software was approved, and our case study on Bob was determined to be “self evaluation” and exempt from IRB.
+  • Informed consent. All participants went through a written informed consent process.
+  • Data management.
+GlucOS does not store any personally identifiable information and uses access control on our Google Cloud Platform infrastructure to ensure that only the authors of the paper can access this anonymized data.
+  • Risk/benefit analysis.
+All participants are already living with T1D and use CGMs and inject insulin, so they already take on the risks inherent in managing T1D.
+The benefits include improved control over their T1D.
+There are risks in using experimental software, but all participants continued to rely on their CGM alerts as a last line of defense.
+  • Participant selection. We selected participants who were university students living with T1D.
+  • Adverse event reporting. Had any participants experienced adverse events during the study, we would have reported it to our IRB for guidance and addressed it according to our safety protocol.
+  • Communication. All surveys were done through our university’s medical school infrastructure for private communication with individuals.
+Since Bob used the full GlucOS software, including for insulin delivery, we included several other safety mechanisms:
+  • Bob increased the frequency of his visits to his Endocrinologist while using the GlucOS software. He went from yearly appointments to quarterly appointments.
+  • During the study, twice Bob got lab work based on blood tests to provide an independent measure of his overall metabolic health.
+This lab work helped to ensure that he was in good health and that there were no hidden negative impacts from his use of GlucOS.
+Overall, GlucOS had a large positive impact on the individuals who used it.
+Bob’s lab-based A1C test (a blood test that measures average glucose levels over time), which includes his time using GlucOS, was 5.8%.
+That is nearly non-diabetic levels of control – healthy individuals will have an A1C of 5.6% or lower.
+All of the students decreased the amount of time spent in hypoglycemia.
+These results are profound for the seven people using GlucOS, and it had a positive impact on their health and quality of life.
+Finally, Bob was a willing and enthusiastic user of the GlucOS system.
+To be clear, Bob is not a graduate student who was forced to use GlucOS so that his advisor could publish a paper.
+Rather, Bob struggled with the cognitive load of T1D management and wanted an automated insulin delivery system to offload part of the burden.
+However, he was unwilling to use any of the other automated insulin delivery systems out there due to concerns over their security.
+Bob found that GlucOS, with its focus on security, correctness, and simplicity, met his needs.
 Appendix B List of pump setup and UX communication messages
@@
 Figure 16: GlucOS’s CGM state machine to handle missing readings from CGMs.
-Appendix C CGM state machine
-Inability to retrieve readings from a CGM also count as runtime failures in addition to those discussed in Section 7.
-However, these failures are easier to handle.
-In this section, we describe the CGM state machine that handles such failures.
-We note that transitions from the state machine described in Section 7 supersede state transitions from the CGM state machine.
+Appendix C Runtime failure management
+Figure 17: GlucOS’s state machine to transition users to the safest manual dosing state when losing trust in insulin delivery (i.e., with a buggy or malicious pump driver) or when the biological invariant is not satisfied.
+In this section, we describe GlucOS’s state machine that combines its security components, i.e., algorithmic security (Section 3), driver security (Section 4) and end-to-end security (Section 5) together to protect users in light of their corresponding failures.
+GlucOS’s state machine consists of three operating states: (1) ML with safety operating in closed loop, (2) reactive safe model operating in closed loop and (3) manual dosing in closed loop.
+ML provides the tightest level of control in managing glucose but poses the largest immediate threat from excess insulin since it preemptively injects insulin by anticipating rise in glucose.
+The reactive safe model provides slightly lesser control, but poses milder immediate threat since it only injects insulin to account for measured excess glucose.
+Manual dosing offers the least amount of control but poses no threat since it only injects the minimum insulin needed for sustenance.
+GlucOS predominantly operates in the state running ML (with safety) in closed loop.
+We transition to the safest manual dosing state when we do not trust insulin delivery (i.e., with a buggy or malicious pump driver).
+We also transition to the safest manual operation when the biological invariant is not satisfied to ensure that we do not risk the possibility of removing too much glucose from our insulin doses.
+If we end up in the manual dosing state, we enforce that only manual user intervention can transition back to other states.
+We visualize these state transitions in Figure 17.
+Inability to retrieve readings from a CGM also count as runtime failures. However, these failures are easier to handle.
 Since ML needs a stream of CGM readings as input, we transition from ML to the reactive safe model if we don’t receive an input from the CGM.
@@
 Figure 18: Report of bugs fixed by formally verifying GlucOS.
-Appendix D Formal verification: invariants and bugs
+Appendix D Did formal verification help with correctness?
+During the formal verification process of our secure insulin delivery system, we uncovered various bugs and inconsistencies that could have compromised the safety and reliability of the system in niche scenarios.
+We identified a total of 9 bugs across different categories.
+3 of these bugs were related to invalid ranges for variables such as negative value returns, domain-specific out of range values, potential division by zero errors etc.
+in the insulin dosing calculations which were resolved by adding appropriate checks and error handling mechanisms to prevent such operations and ensure graceful handling of these cases.
+Additionally, we found 4 instances where the implemented functions violated their intended postconditions along the insulin delivery path, which were addressed by thoroughly reviewing and modifying the implementations to correctly satisfy the postconditions, preserving the intended behavior and meeting our claims for critical funcrions.
+Moreover, our analysis uncovered 2 situations where certain functions lacked proper error handling mechanisms, potentially leading to undefined behavior or system failures.
+By addressing these bugs and incorporating formal verification throughout the development lifecycle, we significantly enhanced the robustness, reliability, and safety of our secure insulin delivery system.
+Appendix E Formal verification: invariants and bugs
 In this section, we list the invariants used to formally verify critical functions of GlucOS. Figure 18 lists the bugs we found and fixed during formal verification.
-D.1 Invariants
+E.1 Invariants
 The MicroBolusing function ensures that micro boluses are always administered at safe time intervals while making sure that they are actually needed (the glucose level of the person is in fact high).

### v3

@@
 GlucOS: Security, correctness, and simplicity for automated insulin delivery
-Abstract
+Abstract.
 We present GlucOS, a novel system for trustworthy automated insulin delivery.
@@
 This work demonstrates the potential for secure, personalized, automated healthcare systems.
-1 Introduction
+Our source code is open source.
+1. Introduction
 Type 1 Diabetes (T1D) is a metabolic disorder where an individual’s pancreas stops producing insulin.
@@
 Computer systems, called automated insulin delivery systems, exist that use subcutaneous sensors to continuously monitor glucose concentrations (glucose is the body’s primary energy source) and inject insulin automatically using an insulin pump device (Figure 1).
-However, insulin is a dangerous hormone where too much insulin can kill people in a matter of hours [22] and too little insulin can kill people in a matter of days [13].
+However, insulin is a dangerous hormone as too much insulin can kill people in a matter of hours (cryer2012severe, ) and too little insulin can kill people in a matter of days (benoit2018trends, ).
 For automated insulin delivery systems, the key challenge is to maintain a balance: provide enough insulin to prevent dangerously high glucose levels while avoiding excessive insulin that could lead to life-threateningly low glucose levels.
 In this context, security means ensuring integrity for insulin dosing against malicious attacks or inadvertant errors.
-Several commercial [17, 5, 3] and open-source [42, 54] automated insulin delivery systems exist today, but none of them consider security as a primary design constraint.
+Several commercial (beta_bionics, ; tandem_iq, ; omnipod, ) and open-source (loop, ; openaps, ) automated insulin delivery systems exist today, but none of them consider security as a primary design constraint.
 (a) Devices for automated insulin delivery. This figure is from breakthrought1d.org.
 (b) Closed-loop architecture.
-(c) Our system for automated insulin delivery.
-Figure 1: Overview of automated insulin delivery devices, closed-loop architecture, and a screenshot of GlucOS.
-The complexity of automated insulin delivery systems and the fact that they make real-time healthcare decisions, makes their security paramount.
-Current research on security for these systems focuses primarily on evaluating algorithms using simulations [71, 72].
-We argue that this approach falls short of addressing the full spectrum of challenges inherent in building trustworthy systems that operate within real-world constraints.
-Real-world systems must cope with runtime platform limitations, changes in human physiology, and the necessity of human interaction with the system.
-Plus, algorithms and simulations model attacks abstractly through their glucose response, missing out on details like which components of the system are likely to be compromised or how the system architecture can prevent damage from these potentially vulnerable components.
-Our experience shows that when faced with these practical challenges, algorithms designed in simulations often prove insufficient or inapplicable.
-These real-world complexities require a fundamentally different approach to designing security for automated insulin delivery systems.
-To address these challenges, we need practical systems that move beyond simulation-based designs and tackle the challenges of real-world systems.
+(c) Overall GlucOS architecture.
+Figure 1. Overview of automated insulin delivery devices, closed-loop architecture, and GlucOS.
 Current open-source automated insulin delivery systems are complex.
-For example, the OpenAPS [54] core function for calculating insulin doses consists of 1192 lines of Javascript code, 63 input and configuration parameters, and 90 branch statements.
+For example, the OpenAPS (openaps, ) core function for calculating insulin doses consists of 1192 lines of Javascript code, 63 input and configuration parameters, and 90 branch statements.
 All of this logic influences the ultimate dosing decision, making reasoning about its correctness, both manually and automatically, challenging.
@@
 In contrast, our implementation of this same function is 25 lines of native and formally verified Swift code.
-None of the open-source automated insulin delivery systems use formal methods, can withstand attacks on components, or have mechanisms or policies for security.
+None of the open-source automated insulin delivery systems uses formal methods, can withstand attacks on components, or has mechanisms or policies for security.
 In this paper, we take on the challenge of building the first trustworthy automated insulin delivery system, called GlucOS.
@@
 We provide clear security boundaries, mechanisms and policies to mitigate attacks on vulnerable components, and formal methods to provide assurances that our implementation is correct.
-Automated insulin delivery systems are different from traditional applications that the security community focuses on, but they are important.
+Automated insulin delivery systems are different from traditional applications that the systems and security communities focus on, but they are important.
 Research on operating systems or web browsers can list scores of CVEs from current systems and show how their system improves security by eliminating the impact of some or all of these vulnerabilities by virtue of their design.
-Automated insulin delivery systems do not have a long list of CVEs that we can reference, although there are a few [46, 47, 48, 49].
+Automated insulin delivery systems do not have a long list of CVEs that we can reference, although there are a few (cve1, ; cve2, ; cve3, ; cve4, ).
 However, our position is that we need to make sure that these systems never get CVEs, because if they do, security vulnerabilities can kill people.
@@
   • Security mechanisms and policies that protect individuals from malicious ML, vulnerable pump drivers, and drastic changes in the human’s glycemic response.
-  • The application of formal methods to prove correct the most critical parts of GlucOS.
-  • A biological invariant that combines end-to-end physiological security checks with formal methods.
-  • A case study describing our experiences from a real-world deployment with seven individuals, and results from simulation to show that our techniques generalize.
-  • A human-centric design that considers the user as an integral part of the system, influencing both the overall architecture and including them in our defensive strategy.
-We give top priority to the health, safety, and ethics of the humans using GlucOS to manage their T1D. Section 9.1 summarizes our safety protocol and Appendix A provides a detailed description.
-There are 8.4 million people living with T1D worldwide [30], underscoring the need for providing trustworthy systems for automated insulin delivery.
-Our work, carried out with rigorous medical and ethical oversight, is a first step in addressing this need.
-2 Overview
+  • A case study describing our experiences from a real-world deployment with seven people, and results from simulation to show that our techniques generalize.
+  • A human-centric design that considers the user as an integral part of the system, influencing both the architecture and including them in our defensive strategy.
+We give top priority to the health, safety, and ethics of the humans using GlucOS to manage their T1D. Section 10.1 summarizes our ethical considerations.
+There are 8.4 million people living with T1D worldwide (t1d_numbers, ), underscoring the need for providing trustworthy systems for automated insulin delivery.
+Our cross-disciplinary work, carried out with rigorous medical and ethical oversight, is a first step in addressing this need.
+2. Overview
 This paper describes our design for GlucOS, a system for trustworthy automated insulin delivery; we have two primary goals.
 First, we want the software to be simple and correct.
-Second, we want to identify the most vulnerable parts of the system and be able to prevent attacks on these components or withstand successful attacks.
+Second, we want to identify the most vulnerable parts of the system and design a system to prevent attacks on these components or withstand successful attacks.
 This section describes our overall architecture for GlucOS that strives to achieve these goals, and our threat model.
-2.1 Locking down the insulin delivery path
+2.1. Locking down the insulin delivery path
 We focus our system architecture on providing security mechanisms and policies on the insulin delivery path.
@@
 Together, these components are both the most complex, and the most sensitive from a security perspective because they are responsible for insulin delivery.
-2.2 Architecture
+2.2. Architecture
 From a system design perspective, we strive to keep our implementation simple and enable our use of formal methods to prove correct the most critical parts.
-For our software system architecture, we use separation principles from the OS and microkernel areas [8, 33] applied to the application layer to push complexity away from the most critical parts of the system.
-This architecture is similar to secure web browsers [31, 67, 63, 59], which also apply OS separation principles at the application layer.
-By decomposing the overall automated insulin delivery system, we strive to keep our trusted computing base simple.
-Figure 2: Overall architecture for GlucOS.
-Figure 2 shows our overall system architecture.
-At the core of our overall architecture is the BioKernel.
-The BioKernel is the component that interacts with the CGM and insulin pump hardware, executes the closed-loop dosing algorithm, runs security and safety checks, and produces event logs that other components use to learn the state of the system.
+For our software system architecture, we use separation principles from the OS and microkernel areas (accetta1986mach, ; hartig1997performance, ) applied to the application layer to push complexity away from the most critical parts of the system.
+This architecture is similar to secure web browsers (grier2008secure, ; wang2009multi, ; tang2010trust, ; reis2019site, ), which also apply OS separation principles at the application layer.
+Figure 1(c) shows how we decompose our automated insulin delivery system, where we strive to keep our trusted computing base simple.
+GlucOS includes a BioKernel that forms our core trusted computing base.
+The BioKernel logs events that other components can ingest to learn the state of the system.
+These other components include apps for training ML models, visualizing statistics, an AI agent for suggesting dosing during meals, and an app for replaying closed-loop algorithm execution.
+Together, these components make up the overall GlucOS system.
+Figure 2. Overall architecture for the BioKernel.
+Figure 2 shows our architecture for the BioKernel.
+The BioKernel is the component that interacts with the continuous glucose monitor (CGM) and insulin pump hardware, executes the closed-loop dosing algorithm, runs security and safety checks, and produces event logs that other components use to learn the state of the system.
 The BioKernel runs the closed-loop algorithm every five minutes as the new CGM readings come in.
@@
 A pump driver converts these commands into low-level I/O to program the pump and adjust insulin delivery.
-To withstand attacks, the BioKernel serves as a reference monitor [10] to enforce security policies on the predictive ML model’s pump commands and the pump driver’s low-level I/O.
+To withstand attacks, the BioKernel serves as a reference monitor (anderson1972computer, ) to enforce security policies on the predictive ML model’s pump commands and the pump driver’s low-level I/O.
 This reference monitor architecture ensures that we can withstand attacks from malicious predictive ML as well as malicious pump drivers without having to know the internals of how they operate.
@@
 Our biological invariant is an end-to-end check of our theoretical calculations of how the individual’s metabolism should behave compared to what we observe in practice, where violations of this invariant suggest a potentially catastrophic failure of the system.
-2.3 Threat model
+2.3. Threat model
 The most serious health risk for people who use synthetic insulin is low glucose, or hypoglycemia.
@@
 Thus, our primary security objective is to reduce or eliminate hypoglycemia that results from an attack on the system.
-Quantitatively, our goal is to maintain the American Diabetes Association’s recommended time in hypoglycemia of less than 4% [12].
+Quantitatively, our goal is to maintain the American Diabetes Association’s recommended time in hypoglycemia of less than 4% (AmericanDiabetesAssociation2021, ).
 As our design philosophy centers around humans, providing flexibility for people to use whatever ML they want by loading personalized models or using any pump hardware they want is important.
@@
 We assume that the Swift type system is correct and that the underlying operating system, the CGM driver and hardware, and the pump hardware we use are free of attacks and their device communications maintain integrity.
-Thus, we do not consider attacks emerging from tampered CGM sensors or pump hardware in our threat model.
+Thus, we do not consider attacks from tampered CGM sensors or pump hardware in our threat model.
 These are important but complementary problems to our work, which focuses on the system architecture and algorithms for automated insulin delivery.
@@
 In our threat model, malicious predictive ML can issue arbitrary insulin doses.
-We also consider malicious pump drivers. In our threat model, malicious pump drivers can issue arbitrary pump I/O operations to add, drop, or modify commands sent to the pump.
+We also consider malicious pump drivers. In our threat model, malicious pump drivers can issue arbitrary pump I/O to add, drop, or modify commands sent to the pump.
 Successful attacks on predictive ML or pump drivers, if unchecked, can be devastating because they directly affect insulin delivery and can be used by an attacker to induce potentially life-threatening hypoglycemia.
-3 Algorithmic security: Withstanding malicious ML
+3. Algorithmic security: Withstanding malicious ML
 Currently, no automated insulin delivery systems use deep neural network ML.
-Some current systems use statistical methods, like linear regression or model predictive control [21, 66, 43, 58, 54].
+Some current systems use statistical methods, like linear regression or model predictive control (physio2011, ; physio2019, ; lunze2013control, ; control2019, ; openaps, ).
 The reason for using this more traditional form of ML is sound.
 It provides consistent, explainable results and has a physiological basis that people can use to vet its decisions.
-However, all of the recent advances in ML, like image classifiers and large language models, have come from deep neural networks, which researchers have shown also work well for predicting metabolic states for use in automated insulin delivery systems [73, 35, 70, 65, 74, 64].
-Despite this promise, none of the current automated insulin delivery systems incorporate more advanced deep neural networks due to the black-box nature of deep neural networks and the potentially dire consequences of mispredictions – an automated insulin delivery system’s unchecked “hallucination” [1] could have lethal consequences if it delivers an inappropriate insulin dose [22, 13].
+However, all of the recent advances in ML, like image classifiers and large language models, have come from deep neural networks, which researchers have shown also work well for predicting metabolic states for use in automated insulin delivery systems (Zhu2022, ; Jacobs, ; Zhang2021, ; Josep2020, ; RL, ; Tyler2022, ).
+Despite this promise, none of the current automated insulin delivery systems incorporate more advanced deep neural networks due to the black-box nature of deep neural networks and the potentially dire consequences of mispredictions – an automated insulin delivery system’s unchecked “hallucination” (chatgpt_hallucination, ) could have lethal consequences if it delivers an inappropriate insulin dose (cryer2012severe, ; benoit2018trends, ).
 In this section, we discuss the design and implementation of GlucOS’s algorithmic security mechanisms to enable predictive models, including deep neural network ML, to run safely in automated insulin delivery systems.
-3.1 Mechanism fundamentals
+3.1. Mechanism fundamentals
 Our key novel insight underpinning our security mechanism design is that all correct insulin delivery algorithms will deliver the same amount of insulin over a sufficiently long period.
@@
 Our approach combines the safety of a reactive safe model and the performance from a predictive model.
-Our security mechanism draws inspiration from previous work on combining simple and complex models for safety [61, 51].
+Our security mechanism draws inspiration from previous work on combining simple and complex models for safety (sha2001using, ; mohan2013s3a, ).
 However, our approach extends these ideas by providing a novel security mechanism specifically designed for insulin delivery models.
 Our key contributions lie in our insight into the security properties of insulin delivery algorithms and our principles for how to apply our mechanism to automated insulin delivery systems (Section 3.2).
-Figure 3: Algorithmic security mechanism for GlucOS.
+Figure 3. Algorithmic security mechanism for GlucOS.
 Figure 3 shows how our algorithmic security mechanism works.
@@
 However, with our mechanism in place, we limit the impact by our three-hour bound and offset by the body’s continuous glucose production.
-3.2 Model principles
+3.2. Model principles
 Our contribution is not in the specific algorithms we use, but rather in the insight around insulin delivery being equal for all correct algorithms and how we use this insight for security.
 In this section, we outline the algorithms we use in our current implementation and define the principles for how to apply security to any predictive insulin dosing algorithms.
-Algorithmically, we use a PID controller [25] from feedback control for our reactive safe model, and we use a deep neural-network from the literature for meal predictions for our predictive ML model [52].
+Algorithmically, we use a PID controller (franklin2002feedback, ) from feedback control for our reactive safe model, and we use a deep neural-network from the literature for meal predictions for our predictive ML model (mosquera2023enabling, ).
 Although we use this model in our current implementation, our system supports any predictive model.
-PID controllers are used in automated insulin delivery systems [68, 38] and research predictive ML models for insulin delivery have been well covered in the literature [73, 35, 70, 65, 74, 64].
+PID controllers are used in automated insulin delivery systems (weinzimer2008fully, ; kushner2018data, ) and research predictive ML models for insulin delivery have been well covered in the literature (Zhu2022, ; Jacobs, ; Zhang2021, ; Josep2020, ; RL, ; Tyler2022, ).
 The design of the reactive safe model is crucial for algorithmic security, as it serves as the anchor for the security mechanism.
@@
 In our evaluation, our predictive model outperforms the reactive safe model both in simulation (Section 9) and for the individual that uses our system for their real-world insulin dosing (Section 10), showing why it is important for automated insulin delivery systems to have a practical way to adopt ML.
-4 Driver security: Withstanding malicious pump drivers
+4. Driver security: Withstanding malicious pump drivers
 Pump drivers provide abstractions for insulin delivery applications to communicate with insulin pumps.
 Much like traditional device drivers, they have complex implementations since they help accomplish a wide range of tasks including pump setup/teardown, managing pump UX, handling insulin delivery/suspension, managing pump errors, accurately tracking insulin doses, and so on.
-Different pump drivers use different implementations with varying amounts of customizations and features to carry out these tasks.
-While CVEs do not currently exist for pump drivers, there is strong possibility for bugs or vulnerabilities in their implementation given their complexity and the fact that they not been formally verified.
-Since drivers control insulin pumps, these vulnerabilities can harm people using them, even lead to the death.
+Given this complexity, there is a strong possibility for bugs or vulnerabilities to be present in existing pump drivers.
+These vulnerabilities can harm or even lead to the death of individuals since pump drivers control insulin pumps.
 For example, consider a scenario where the user decides to inject 3 units (3U) of insulin using the BioKernel.
-The BioKernel would relay this as a command to the pump through abstractions provided by the pump driver.
-If the pump driver inadvertently or intentionally issues this command twice, 6U of insulin would be delivered to the user, which would put the user in danger by dosing too much insulin.
-We need a mechanism to verify that the pump driver is executing correctly.
+The BioKernel would relay this to the pump through abstractions provided by the pump driver.
+If the pump driver inadvertently or intentionally issues this command twice, 6U would be delivered, which could endanger the user by dosing too much insulin.
 Our security goals for GlucOS’s driver security are to protect users from inadvertent bugs in pump drivers and safeguard against potentially malicious pump drivers.
-To achieve these goals, we (1) introduce driver security logic in the BioKernel that interposes on pump driver-pump communications, (2) validate all commands against known high-level pump commands, and (3) detect and handle anomalies by stopping any added/modified commands and proactively handling dropping critical commands.
-This approach allows people to choose between various pump drivers while maintaining security, balancing user flexibility and safety with automated insulin delivery.
-It also keeps our trusted computing base small, simplifying formal verification.
-4.1 Driver security logic design
-Figure 4: We interpose on communications between the pump driver and the pump to handle untrusted pump drivers to deliver insulin.
-Message Message
-name description
-Request insulin delivery
-getStatus status from the pump
-Insulin dose delivery
-statusResponse information from the pump
-Insulin dose delivery
-errorResponse error from the pump
-Set default rate of
-setBasalSchedule insulin delivery
-Set temporary rate
-setTempBasal of insulin delivery
-bolus Deliver instant insulin dose
-cancelDelivery Cancel insulin delivery
-Figure 5: Types of insulin related messages exchanged between between a pump driver and an insulin pump.
-Figure 4 provides an overview of our driver security logic.
-The BioKernel registers high-level commands, like “deliver 2U of insulin” with our driver security logic, and sends them to the pump driver.
+To achieve these goals, we introduce a shim between the pump driver and the pump, which validates all communications between them.
+We design our driver security mechanism to interpose on existing pump driver-pump communication instead of directly programming the pump.
+This design ensures that our trusted computing base remains small and simple by not including the complexity of pump driver implementations, thereby enabling formal verification.
+This design also provides flexibility with altering or introducing new functionality to pump drivers.
+The most interesting aspect of our driver security mechanism was having to overrule our principle of not programming the pump in certain cases in the interest of protecting the individual.
+These cases involve commands pertaining to canceling ongoing insulin delivery.
+We could not directly adopt the design of other device driver security mechanisms (ref_validation, ; tang2010trust, ; expressos, ) for our driver security mechanism since these cases are specific to insulin pumps.
+We focus our security efforts on pump drivers because they are complex (14.5k LoC), multi-threaded, event-driven, and use two-way communication between the pump and the driver.
+In contrast, CGM drivers are simple (1.9k LoC) and only read sensor values from the device.
+In our future work, we will formally verify our CGM driver using the same techniques we use for the insulin delivery path (Section 7).
+4.1. Driver security mechanism design
+Figure 4. We interpose on communications between the pump driver and the pump to handle untrusted pump drivers.
+Figure 4 provides an overview of our driver security mechanism.
+The BioKernel registers high-level commands, like “deliver 2U of insulin” with our driver security mechanism, and sends them to the pump driver.
 The pump driver converts these commands into I/O messages to program the pump.
-However, these I/O messages come back to the driver security logic, which checks to make sure that they are consistent with the previously registered high-level pump command, before forwarding them to the pump hardware.
+The driver security mechanism then checks to make sure that these I/O messages are consistent with the previously registered high-level pump commands before forwarding them to the pump hardware.
 We only validate commands for insulin delivery since they pose the biggest threats to the user.
-We do not validate communication pertaining to pump setup / UX since users can directly detect the respective manipulations.
-Figure 5 lists the different types of messages pertaining to insulin delivery exchanged between a pump driver and the OmniBLE insulin pump, which we use in our current implementation.
-We list other message types in Appendix B.
-By interposing on the pump I/O message path, the driver security logic can ensure that the driver has not added, modified, or dropped any pump commands.
-Detecting added or modified commands is relatively straightforward.
-Our driver security logic in the BioKernel knows all high-level pump commands that it sends to the driver.
-Thus, we have a simple logic to confirm the corresponding I/O messages from the driver.
-The driver security logic stops any anomalous I/O messages before sending them to the pump.
-Detecting dropped commands is more nuanced when the command that the driver drops is a command to cancel an ongoing insulin delivery.
-Users can cancel insulin delivery if they accidentally triggered an insulin delivery command or incorrectly set its parameters.
-Since excess insulin delivery is dangerous, the driver security logic automatically sends these cancel commands to the pump when it detects a dropped insulin cancellation command.
-There are two commands that deal with canceling insulin delivery: cancelDelivery and setting the suspend option on setTempBasal.
-The cancelDelivery command cancels a previous insulin delivery command (if the delivery is still ongoing) and the suspend option on the setTempBasal command suspends all insulin delivery until the user resumes delivery.
-The most interesting lesson learned from our experience is in coping with dropped pump commands.
-In our original design, our goal was to ensure that our driver security logic only understood I/O commands and prevented any malicious I/O commands from going through.
-However, our driver security logic ended up needing to issue its own I/O commands to deal with dropped insulin delivery cancellation commands.
-This more proactive approach to driver security is different than previous approaches that also use specifications for runtime driver verification [69, 63, 44], which simply infer I/O semantics and do not issue their own I/O commands.
-In general, we avoid programming the pump directly from the driver security logic.
-But in this case, we decided to program the pump directly from the driver security logic for human safety.
-When we detect a malicious driver, we transition the system into a safe manual mode where the individual needs to take over control of their own insulin delivery.
-They can still use the BioKernel to manage their insulin, but they would be unable to use the automated closed-loop algorithms.
-In our evaluation, our system had no false positives and detected all attacks (Section 8.2).
-See Appendix C for more details on how we maintain safety even when detecting an attack.
-4.2 Driver security logic implementation
-We focus our implementation on Loop’s OmniBLE [4] pump driver.
-OmniBLE allows iOS applications to communicate with Omnipod pumps over Bluetooth.
-We include OmniBLE’s cryptography and Bluetooth communication modules as part of the BioKernel and incorporate the reference monitor just before encrypting messages sent to the pump/ decrypting messages received from the pump.
-OmniBLE does not incorporate automatic retry in its implementation.
-However, OmniBLE occasionally sends additional commands as part of its operation.
-For example, when sending a setTempBasal command, if OmniBLE’s internal states indicate that delivery is still in progress as part of a previous setTempBasal command, it first issues a cancelDelivery command to cancel that delivery.
-To remove ambiguity in terms of whether the cancelDelivery command was issued as part of setTempBasal, we enforce OmniBLE to always issue the cancelDelivery command.
-5 End-to-end security: Biological invariant
+On detecting added, modified or dropped commands from the pump driver, the driver security mechanism rejects further commands and transitions the system to a manual mode where the user takes control.
+Dropped commands pertaining to canceling ongoing insulin delivery requires separate consideration since excess insulin is dangerous and cannot be withdrawn once it has been delivered.
+Users will cancel insulin delivery if they accidentally issued an insulin delivery command or incorrectly set the insulin dosage.
+In cases where a malicious driver drops a cancel command, the driver security mechanism directly issues I/O messages to the pump to cancel insulin delivery before transitioning the pump to a manual mode.
+We make the tradeoff of increasing the size of our trusted computing base to handle these cases in the interest of protecting the user.
+Fortunately, there are only two commands to cancel insulin delivery in our current implementation, so the increase to our trusted computing base is small.
+5. End-to-end security: Biological invariant
+When implementing security systems for humans, even the perfect algorithmic and driver security mechanisms are unable to handle all situations due to changes in human physiology.
 A key novelty in our overall approach is our introduction of an end-to-end check that ensures the BioKernel’s glucose calculations against observations of the user’s metabolic state.
-This raises an interesting challenge for formal verification as it requires modeling the true biological state of the user – a parameter that is both unknown at runtime and distinct from measured parameters which are part of the software implementation.
-We refer to the true safety invariant we enforce on the user’s state as the biological invariant, and proving it is both a key part of our implementation and one of the key conclusions of our formal verification efforts.
 If there is a large mismatch between our calculations and what we observe, then it means that our calculations are wrong.
-These deviations can happen from drastic changes in the user’s glucose absorption, like during exercise when the body uses glucose without needing insulin [14], or if the pump insertion site hits a vein and delivers insulin directly into the blood stream [62].
+These deviations can happen from drastic changes in the user’s glucose absorption, like during exercise when the body uses glucose without needing insulin (borghouts2000exercise, ), or if the pump insertion site hits a vein and delivers insulin directly into the blood stream (steyn2023accelerated, ).
 Regardless of the cause, the end-to-end check detects these conditions and prevents harm to the user.
-For the purposes of the following example, suppose the end-to-end check is that the measured glucose levels are always within 30mg/dl of the predicted level; we refer to this as the implementation invariant that is enforced in an end-to-end way across the system:
-$|g_{\mathsf{measured}}-g_{\mathsf{predicted}}|<30\mathsf{mg/dl}\quad(1)$
-Naively, one might look at the $30\mathsf{mg/dl}$ on the right-hand side and think that the user’s true glucose levels are within $30\mathsf{mg}$ of the actual glucose levels.
-But this would be wrong!
-The problem is that the measured glucose level is subject to measurement error; the state-of-the-art Dexcom G7 CGM has a mean absolute relative difference of 8.2% from the true glucose level in the body [29].
-For the sake of example, suppose that the measurement error on the CGM reading is around 5%; we refer to this as the CGM invariant:
-$|g_{\mathsf{measured}}-g_{\mathsf{actual}}|<5\mathsf{mg/dl}\quad(2)$
-and from this we can derive the true conclusion on the user’s glucose levels, which is what we refer to as the biological invariant.
-We stress that this invariant involves the parameter $g_{\mathsf{actual}}$ that is not available in software; it instead ensures a property (which we are able to formally prove) about the real level of glucose in the biological system.
-That statement we prove in the software is actually the following, not (1):
-$|g_{\mathsf{measured}}-g_{\mathsf{predicted}}|<35\mathsf{mg/dl}\quad(3).$
-When we run BioKernel on a real human, we don’t want to assume safety with respect to the measured glucose levels (invariant (1)); we actually want to ensure safety with respect to the real glucose levels in the human that is using the system (invariant (3)).
-Failing to distinguish these cases could be a severe software bug, and would jeopardize the end-to-end safety that we aim to provide by failing to account for CGM error.
-In the rest of the section, we describe the invariants that we enforce in the implementation (including invariant (1) as a simple case) and what additional invariants we prove in our formal verification development (including invariant (3)) below.
-Implementation invariant.
-In the implementation, we check that changes in glucose levels that we observe from CGM readings and the individual’s insulin delivery history are consistent with our theoretical calculations of how much the person’s glucose should change given these same inputs.
+We call this safety check that accounts for changing human physiology a biological invariant.
+In our implementation, we check that changes in glucose levels that we observe from CGM readings and the individual’s insulin delivery history are consistent with our theoretical calculations of how much the person’s glucose should change given these same inputs.
 If their glucose is dropping by more than our theoretical calculations show, then they are more sensitive to insulin then our system believes they should be.
-We set a threshold of 30 mg/dl per hour, for the maximum allowable divergence based on clinical evaluations showing detectability of significant glycemic deviations above this level [37, 34].
+We set a threshold of 30 mg/dl per hour, for the maximum allowable divergence based on clinical evaluations showing detection of significant glycemic deviations above this level (30mg_1, ; 30mg_2, ).
 We take a three-tiered approach to dealing with deviations in our glucose change calculations.
@@
 Third, if they remain at 30 mg/dl or more for more than two hours, we alert the individual and transition the system into a manual mode that requires the user to dose insulin manually.
-Verification.
-In the verification development (described in Section 7), we prove that the biological invariant holds, given error bounds set by the CGM manufacturer.
-First, in the ported Dafny code, we locate the portions which enforce implementation invariants described above, including invariant (1).
-The code includes the two parameters $g_{\mathsf{measured}}$ and $g_{\mathsf{actual}}$, as these are available to the software implementation.
-Second, we model actual glucose by putting error bounds on our CGM readings, in the CGM invariant (2).
-This CGM invariant is inserted as a precondition on all functions that operate on CGM values.
-Third, we track the individual’s actual glucose level using a ghost variable in Dafny.
-Ghost variables are variables that are only available to the prover, these do not exist in the real software system.
-Fourth and lastly, we express as a postcondition the biological invariant (4), which asserts that the absolute difference between the theoretical model glucose and the ground-truth actual glucose cannot exceed a threshold of 35 mg/dl per hour.
-When the code is compiled, the Dafny verifier proves that the biological invariant holds on all inputs to the verified function.
-Evaluation results.
-Our evaluation shows that the biological invariant is effective at keeping humans safe when they face drastic changes in insulin sensitivity (Section 8.3), and that our thresholds of 30 mg/dl per hour and two hours were effective for the human that used GlucOS to manage their T1D without any false positives (Section 9.3).
-Novelty.
-We beleive that having a parameter real biological state of the system is an important modeling decision; and it diverges, for example, from previous work which only models the biological data in simulation [71, 72].
-Traditional software verification techniques like model checking [36] and deductive verification [27] have been explored in the past for hybrid system verification [26] that interact with the human body, but they do not completely account for physiological variability between individuals or model inaccuracies during the insulin delivery process.
-Also, researchers have used real-world invariants for security in the context of autonomous vehicles [18], but these were based on quad-copter physics not human physiology and did not include formal verification of their implementation.
-6 Last line of defense: Humans
+6. Last line of defense: Humans
 In this section, we describe our design and implementation for incorporating humans as a critical component of our multi-layered security system.
 While our algorithmic, system-level, and biological invariant defenses form the primary barriers against attacks, humans serve as the last line of defense, capable of mitigating attacks that manifest as glycemic imbalance.
-When an individual’s glucose levels are out-of-range, they need to intervene to bring them back in range.
+We discuss our most interesting findings around preferences and the architectural iterations we make to our design based on our real-world experiences.
+6.1. Predictive alerts for humans
+When an individual’s glucose levels are not within the range of healthy values, they need to intervene to bring them back in range.
 Low glucose, or hypoglycemia, occurs from too much insulin, and there is no way to remove insulin that the automated insulin delivery system has already delivered.
@@
 Given that people with T1D are already active participants in managing glycemic imbalance, we focus on how we can use this fact to our advantage to create a more comprehensive security system.
-Figure 6: GlucOS predicts hypoglycemia and notifies the individual before it happens so they can proactively prevent it.
+Figure 5. GlucOS predicts hypoglycemia and notifies the user before it happens so they can proactively prevent it.
 We use predictive yet deterministic mobile notifications to alert people when they are likely to experience hypoglycemia before they reach it (Figure 5), as hypoglycemia presents the most immediate health risk and could be an indicator of an attack.
 Alerts for hypoglycemia are not new.
-All CGM software comes with non-predictive hypoglycemia alerts by default [2] and glucose prediction algorithms have been well covered in the literature [65, 70, 11].
+All CGM software comes with non-predictive hypoglycemia alerts by default (dexcom_alerts, ) and glucose prediction algorithms have been well covered in the literature (Josep2020, ; Zhang2021, ; arora2023multivariate, ).
 Our contribution is using predictive alerts before someone reaches hypoglycemia as a security defense.
@@
 Our finding has the potential to have broader implications for security in healthcare systems.
-Often, security systems favor uniformity and consistency [23, 28], but when managing people’s health we found that allowing people to balance alert fatigue and security increased engagement and potentially increased protection, based on each individual’s needs.
-7 Formal verification
-To provide additional assurance that our implementaiton is correct, we port our Swift code to Dafny and use Hoare logic.
-As the system we are verifying is complex, our goal is to locate the pieces and target verification in places where formal methods can have the greatest impact; this is a common reality of applying formal methods at scale, see for example [19].
+Often, security systems favor uniformity and consistency (DHILLON2016656, ; FURNELL20165, ), but when managing people’s health we found that allowing people to balance alert fatigue and security increased engagement and potentially increased protection, based on each individual’s needs.
+6.2. Architectural implications
+One interesting design iteration we made was moving the alerting logic into the BioKernel itself.
+In our first design, consistent with our separation principles, the alerting logic resided in a separate app.
+This alerting app learned the state of the system using our cloud-based event logging system.
+However, one of our participants went on a canoeing trip and was disconnected from the internet, so they could not get alerts.
+Ideally iOS would provide inter-process communication mechanisms to pass data from the BioKernel directly to the alerting app, but iOS does not have the appropriate APIs for this style of communication.
+Thus, we moved the alerting logic into the BioKernel, violating our principle of separation, but enabling our users to get alerts even when they are disconnected from the internet.
+Moving the alerting logic into the BioKernel did have security implications.
+It adds complexity to our trusted computing base, but we carefully manage this tradeoff.
+We minimize the added complexity by using traditional, well-understood algorithms for prediction rather than more complex ML models and reuse existing functionality within the BioKernel when possible to keep the implementation simple.
+This approach allowed us to maintain a high level of confidence in the security and correctness of our alerting system while still providing the necessary functionality.
+7. Formal verification
+To provide additional assurance that our implementation is correct, we port our Swift code to Dafny and use Hoare logic.
+As the system we are verifying is complex, our goal is to locate the pieces and target verification in places where formal methods can have the greatest impact; this is a common reality of applying formal methods at scale, see for example (chong2020code, ).
 Since we are porting the Dafny code to Swift, to reduce the chance of human error, we use a line-by-line port as much as possible and add unit tests for each function to provide confidence that the two implementations are equivalent.
 We port updated code back to Swift manually because Dafny does not include code generation for Swift.
-These efforts paid off: we found and fixed nine bugs in our original implementation (Appendix D); in addition to proving various invariants in the code for our insulin delivery path and security mechanisms.
+These efforts paid off: we found and fixed nine bugs in our original implementation; in addition to proving various invariants in the code for our insulin delivery path and security mechanisms.
 However, the most interesting part was where these bugs were: all of our core algorithms were correct, but we found bugs in seemingly less critical parts of the system, like user settings, which had subtle interactions with the core algorithms.
-For a full list of the invariants that we define and prove, see Appendix E.
-8 Evaluation
-8.1 How effective is GlucOS’s algorithmic security mechanism?
-We evaluate GlucOS’s algorithmic security mechanism on 21 virtual humans from an FDA-approved simulator [7]. We use simulation since this evaluation involves injecting dangerous amounts of insulin.
-To be effective, GlucOS’s algorithmic security mechanism should protect users against untrusted ML algorithms that administer incorrect insulin doses (overdose or underdose), while minimally impeding algorithms that are efficient at managing T1D.
-Accordingly, we first run simulations with a predictive ML model inspired from prior research [52], without incorporating our security mechanism to set a baseline of an algorithm that is efficient at managing T1D.
-Then, in the same simulation scenarios, we introduce an order of magnitude difference (increase and decrease) to the insulin doses computed by the baseline algorithm, without incorporating our security mechanism to quantify the risks of untrusted algorithms.
-We then repeat both sets of simulations while incorporating our security mechanism to quantify the impact of the mechanism on correct and incorrect dosing.
-The goal of T1D management is to maintain glucose levels that are in a safe, target range (70-180 mg/dl) for the majority of the time – an outcome that has been linked to positive long-term health outcomes [53].
-Glucose levels 54-69 mg/dl (level 1 hypoglycemia) or less than 54 mg/dl (level 2 hypoglycemia) put individuals at risk for acute complications like impaired consciousness, seizure, and death, while levels 181-250 mg/dl (level 1 hyperglycemia) or greater than 250 mg/dl (level 2 hyperglycemia) increase risks for long-term vascular damage.
-The American Diabetes Association (ADA) recommends that individuals spend at least 70% of their time in range [6, 24] in order to remain healthy.
-8.1.1 Without algorithmic security mechanism
+Our implementation operates on CGM readings, but in our application of formal methods, our goal is to reason about the individuals actual glucose level, taking into account sensor errors.
+This raises an interesting challenge for formal verification as it requires modeling the true biological state of the user – a parameter that is both unknown at runtime and distinct from measured parameters that are part of the software implementation.
+We prove that the biological invariant holds, given error bounds set by the CGM manufacturer (garg2022accuracy, ).
+To reason about actual glucose levels formally, we use a ghost variable in Dafny.
+Ghost variables are variables that are only available to the prover, these do not exist in the real software system.
+By defining the actual glucose levels using a ghost variable, we can include it in preconditions and postconditions to prove properties about our software relative to actual glucose levels.
+We believe that having a parameter for the real biological state of the system is an important modeling decision; and it diverges, for example, from previous work that only models the biological data in simulation (zhou2021data, ; zhou2023knowsafe, ).
+Traditional software verification techniques like model checking (cps_1, ) and deductive verification (cps_3, ) have been explored in the past for hybrid system verification (cps_2, ) that interact with the human body, but they do not completely account for physiological variability between individuals or model inaccuracies during the insulin delivery process.
+Also, researchers have used real-world invariants for security in the context of autonomous vehicles (choi2018detecting, ), but these were based on quad-copter physics not human physiology and did not include formal verification of their implementation.
+8. Implementation
+In our implementation, the BioKernel is an iOS app that consists of 5.1k lines of code in our trusted computing base.
+It communicates with a Dexcom G6 or G7 CGM and Omnipod DASH insulin pump via Bluetooth low energy, and runs the core closed-loop algorithm.
+For the BioKernel, we use the Swift programming language, the Actor abstraction for data-race-free concurrency, and JSON files stored on disk to capture persistent state.
+Although the BioKernel is a fully featured and stand alone automated insulin delivery system, GlucOS uses an event logging system (Figure 1(c)) that enables separate apps to extend functionality while remaining isolated from the BioKernel.
+9. Evaluation
+In this section, we evaluate GlucOS’s algorithmic and driver security mechanisms. In the next section (Section 10), we evaluate all components of GlucOS in the real world.
+9.1. How effective is GlucOS’s algorithmic security mechanism?
+We evaluate GlucOS’s algorithmic security mechanism on 21 virtual humans from an FDA-approved simulator (simglucose, ).
+GlucOS’s algorithmic security mechanism should protect users against untrusted ML algorithms that administer incorrect insulin doses (overdose or underdose), while minimally impeding algorithms that are efficient at managing T1D.
+We consider extreme cases of attack in our evaluation.
+While subtle attacks are also possible, we focus on extreme attacks since they have the most adverse consequences.
+Concreteny, we first run simulations with a predictive ML model inspired from prior research (mosquera2023enabling, ), without incorporating our security mechanism to set a baseline of an effctive algorithm in managing T1D.
+Then, in the same simulation scenarios, we introduce an order of magnitude increase to the insulin doses computed by the baseline algorithm, without incorporating our security mechanism to quantify the risks of untrusted algorithms.
+We then repeat these experiments with our security mechanism to quantify the impact of the mechanism on correct and incorrect dosing.
+The American Diabetes Association (ADA) recommends that individuals with T1D spend at least 70% of their time in the range (TIR) of 70-180 mg/dl to remain healthy (nathan2014diabetes, ; endocrine_tir, ; journal_glycemic_targets, ).
+Glucose levels less than 70 mg/dl (hypoglycemia) put individuals at risk for acute complications like impaired consciousness, seizure, and death, while levels greater than 180 mg/dl (hyperglycemia) increase risks for long-term vascular damage.
+9.1.1. Without algorithmic security mechanism
 Predictive ML model baseline With the predictive ML model, we followed all ML best practices in carefully tuning hyperparameters and evaluating the model on simulation scenarios that were different from those used in training.
-Figure 7 shows the average proportion of time spent in range (TIR) by virtual humans from different age groups (adults, adolescents and children) when employing the predictive ML model.
+Figure 6 shows the average proportion of time spent in the safe target range (TIR) by virtual humans from different age groups when employing the predictive ML model.
 We see that individuals across all groups spend over 85% of their time in range compared to 77% when using only the reactive safe model.
@@
 Children 77.31% 86.6% 85.42%
-Figure 7: Average proportion of time spent in range (TIR) across 21 virtual humans with a reactive safe model, a predictive ML model, and clamping the predictive ML with the reactive safe model for security.
+Figure 6.
+Average proportion of time spent in range (TIR) across 21 virtual humans with a reactive safe model, a predictive ML model, and clamping the predictive ML with the reactive safe model for security.
 Incorrect ML dosing Since GlucOS is designed to be flexible with any predictive algorithm, it could be subject to attacks from malicious ML algorithms.
-To simulate attacks, we intentionally introduce an order of magnitude difference to the amount of insulin predicted by the baseline predictive ML model.
-In one experiment, we dose one-tenth of the amount of insulin and in another experiment, we dose ten times the amount of insulin computed by the predictive ML model.
-We run these malicious ML algorithms within the simulator on each virtual human on the same scenarios as the previous experiment.
-Figure 8 and Figure 9 show the CGM readings on three randomly chosen virtual humans on a particular scenario for the experiments where we dose less and dose more respectively.
-When dosing less, we see that all of them spend most of their time above Level 1 hyperglycemia.
-When dosing more, all individuals fully go below Level 2 hypoglycemia, which would have killed them.
-Figure 8: Variation in CGM readings for three virtual humans when running a predictive ML algorithm that intentionally doses one-tenth the amount of the required insulin.
-The CGM readings staying mostly above the Level 1 hyperglycemic range.
-Figure 9: Variation in CGM readings for three virtual humans when running a predictive ML algorithm that intentionally doses ten times the amount of the required insulin.
-The CGM readings staying below the Level 2 hypoglycemic range would have likely been fatal to all three individuals.
-Across all virtual humans, we see an average drop in TIR from 88.44% with correct dosing to 23.32% when dosing less and an average drop in TIR from 88.44% to 0% when dosing more.
-Correspondingly, when dosing less, we observe an average increase in the time in hyperglycemia from 3.77% with correct dosing to 65.05%.
-Similarly, when dosing more, we observe an average increase in the time in hypoglycemia from 0% to 100%.
-These results further motivate the need for security mechanisms when providing individuals with the flexibility to use any algorithm to manage T1D.
-8.1.2 With algorithmic security mechanism
+In this section, we discuss an experiment with a malicious algorithm that intentionally doses too much insulin.
+Concretely, our malicious algorithm doses ten times the amount of insulin computed by the previously described predictive ML model.
+We ran this algorithm within the simulator on the same virtual human on the same scenarios as the previous experiment.
+Figure 7 shows the CGM readings on three randomly chosen virtual humans on a particular scenario when we dose more.
+The CGM readings of all 3 individuals fully go below the hypogycemic bound, which would have killed them in the real-world.
+Figure 7.
+Variation in CGM readings for three virtual humans when running a predictive ML algorithm that intentionally doses ten times the amount of the required insulin.
+The CGM readings staying below the hypoglycemic bound would have likely been fatal to all three individuals.
+Across all virtual humans, we see an average drop in TIR from 88.44% to 0% when dosing too much insulin.
+Correspondingly, we observe an average increase in the time spent in hypoglycemia from 0% to 100%.
+We also observed similar results with experiments where we intentionally underdosed insulin.
+In these experiments, all virtual humans spend most of their time above the hyperglycemic bound which wouldn’t have been fatal but would have led to severe health consequences in the real world.
+These results motivate the need for our algorithmic security mechanism.
+9.1.2. With algorithmic security mechanism
 Impact on incorrect dosing.
-We repeat the experiments from the previous section with the same scenarios.
-We run the same models for incorrect dosing, but also run our reactive safe model (Section 3) to set bounds on the amount of insulin the malicious algorithms can inject.
-Figure 10 and Figure 11 show CGM readings on the same three virtual humans on the same scenarios from the experiment in the previous section.
-With both types of incorrect dosing, we see that all of them spend over 70% of their time in range.
-These results show that GlucOS protects individuals even in presence of malicious incorrect dosing.
-On average, across all virtual humans, we only see a drop in TIR from 88.44% with correct dosing to 73.89% when employing safety while dosing less and a drop in TIR from 88.44% to 78.22% when dosing more.
-While the times spent in range is less than that of correct dosing, the impact of incorrect dosing is significantly dampened with individuals spending the recommended amount of time in range, despite facing an active attack.
-Figure 10: Variation in CGM readings for three virtual individuals when clamping an algorithm that intentionally tries to dose one-tenth the amount of the required insulin.
-On average, the CGM readings stay in range over 70% of the time.
-Figure 11: Variation in CGM readings for three virtual humans when clamping an algorithm that intentionally tries to dose ten times the amount of the required insulin.
+We repeat experiments from the previous section with the same scenarios.
+We run the same malicous model to inject too much insulin, but also run our reactive safe model (Section 3) to bound insulin injection.
+Figure 8 show CGM readings on the same three virtual humans on the same scenarios as the experiment in the previous section.
+We see that all of them spend over 70% of their time in range, showing that GlucOS protects individuals even in presence of malicious overdosing.
+On average, across all virtual humans, we only see a drop in TIR from 88.44% with correct dosing to 78.22% when employing safety while overdosing.
+We also report similar results on experiments with a malicious algorithm that underdoses insulin.
+While the times spent in range are less than those of correct dosing, the impact of incorrect dosing is significantly dampened with individuals spending the recommended amount of time in range, despite facing active attacks.
+Figure 8.
+Variation in CGM readings for three virtual humans when clamping an algorithm that intentionally tries to dose ten times the amount of the required insulin.
 On average, CGM readings stay in range over 70% of the time.
@@
 We evaluate the impact of GlucOS’s algorithmic security mechanism on benign algorithms that are effective in managing T1D such as the baseline predictive ML model.
-We repeat the experiment of running the predictive ML model on all virtual humans in the same random scenarios while maintaining the same parameters for the clamps and reactive model used to evaluate incorrect dosing.
+We repeat the experiment of running the predictive ML model on all virtual humans in the same scenarios with the same parameters for the clamps and reactive model used to evaluate incorrect dosing.
 Comparing data in Figure 6 we see that there is only a slight drop in the average proportion of time spent in range by virtual humans across all age groups when employing GlucOS’s algorithmic security mechanism.
@@
 This shows that GlucOS’s algorithmic security guarantees do not come at the cost of effective T1D management from predictive ML.
-Figure 12: We attached an Omnipod DASH insulin pump to a stuffed animal and modified the OmniBLE pump driver to intentionally add, drop and modify insulin commands.
+Figure 9.
+We attached an Omnipod DASH insulin pump to a stuffed animal and modified the OmniBLE pump driver to intentionally add, drop and modify insulin commands.
 Our driver security mechanism detected all manipulations.
-8.2 Does GlucOS ensure pump operation while overcoming malicious pump drivers?
-We first evaluate GlucOS’s driver security mechanism in terms of its impact on benign pump drivers that accurately relay pump communications.
-We then evaluate the impact on detecting additional, modified, or dropped communication from a malicious pump driver.
+9.2. Does GlucOS ensure pump operation while overcoming malicious pump drivers?
 Our mechanism should ensure that all communications go through with a benign pump driver and should catch all manipulated communication with a malicious pump driver.
-We do not use the simulator for this evaluation since it does not incorporate a pump driver.
 Impact on malicious pump drivers.
-As shown in Figure 12, we attached an Omnipod DASH insulin pump to a stuffed animal.
-We then modified the source code of the OmniBLE pump driver to intentionally add, drop and modify insulin commands.
-We detected all manipulated pump commands showing that our driver security mechanism has 0 false negatives in detecting manipulated commands.
+We attached an Omnipod DASH insulin pump to a stuffed animal (Figure 9).
+We then modified the OmniBLE pump driver to intentionally add, drop and modify insulin commands.
+We detected all manipulated commands thereby showing that our driver security mechanism has 0 false negatives in detecting them.
 Impact on benign pump drivers.
-For one week, we logged messages from the BioKernel to unmodified OmniBLE as well as messages sent from OmniBLE to the pump on an individual running GlucOS (See Section 9).
-We report 100% consistency between the commands sent to OmniBLE and commands issued from OmniBLE showing that GlucOS’s driver security mechanisms do not incur any false positives.
+For one week, we logged messages passed between the BioKernel, the unmodified OmniBLE pump driver and the pump on an individual running GlucOS (Section 10).
+We report 100% consistency between the commands sent to OmniBLE and commands issued from OmniBLE showing that GlucOS’s driver security mechanism does not incur any false positives.
 These results show that GlucOS can ensure pump operation while withstanding attacks from malicious drivers.
-Age Group TIR without
-biological invariant
-Adults 49.93%
-Adolescents 55.83%
-Children 67.41%
-Figure 13: Average proportion of time spent in range across 21 virtual humans when fluctuating their insulin sensitivity.
-The low values for TIR stress the importance of an end-to-end security mechanism, i.e., the biological invariant.
-8.3 Is the biological invariant important?
-We run experiments on the simulator to assess the risks of not considering our end-to-end security mechanism, i.e., the biological invariant.
-Figure 13 shows the average TIR across age groups.
-Across all virtual humans, we record an average TIR of 57.56%, even with our algorithmic security mechanism, when fluctuating insulin sensitivity.
-This corresponds to an average drop of 29.28% as a result of not considering the biological invariant.
-These results emphasize the importance of incorporating an end-to-end invariant and undermines results obtained from pure simulation.
-We present case-studies involving the biological invariant with a real user using GlucOS in Section 9.3.
-9 Using GlucOS in the real world
+10. Using GlucOS in the real world
 In this section, we report on our experiences working with seven individuals using GlucOS to help manage their T1D.
@@
 One individual, who we will call Bob, has been using the full GlucOS system since November 2023 for insulin delivery and predictive alerting.
+10.1. Ethical considerations
+In this section, we outline the steps we take to ensure that we uphold high ethical standards and ensure that the people using our system, are safe.
 Our user study for the six students was approved by our university’s IRB, and Bob’s use GlucOS was determined by our university’s IRB to be “self evaluation” and exempt from IRB.
-In our implementation, the BioKernel is an iOS app that consists of 5.1k lines of code in our trusted computing base.
-It communicates with a Dexcom G6 or G7 CGM and Omnipod DASH insulin pump via Bluetooth low energy, and runs the core closed-loop algorithm.
-For the BioKernel, we use the Swift programming language, the Actor abstraction for data-race-free concurrency, and JSON files stored on disk to capture persistent state.
-Although the BioKernel is a fully featured and stand alone automated insulin delivery system, GlucOS uses an event logging system that enables separate apps to implement complementary features while remaining isolated from the BioKernel.
-9.1 Ethical considerations
-In this section, we outline the steps we take to ensure that we uphold high ethical standards and ensure that the people using our system, are safe.
 One of the co-authors of this paper is a board-certified Endocrinologist, who specializes in T1D.
-They are both a clinician with an active practice seeing patients living with T1D and a researcher who focuses on research on T1D.
-They designed the safety protocol, which defines the testing we conduct to ensure that the individuals using our system are safe, and the specific criteria we use to stop the study, if needed.
-Appendix A includes a more detailed discussion of our ethical considerations and safety protocol.
-9.2 Is an ML-based closed-loop system safe and effective in practice?
+They designed the safety protocol, which defines the specific criteria we use to stop the study, if needed.
+The Appendix will include a more detailed discussion of our ethics and safety protocols.
+10.2. Is an ML-based closed-loop system safe and effective in practice?
 Bob has been using GlucOS in a closed loop since November 23rd, 2023.
@@
 We report on one week’s worth of data from the predictive ML model starting on January 31st ending on February 6th, 2024.
-Glucose Range Reactive Reactive ML
-range (mg/dl) label 67 days last wk. 1 wk
-<54 L2 Hypo $0$%
-54-69 L1 Hypo 0.59% 0.48% 0%
-70-180 In range 96.35% 91.90% 97.22%
-181-250 L1 Hyper 2.79% 7.62% 2.78%
->250 L2 Hyper 0.18% 0% 0%
-Figure 14: The percent of time spent in different glucose ranges for the reactive safe model and the predictive ML model. Bold values represent the best results for each range.
-Figure 14 shows the amount of time Bob spent in different glucose level ranges over the duration of our study.
-Data from the reactive safe model covers 67 days, followed by one week of data from our predictive ML model.
-We also show data for the last week of using the reactive safe model to show trends over time.
 Overall, Bob had the best results when using ML.
+His time spent in range increased from 96.3% to 97.2%, and notably he spent no time in hypoglycemia (below 70 mg/dl), down from 0.67% when not using ML.
 Next, we measure how often the predictive ML model was responsible for programming the pump vs our reactive safe model.
@@
 We believe that this distribution represents a suitable balance for providing the ML with enough flexibility to improve outcomes, while having tight enough bounds to limit potential damage from bad or malicious ML predictions.
-9.3 Do biological invariant violations trigger in practice?
+10.3. Do biological invariant violations trigger in practice?
 To determine if biological invariant violations trigger in practice, we review 90 days of data for Bob, from May 29th, 2024 to August 27th, 2024.
@@
 We reviewed one weeks worth of violations with Bob manually to determine the cause. During the most recent week of data, Bob had five sequences of 30 minutes or longer and all were caused by exercise.
-9.4 Are predictive alerts useful for security?
+10.4. Are predictive alerts useful for security?
 To measure predictive alerts as a security mechanism (Section 6), we recruited six students who live with T1D and had them use the GlucOS predictive alerts to complement their default CGM alerts.
@@
 For the two month study, our participants decreased their average time in hypoglycemia from 2.7% down to 1.6%, and all individuals experienced a decrease in the amount of time spent in hypoglycemia.
-For context, participants in a state-of-the-art automated insulin delivery clinical trial experienced a decrease of time in hypoglycemia of 0.9% [15], showing that our 1.1% decrease is substantial.
+For context, participants in a state-of-the-art automated insulin delivery clinical trial experienced a decrease of time in hypoglycemia of 0.9% (brown2019six, ), showing that our 1.1% decrease is substantial.
 Overall, our results suggest that predictive alerts can be a useful last line of defense for trustworthy automated insulin delivery systems.
-10 Related work
+10.5. How does the BioKernel perform?
+We look at two aspects of system performance when evaluating our real-world use of GlucOS.
+First, we measure how long our closed-loop algorithm takes to run and second, we count how often the closed-loop algorithm runs successfully end-to-end.
+We use Bob’s iPhone 14 to collect all of the data we show here and these numbers are averages across the entire eleven month period.
+On average, the BioKernel spends 4.3s running the closed loop algorithm, where almost half of that time is dominated by CGM and pump communications.
+Of the components within the closed-loop algorithm, the ML predictions take 2.2s on average and the safety logic takes 19ms (0.019s).
+Since the system runs in the background, we believe that iOS runs the device with reduced computational capacity to save power.
+But overall, 4.3s of wall-clock time for every five minutes our closed-loop algorithm runs with only a tiny fraction of that coming from our safety logic represents practical system performance.
+For successful closed-loop end-to-end execution, over the evaluation period the closed-loop algorithm runs successfully 96.5% of the time, with 2.1% of the runs failing due to pump communication, which is our largest class of failure.
+11. Related work
 In addition to the work already discussed in this paper, there are several other related works.
-Previous research has looked at the security of implanted medical devices in general [32, 16, 60], in addition to looking at insulin pumps in particular [39, 56], with more recent work looking at providing improved security [45, 9].
-Also, recent work has looked at applying formal methods to insulin pumps for high assurance [55].
+The complexity of automated insulin delivery systems and the fact that they make real-time healthcare decisions, makes their security paramount.
+Current research on security for these systems focuses primarily on evaluating algorithms using simulations (zhou2021data, ; zhou2023knowsafe, ).
+We argue that this approach falls short of addressing the full spectrum of challenges inherent in building trustworthy systems that operate within real-world constraints.
+Real-world systems must cope with runtime platform limitations, changes in human physiology, and the necessity of human interaction with the system.
+Plus, algorithms and simulations model attacks abstractly through their glucose response, missing out on details like which components of the system are likely to be compromised or how the system architecture can prevent damage from these potentially vulnerable components.
+Our experience shows that when faced with these practical challenges, algorithms designed in simulations often prove insufficient or inapplicable.
+These real-world complexities require a fundamentally different approach to designing security for automated insulin delivery systems.
+To address these challenges, GlucOS moves beyond simulation-based designs and tackles the challenges of real-world systems.
+Previous research has looked at the security of implanted medical devices in general (halperin2008security, ; burleson2012design, ; rushanan2014sok, ), in addition to looking at insulin pumps in particular (li2011hijacking, ; paul2011review, ), with more recent work looking at providing improved security (marin2016feasibility, ; ahmad2018securing, ).
+Also, recent work has looked at applying formal methods to insulin pumps for high assurance (panda2021secure, ).
 These works focus on devices and their communication channels.
 In contrast, with GlucOS we assume that these devices are correct and secure and focus our efforts on the software we use to run the automated insulin delivery system.
-Researchers have started to look at combining untrustworthy ML with more traditional forms of automated decision making, like Markov Decision Processes, from a theoretical perspective [41, 50, 57, 20, 40].
-These theoretical works focus on providing algorithms to prove desirable properties about a combined ML / Markov system, like robustness and stability.
-In contrast, we take advantage of domain knowledge and formulate a simple and practical system for one specific problem: automated insulin delivery.
-11 Conclusion
+12. Conclusion
 The design of GlucOS, while grounded in established OS and security principles, incorporates several non-obvious elements that emerged from our deep engagement with the problem domain and real-world testing.
@@
 This interdisciplinary research drew on deep expertise from computer security, medicine, and ethics to create a practical system.
-Our work improves the security of automated insulin delivery systems and also provides a framework for thinking about security in other critical healthcare technologies.
 Our exploration is an extreme example of personalized, distributed, and automated healthcare using trustworthy computer systems.
-Appendix A Ethical and safety considerations
-One individual, who we will call Bob, used GlucOS for 9 months.
-Six university students also used the predictive alerts from GlucOS for 2 months.
-We put in place a protocol where we would stop our study and ask people to stop using our system and revert back to their previous management regime.
-In our protocol, we have formal in-person meetings with Bob every week to check his data and make sure that he is not incurring risk as a result of using GlucOS, and weekly surveys for the six students, plus monthly meetings.
-We would have stopped our study if any of the following conditions occurred:
-  • Excessive hypoglycemia.
-We define excessive hypoglycemia as having spent 8% or more of time during a week with CGM readings below 70 mg/dl or more than 1% below 54 mg/dl.
-These amounts of time spent in hypoglycemia would have represented a significant increase in time spent in hypoglycemia for participants.
-  • Insufficient CGM data. If they had less than 70% of the time during a week with CGM readings, then the lack of CGM data would represent a decrease for the participants.
-  • Insufficient closed-loop runs.
-We designed GlucOS to run closed-loop algorithms every five minutes, or 288 times per day.
-If Bob had a day where our closed-loop algorithm ran successfully less than 200 times, it would represent a fundamental flaw in the system.
-  • Serious complications due to diabetes.
-Serious complications include any hospitalization for diabetes related issues, severe hypoglycemia where a participant was unable to recover from hypoglycemia themselves and had to get help from someone else, diabetic ketoacidosis, or seizure.
-  • New medical diagnosis requiring significant attention. If anyone had received a new medical diagnosis during the study, the study would have had the potential to be a distraction for participants.
-  • New mental health diagnosis. If anyone were struggling with mental health during the study, the study would have the potential to be a distraction.
-In addition to our safety protocol, we followed best practices for ethical research to ensure that we conducted our study ethically.
-  • IRB.
-Our study went through our university’s IRB process, where the study including six university students who used GlucOS’s alerting software was approved, and our case study on Bob was determined to be “self evaluation” and exempt from IRB.
-  • Informed consent. All participants went through a written informed consent process.
-  • Data management.
-GlucOS does not store any personally identifiable information and uses access control on our Google Cloud Platform infrastructure to ensure that only the authors of the paper can access this anonymized data.
-  • Risk/benefit analysis.
-All participants are already living with T1D and use CGMs and inject insulin, so they already take on the risks inherent in managing T1D.
-The benefits include improved control over their T1D.
-There are risks in using experimental software, but all participants continued to rely on their CGM alerts as a last line of defense.
-  • Participant selection. We selected participants who were university students living with T1D.
-  • Adverse event reporting. Had any participants experienced adverse events during the study, we would have reported it to our IRB for guidance and addressed it according to our safety protocol.
-  • Communication. All surveys were done through our university’s medical school infrastructure for private communication with individuals.
-Since Bob used the full GlucOS software, including for insulin delivery, we included several other safety mechanisms:
-  • Bob increased the frequency of his visits to his Endocrinologist while using the GlucOS software. He went from yearly appointments to quarterly appointments.
-  • During the study, twice Bob got lab work based on blood tests to provide an independent measure of his overall metabolic health.
-This lab work helped to ensure that he was in good health and that there were no hidden negative impacts from his use of GlucOS.
-Overall, GlucOS had a large positive impact on the individuals who used it.
-Bob’s lab-based A1C test (a blood test that measures average glucose levels over time), which includes his time using GlucOS, was 5.8%.
-That is nearly non-diabetic levels of control – healthy individuals will have an A1C of 5.6% or lower.
-All of the students decreased the amount of time spent in hypoglycemia.
-These results are profound for the seven people using GlucOS, and it had a positive impact on their health and quality of life.
-Finally, Bob was a willing and enthusiastic user of the GlucOS system.
-To be clear, Bob is not a graduate student who was forced to use GlucOS so that his advisor could publish a paper.
-Rather, Bob struggled with the cognitive load of T1D management and wanted an automated insulin delivery system to offload part of the burden.
-However, he was unwilling to use any of the other automated insulin delivery systems out there due to concerns over their security.
-Bob found that GlucOS, with its focus on security, correctness, and simplicity, met his needs.
-Appendix B List of pump setup and UX communication messages
-Message Message Message
-name type description
-setupPod Setup Setup a new pump
-assignAddress Setup Pair pump with phone
-deactivatePod Setup Teardown pump
-acknowledgeAlert UX Acknowledge alert from pump
-configureAlerts UX Configure alerts from pump
-beepConfig UX Configure pump beeps
-Figure 15: Types of messages exchanged between a pump driver and pump pertaining to setup/UX.
-We list insulin message types pertaining to pump setup and UX in Figure 15.
-Figure 16: GlucOS’s CGM state machine to handle missing readings from CGMs.
-Appendix C Runtime failure management
-Figure 17: GlucOS’s state machine to transition users to the safest manual dosing state when losing trust in insulin delivery (i.e., with a buggy or malicious pump driver) or when the biological invariant is not satisfied.
-In this section, we describe GlucOS’s state machine that combines its security components, i.e., algorithmic security (Section 3), driver security (Section 4) and end-to-end security (Section 5) together to protect users in light of their corresponding failures.
-GlucOS’s state machine consists of three operating states: (1) ML with safety operating in closed loop, (2) reactive safe model operating in closed loop and (3) manual dosing in closed loop.
-ML provides the tightest level of control in managing glucose but poses the largest immediate threat from excess insulin since it preemptively injects insulin by anticipating rise in glucose.
-The reactive safe model provides slightly lesser control, but poses milder immediate threat since it only injects insulin to account for measured excess glucose.
-Manual dosing offers the least amount of control but poses no threat since it only injects the minimum insulin needed for sustenance.
-GlucOS predominantly operates in the state running ML (with safety) in closed loop.
-We transition to the safest manual dosing state when we do not trust insulin delivery (i.e., with a buggy or malicious pump driver).
-We also transition to the safest manual operation when the biological invariant is not satisfied to ensure that we do not risk the possibility of removing too much glucose from our insulin doses.
-If we end up in the manual dosing state, we enforce that only manual user intervention can transition back to other states.
-We visualize these state transitions in Figure 17.
-Inability to retrieve readings from a CGM also count as runtime failures. However, these failures are easier to handle.
-Since ML needs a stream of CGM readings as input, we transition from ML to the reactive safe model if we don’t receive an input from the CGM.
-As long as the number of missing CGM readings is low (say, within 30 minutes), we can interpolate to account for the missing readings to operate the reactive safe model.
-Once we have too many missing CGM readings that make us lose confidence in interpolation, we transition to manual dosing.
-From manual dosing, we transition back to the reactive safe model on receiving a CGM reading.
-Lastly, from the reactive safe model, we transition back to ML once we have enough CGM readings that are required by the ML model.
-Figure 16 summarizes these transitions.
-Bug Type Description
-Division by Zero Error Potential division by zero error when calculating insulin dosage if the user entered the same start and end time for a basal rate, resulting in a duration of zero.
-Negative Value Handling The insulin dosing calculation for microbolusing and basal rate erroneously allowed negative values.
-Out-of-Range Value Handling The microbolusing function failed to handle cases where one of its calculation components was out of the expected range or resulted in a negative value.
-Violation of Intended Postcondition The function responsible for setting the insulin delivery duration violated its intended postcondition by allowing the end date to be earlier than the start date.
-Violation of Intended Postcondition The function responsible for calculating the total insulin dose violated its intended postcondition by returning an incorrect upper bound value under certain conditions.
-Violation of Intended Postcondition The function responsible for updating the user’s insulin dosage violated its intended postcondition by failing to validate the input parameters correctly.
-Violation of Intended Postcondition The function responsible for calculation of ‘unitsDelivered‘ needed additional postconditions accurately reflect the amount of insulin delivered over the time gap.
-Lack of Error Handling Additional error handling was added to insulin delivered over a time segment function to gracefully catch undesirable state.
-Lack of Error Handling Additional error handling was added to microbolusing function to gracefully catch undesirable states.
-Figure 18: Report of bugs fixed by formally verifying GlucOS.
-Appendix D Did formal verification help with correctness?
-During the formal verification process of our secure insulin delivery system, we uncovered various bugs and inconsistencies that could have compromised the safety and reliability of the system in niche scenarios.
-We identified a total of 9 bugs across different categories.
-3 of these bugs were related to invalid ranges for variables such as negative value returns, domain-specific out of range values, potential division by zero errors etc.
-in the insulin dosing calculations which were resolved by adding appropriate checks and error handling mechanisms to prevent such operations and ensure graceful handling of these cases.
-Additionally, we found 4 instances where the implemented functions violated their intended postconditions along the insulin delivery path, which were addressed by thoroughly reviewing and modifying the implementations to correctly satisfy the postconditions, preserving the intended behavior and meeting our claims for critical funcrions.
-Moreover, our analysis uncovered 2 situations where certain functions lacked proper error handling mechanisms, potentially leading to undefined behavior or system failures.
-By addressing these bugs and incorporating formal verification throughout the development lifecycle, we significantly enhanced the robustness, reliability, and safety of our secure insulin delivery system.
-Appendix E Formal verification: invariants and bugs
-In this section, we list the invariants used to formally verify critical functions of GlucOS. Figure 18 lists the bugs we found and fixed during formal verification.
-E.1 Invariants
-The MicroBolusing function ensures that micro boluses are always administered at safe time intervals while making sure that they are actually needed (the glucose level of the person is in fact high).
-  • Ensure no micro bolus in the last 4.2 minutes.
-  • Ensure glucose is at least 20 mg/dl above the target glucose level.
-  • Ensure predicted glucose is greater than the current glucose level minus 2.
-  • Ensure micro bolus amount is within the allowed range.
-  • Ensure bolus amount is rounded correctly if pumpManager is available.
-We assume that the BioKernel is the only software issuing bolus commands.
-The GuardRails function is responsible for ensuring a safe basal rate value, mindful of the user’s glucose levels. It enforces several invariants:
-  • Ensure the result of rounding newBasalRateRaw to the supported basal rate is within the allowable range.
-  • Ensure the new basal rate does not exceed the maximum basal rate in the settings.
-  • Ensure the new basal rate is not negative.
-  • Ensure that if either the current glucose level or the predicted glucose level falls below or equals the shut-off glucose threshold, the new basal rate is set to 0.0.
-The function operates under the assumption that all paths must go through the guard rails, with a specific evaluation comparing with the pump invariant.
-Additionally, it assumes the integrity of settings, trusting the underlying OS and file system and storage stack.
-The insulinDeliveredForSegment function ensures that intersectionStart is not greater than intersectionEnd when computing the intersection between self.startDate and self.endDate with startDate and endDate.
-  • Verifies that intersectionDuration is non-negative and does not exceed doseDuration, ensuring it accurately reflects the overlapping time duration.
-  • Checks that the units of insulin (units) used for calculation are either deliveredUnits or programmedUnits, ensuring consistency in how the insulin amount is derived.
-  • Verifies that the computation (units * intersectionDuration / doseDuration) accurately represents the amount of insulin delivered during the intersection period.
-The function assumes that the input parameters (self.startDate, self.endDate, startDate, endDate, units, and doseDuration) are valid and within the expected ranges.
-It is also assumed that the underlying data structures and calculations are accurate and consistent throughout the system.
-The CreateBasalDose function requires a time gap greater than 1 second to proceed and return a valid DoseEntry.
-  • Ensures that the calculation of basalRatePerSecond from basalRate accurately represents the rate of insulin delivery per second.
-  • Ensures that the calculation of unitsDelivered accurately reflects the amount of insulin delivered over the time gap.
-  • The DoseEntry constructed should have consistent attributes and adhere to the specified type, units (unitsPerHour), and mutability (false).
-The function assumes that the input parameters (basalRate, startDate, and endDate) are valid and within the expected ranges.
-It also assumes that the underlying time calculations and conversions between units are accurate and consistent.
-The inferBasalDoses function ensures that basalDoses contains only DoseEntry instances where type is .tempBasal, .resume, or .suspend, sorted in ascending order by startDate.
-  • InsulinType should be determined correctly based on the last .tempBasal or .bolus type dose in doses, defaulting to .humalog if none are found.
-  • Each inferred basal dose added to inferredBasalDoses must be created using the createBasalDose function with valid parameters and added in the correct chronological order.
-  • If the last dose in basalDoses is not of type .suspend, a valid basal dose must be inferred from its endDate to at.
-The function assumes that the input parameters (doses, basalDoses, at, and pumpRecordsBasalProfileStartEvents) are valid and within the expected ranges.
-It also assumes the correctness of the underlying data structures and the createBasalDose function.
-The insulinOnBoard function ensures that doses contains unique and valid DoseEntry instances filtered up to at.
-  • The total iob should accurately represent the cumulative insulin on board from all valid dose entries in doses.
-  • basalDoses should contain inferred basal doses that accurately reflect insulin on board contributions when pumpRecordsBasalProfileStartEvents is false.
-  • The final return value of insulinOnBoard should be the sum of iob and basalIob, representing the total insulin on board at at.
-The function assumes that the input parameters (doses, at, and pumpRecordsBasalProfileStartEvents) are valid and within the expected ranges.
-It also assumes the correctness of the underlying data structures and the inferBasalDoses function.
-The ActiveBolus function ensures that data is read consistently from disk before proceeding with any operations.
-  • Verifies that DeduplicatedDoses(events: eventLog, at: at) correctly filters out duplicates and retains only relevant dose entries up to at.
-  • Ensures that doses.filter accurately selects bolus entries that were active at the specified at.
-  • Verifies that doses.last correctly returns the last active bolus entry, ensuring it is non-null if a bolus exists within the specified criteria.
-The function assumes that the input parameters (eventLog and at) are valid and within the expected ranges.
-It also assumes the correctness of the underlying data structures and the DeduplicatedDoses function, as well as the consistency and reliability of disk read operations.
-The TempBasal (Safety Clamps) function ensures that the events array includes only events from safetyStates that fall within the time range [start - duration, at).
-  • Verifies that the historicalMlInsulin value represents the total units of insulin delivered by the machine learning system over the specified time horizon.
-  • Ensures that the machine learning and safety temp basal rates are correctly converted to units of insulin based on the given duration.
-  • Ensures that the upper and lower bounds for insulin units are correctly adjusted based on historicalMlInsulin.
-  • Ensures that the difference between machine learning and safety temp basal units is correctly clamped within the calculated bounds.
-  • The clamped delta units should be converted back to a temp basal rate and added to the safety temp basal rate.
-  • The returned SafetyTempBasal object should correctly encapsulate the adjusted temp basal rate and historical machine learning insulin.
-The function assumes that the input parameters (safetyStates, start, duration, at, mlTempBasal, safetyTempBasal, and lastHistoricalMlInsulin) are valid and within the expected ranges.
-It also assumes the correctness of the underlying data structures and the conversion functions between units, as well as the proper configuration and appropriateness of the safety clamp parameters and thresholds for the user.

### v4

@@
 Fundamentally, this paper is about a system we designed, implemented, and deployed on real humans and the lessons learned from our experiences.
-GlucOS combines algorithmic security, driver security, and end-to-end verification to protect against malicious ML models, vulnerable pump drivers, and drastic changes in human physiology.
+GlucOS introduces a novel architecture that allows users to personalize diabetes management using any predictive model (including ML) for insulin dosing while simultaneously protecting them against malicious models.
+We also introduce a novel holistic security mechanism that adapts to unprecedented changes to human physiology.
 We use formal methods to prove correctness of critical components and incorporate humans as part of our defensive strategy.
 Our evaluation includes both a real-world deployment with seven individuals and results from simulation to show that our techniques generalize.
+We highlight that our results are not from a lab study, with people using GlucOS to manage Type 1 Diabetes in their daily lives.
 Our results show that GlucOS maintains safety and improves glucose control even under attack conditions.
 This work demonstrates the potential for secure, personalized, automated healthcare systems.
-Our source code is open source.
+Our entire source code is available at this link.
 1. Introduction
@@
 To compensate, they inject synthetic insulin.
-Computer systems, called automated insulin delivery systems, exist that use subcutaneous sensors to continuously monitor glucose concentrations (glucose is the body’s primary energy source) and inject insulin automatically using an insulin pump device (Figure 1).
+Mobile applications, called automated insulin delivery systems, use subcutaneous sensors to continuously monitor glucose concentrations (glucose is the body’s primary energy source) and regulate glucose by automatically injecting insulin via an insulin pump (Figure 1).
 However, insulin is a dangerous hormone as too much insulin can kill people in a matter of hours (cryer2012severe, ) and too little insulin can kill people in a matter of days (benoit2018trends, ).
 For automated insulin delivery systems, the key challenge is to maintain a balance: provide enough insulin to prevent dangerously high glucose levels while avoiding excessive insulin that could lead to life-threateningly low glucose levels.
-In this context, security means ensuring integrity for insulin dosing against malicious attacks or inadvertant errors.
-Several commercial (beta_bionics, ; tandem_iq, ; omnipod, ) and open-source (loop, ; openaps, ) automated insulin delivery systems exist today, but none of them consider security as a primary design constraint.
+In this context, security means ensuring integrity for insulin dosing against malicious attacks or inadvertent errors.
 (a) Devices for automated insulin delivery. This figure is from breakthrought1d.org.
@@
 Figure 1. Overview of automated insulin delivery devices, closed-loop architecture, and GlucOS.
-Current open-source automated insulin delivery systems are complex.
+Several commercial (beta_bionics, ; tandem_iq, ; omnipod, ) and open-source (loop, ; openaps, ) automated insulin delivery systems exist today, but none of them consider security as a primary design constraint.
+The complexity of these systems makes it difficult to reason about their correctness, which is crucial for security.
 For example, the OpenAPS (openaps, ) core function for calculating insulin doses consists of 1192 lines of Javascript code, 63 input and configuration parameters, and 90 branch statements.
-All of this logic influences the ultimate dosing decision, making reasoning about its correctness, both manually and automatically, challenging.
-Automated insulin delivery systems that use the OpenAPS algorithm embed it into native apps using a WebView for Javascript interpretation, adding to the overall complexity and increasing the attack surface.
-In contrast, our implementation of this same function is 25 lines of native and formally verified Swift code.
-None of the open-source automated insulin delivery systems uses formal methods, can withstand attacks on components, or has mechanisms or policies for security.
+Furthermore, automated insulin delivery systems that use the OpenAPS algorithm embed it into native apps using a WebView for Javascript interpretation, increasing the overall complexity and attack surface.
+Vulnerabilities or attacks on such systems can be severe and fatal.
 In this paper, we take on the challenge of building the first trustworthy automated insulin delivery system, called GlucOS.
-Our design includes: (1) an architecture where we apply separation principles for isolated and simple components, (2) a novel security mechanism and policy to enable secure ML-driven insulin delivery, (3) the ability to withstand attacks on insulin pump driver software, and (4) the application of formal methods to prove correct our insulin delivery path, the most sensitive part of the system.
-Despite these efforts, we identify scenarios in automated insulin delivery that require users to take action for security.
+Our design includes: (1) an architecture where we apply separation principles for isolated and simple components, (2) a novel security mechanism and policy to enable secure insulin delivery, (3) a mechanism to account for unprecedented changes in human physiology, and (4) the application of formal methods to prove correct the implementation of our insulin delivery path, the most sensitive part of the system.
+Despite these efforts, we identify scenarios that require users to take action for security.
 Thus, we incorporate humans as part of our defensive strategy, providing them with agency and trust to take the required, corrective actions.
+In contrast to existing systems, GlucOS enables the use of any algorithm (including ML) to calculate insulin doses, while ensuring security by moderating insulin doses using 25 lines of formally verified Swift code.
 Our approach is distinguished by its holistic consideration of the entire problem of automated insulin delivery.
 We provide clear security boundaries, mechanisms and policies to mitigate attacks on vulnerable components, and formal methods to provide assurances that our implementation is correct.
-Automated insulin delivery systems are different from traditional applications that the systems and security communities focus on, but they are important.
-Research on operating systems or web browsers can list scores of CVEs from current systems and show how their system improves security by eliminating the impact of some or all of these vulnerabilities by virtue of their design.
-Automated insulin delivery systems do not have a long list of CVEs that we can reference, although there are a few (cve1, ; cve2, ; cve3, ; cve4, ).
-However, our position is that we need to make sure that these systems never get CVEs, because if they do, security vulnerabilities can kill people.
-Thus, we focus on building a new system, using security first principles and formal methods, to avoid security vulnerabilities and mitigate their damage from the start.
+Unlike prior research that exclusively focuses on simulation for evaluation (zhou2021data, ; zhou2023knowsafe, ), we also evaluate our approach with an end-to-end system used by individuals to manage diabetes with all the uncertainties of the real world.
+Given their life-threatening stakes, automated insulin delivery systems require special consideration beyond the traditional purview of the systems and security communities.
+Research on operating systems or web browsers can list scores of CVEs from current systems and show how their system improves security by eliminating their impact.
+Our position is that we need to make sure that automated insulin delivery systems are never assigned CVEs, because vulnerabilities can kill people.
+Thus, we focus on building a new system, using security-first principles, to avoid security vulnerabilities and mitigate their damage from the start.
 Our main contribution lies in a new end-to-end system that we built and the lessons learned from our experience designing, implementing, and deploying GlucOS to humans to manage their T1D.
 Our novel contributions include:
-  • Security mechanisms and policies that protect individuals from malicious ML, vulnerable pump drivers, and drastic changes in the human’s glycemic response.
+  • Security mechanisms and policies that protect individuals from malicious dosing algorithms, as well as drastic changes in human physiology.
   • A case study describing our experiences from a real-world deployment with seven people, and results from simulation to show that our techniques generalize.
@@
 2. Overview
+Doctors and healthcare professional recognize Type-1 Diabetes as a predominantly self-managed condition and support the adoption of automated insulin delivery systems for management (pho, ; khor2023individualized, ; midyett2023one, ).
 This paper describes our design for GlucOS, a system for trustworthy automated insulin delivery; we have two primary goals.
@@
 2.1. Locking down the insulin delivery path
-We focus our system architecture on providing security mechanisms and policies on the insulin delivery path.
-The insulin delivery path includes models for calculating the amount of insulin still active in the individual (insulin can take up to six hours to fully absorb), closed-loop algorithms for calculating insulin dosing, deep neural network ML to predict future metabolic states, and drivers for programming the insulin pump.
+Our system architecture focuses on providing security mechanisms and policies on the insulin delivery path.
+The insulin delivery path in automated insulin delivery systems includes models for calculating the amount of insulin still active in the individual (insulin can take up to six hours for complete absorption), closed-loop algorithms for calculating and delivering insulin doses that could use deep neural networks to predict future metabolic states.
 Together, these components are both the most complex, and the most sensitive from a security perspective because they are responsible for insulin delivery.
@@
 The BioKernel is the component that interacts with the continuous glucose monitor (CGM) and insulin pump hardware, executes the closed-loop dosing algorithm, runs security and safety checks, and produces event logs that other components use to learn the state of the system.
-The BioKernel runs the closed-loop algorithm every five minutes as the new CGM readings come in.
-Once the BioKernel receives new CGM readings, it collects the current metabolic state of the individual and sends it to a predictive ML model.
-The predictive ML model calculates the amount of insulin to inject based on its assessment of the individual’s current metabolic state and predictions of future metabolic states.
+The BioKernel handles the typical operations of automated insulin delivery systems.
+It runs the closed-loop algorithm every five minutes as the new CGM readings come in.
+Once the BioKernel receives new CGM readings, it collects the current metabolic state of the individual and sends it to a predictive model.
+The predictive model calculates the amount of insulin to inject based on its assessment of the individual’s current metabolic state and predictions of future metabolic states.
 The BioKernel then issues commands to the pump to deliver the calculated amount of insulin.
 A pump driver converts these commands into low-level I/O to program the pump and adjust insulin delivery.
-To withstand attacks, the BioKernel serves as a reference monitor (anderson1972computer, ) to enforce security policies on the predictive ML model’s pump commands and the pump driver’s low-level I/O.
-This reference monitor architecture ensures that we can withstand attacks from malicious predictive ML as well as malicious pump drivers without having to know the internals of how they operate.
-For predictive ML, our reference monitor enforces algorithmic security, where we bound the amount of insulin calculated by ML to stay within dynamic safety bounds computed by a simple reactive safe model.
-For the pump driver, we interpose on the I/O path to ensure that only the pump commands issued by the BioKernel are programmed into the pump.
-These checks constitute our pump invariants, which, when violated switch the pump to operate in a safe mode.
-Our use of formal methods focuses on the software related to delivering insulin, which is the most risky operation in any automated insulin delivery system.
+To withstand attacks, the BioKernel serves as a reference monitor (anderson1972computer, ) to enforce security policies on the predictive model’s pump commands.
+This reference monitor architecture ensures that we can withstand attacks from malicious predictive models without having to know the internals of how they operate.
+Our reference monitor enforces algorithmic security, where we bound the amount of insulin injscted by the model to stay within dynamic safety bounds computed by a simple reactive safe model.
+Our use of formal methods focuses on the software in the insulin delivery path.
 We use Hoare logic to prove key functions correct, define system states and transitions to handle runtime verification check failures safely, and we introduce the notion of a novel biological invariant.
-Our biological invariant is an end-to-end check of our theoretical calculations of how the individual’s metabolism should behave compared to what we observe in practice, where violations of this invariant suggest a potentially catastrophic failure of the system.
+Our biological invariant is an end-to-end check of our theoretical calculations of how the individual’s metabolism should behave compared to what we observe in practice.
+Violations of this invariant suggest the need to temporarily disable automated insulin delivery in response to an unprecedented change to human physiology.
 2.3. Threat model
-The most serious health risk for people who use synthetic insulin is low glucose, or hypoglycemia.
+Automated insulin delivery systems inject synthetic insulin to regulate glucose concentration in the body.
+In the long run, high glucose or hyperglycemia leads to dangerous health outcomes such as kidney failure, heart disease, amputation etc.
+The most serious and immediate health risk for people who use synthetic insulin is low glucose, or hypoglycemia.
 Hypoglycemia results from an overdose of insulin.
@@
 If untreated, severe hypoglycemia can lead to loss of consciousness, seizure, or death in a matter of hours.
-Thus, our primary security objective is to reduce or eliminate hypoglycemia that results from an attack on the system.
-Quantitatively, our goal is to maintain the American Diabetes Association’s recommended time in hypoglycemia of less than 4% (AmericanDiabetesAssociation2021, ).
-As our design philosophy centers around humans, providing flexibility for people to use whatever ML they want by loading personalized models or using any pump hardware they want is important.
-This flexibility helps guide our focus on providing novel security mechanisms and policies for these components.
-Our goal is to make sure that they stay safe even if the model they pick is malicious or the pump driver they choose has security vulnerabilities.
-Attackers may target automated insulin delivery systems because they can directly harm the human using the system.
-We assume that the Swift type system is correct and that the underlying operating system, the CGM driver and hardware, and the pump hardware we use are free of attacks and their device communications maintain integrity.
-Thus, we do not consider attacks from tampered CGM sensors or pump hardware in our threat model.
-These are important but complementary problems to our work, which focuses on the system architecture and algorithms for automated insulin delivery.
-We consider attacks from malicious predictive ML algorithms.
-These attacks can come from trojan attack logic embedded directly in a model, or from an adversarial attack on predictive ML.
-For adversarial attacks, although we assume that the CGM and insulin pump remain attack free, an adversary can manipulate these entities by serving an individual who is using our system with calculated amounts of carbohydrates in their food (e.g., at a restaurant), thus influencing the resulting CGM readings and corresponding insulin commands.
-In our threat model, malicious predictive ML can issue arbitrary insulin doses.
-We also consider malicious pump drivers. In our threat model, malicious pump drivers can issue arbitrary pump I/O to add, drop, or modify commands sent to the pump.
-Successful attacks on predictive ML or pump drivers, if unchecked, can be devastating because they directly affect insulin delivery and can be used by an attacker to induce potentially life-threatening hypoglycemia.
-3. Algorithmic security: Withstanding malicious ML
-Currently, no automated insulin delivery systems use deep neural network ML.
-Some current systems use statistical methods, like linear regression or model predictive control (physio2011, ; physio2019, ; lunze2013control, ; control2019, ; openaps, ).
+Thus, our primary security objective is to reduce or eliminate hypoglycemia.
+Since personalization is important for T1D self-management (glucomaker, ), our design philosophy centers around humans.
+We provide flexibility for people to use any model for insulin dosing that suits their preferences.
+Our goal is to make sure that they stay safe even when they pick vulnerable or malicious models.
+Attackers may target automated insulin delivery systems to directly harm the human using the system.
+For safety, we focus on protecting users from incorrect insulin doses that emerge from both, poorly chosen insulin dosing models, as well as inadvertent mistakes from benign models.
+Malicious or vulnerable models, if unchecked, can be devastating because they directly affect insulin delivery and can lead to life-threatening hypoglycemia.
+We build GlucOS as a complete iOS app, and accordingly assume that the Swift type system, that forms the backbone of the iOS ecosystem, is secure.
+We only consider FDA-approved CGMs and pumps.
+With the FDA being stringent with security guidelines for medical devices (fda_medical, ), we assume that CGM and pump hardware are free of attacks.
+However, we assume that CGMs and pumps are prone to measurement errors within their standard error bounds (garg2022accuracy, ; dexcom_error, ; Zisser2011OmniPodAccuracy, ).
+We discuss malicious pump drivers and ways to overcome them in Appendix D.
+3. Design principles
+Four key principles guide our design of GlucOS:
+  • Principle 1: Focus on simplicity for security mechanisms by using domain knowledge.
+Rather than design a general security mechanism for a broad class of problems, we focus specifically on the highly-risky domain of automated insulin delivery.
+Using domain-specific insights, we provide simple security mechanisms that both humans and theorem proving software can easily reason through.
+  • Principle 2: Design the system to empower humans to provide security.
+Automated insulin delivery represents an extreme example of high-stakes healthcare.
+For people to be able to use a secure system to manage their health, they need to have agency over the security of their system.
+Rather than simply “alerting the human” if something unexpected happens, we empower the humans using GlucOS to be active participants in the security of their system.
+  • Principle 3: Consider asymmetries in risk.
+Insulin dosing possesses an inherent asymmetry where overdosing poses greater immediate danger than underdosing (cryer2012severe, ; benoit2018trends, ).
+With GlucOS, we optimize our mechanisms to be more aggressive when protecting individuals from insulin overdose.
+  • Principle 4: Adaptability for security.
+Although we carefully design our software to be correct, humans and glucose metabolism can change quickly.
+Any change that affects the correctness of models of glucose metabolism would compromise the security of the entire system.
+Thus, we design GlucOS to adapt to constantly changing human physiology to provide security both in our formulation of the problem and our use of formal methods.
+Although we do strive to apply these principles to the entire design, there are times when there is tension between them.
+For example, our biological invariant (Section 5) helps with adaptability (Principle 4) by detecting when our physiological models are wrong and considers asymmetries in risk (Principle 3).
+But to ensure that the human remains empowered (Principle 2), we use necessarily complex logic to deal with cases when the biological invariant does not hold (violation of Principle 1).
+These types of tensions are fundamental to a real-world system that manages human health and underscores the importance of building real systems to gain experience to navigate these tradeoffs.
+4. Algorithmic security: Withstanding malicious dosing algorithms
+We have dual goals for our algorithmic security mechanism of being both defensive and enabling.
+Our security mechanism is defensive by protecting people from inappropriate insulin doses emerging from bugs, corner cases, or even attacks merging from complex predictive algorithms – be it deep neural networks or massive sets of hand-crafted rules.
+Our algorithmic security mechanism is enabling in allowing people to employ any predictive algorithm to personalize the system to suit their individual preferences.
+To enable predictive algorithms, GlucOS’s algorithmic security mechanism introduces a novel architecture that separates the use of predictive models for insulin dosing from accounting and safety logic to ensure that they are operating safely.
+This architecture provides flexibility in using any predictive model while enforcing bounds on insulin doses with a simple and formally verified model for safety.
+4.1. Need for algorithmic security
+Closed-loop automated insulin delivery systems have traditionally used reactive physiological models (such as PID controllers (weinzimer2008fully, ; kushner2018data, )) to calculate insulin doses.
+These models essentially use standard physiological models to react to offset the amount of excess glucose in the body.
+This makes it easy to reason about the correctness of their insulin doses.
+However, because synthetic insulin is slow-acting and takes up to six hours to fully absorb, the reactive nature of these models limits their effectiveness in regulating elevated glucose concentrations.
+Thus, having the ability to inject insulin proactively by predicting glucose concentrations leads to more effective glucose regulation.
+But, predictive dosing opens up the individual to the risk of potentially dosing too much insulin, and current insulin delivery systems have been reluctant to adopt the most advanced forms of predictive models.
+Researchers have shown that modern advances in ML, such as deep neural networks, work well in making predictions for automated insulin delivery (Zhu2022, ; Jacobs, ; Zhang2021, ; Josep2020, ; RL, ; Tyler2022, ).
+Despite their promise, none of the existing automated insulin delivery systems use deep neural networks.
+Instead, they use statistical methods, like linear regression or model predictive control (physio2011, ; physio2019, ; lunze2013control, ; control2019, ; openaps, ).
 The reason for using this more traditional form of ML is sound.
 It provides consistent, explainable results and has a physiological basis that people can use to vet its decisions.
-However, all of the recent advances in ML, like image classifiers and large language models, have come from deep neural networks, which researchers have shown also work well for predicting metabolic states for use in automated insulin delivery systems (Zhu2022, ; Jacobs, ; Zhang2021, ; Josep2020, ; RL, ; Tyler2022, ).
-Despite this promise, none of the current automated insulin delivery systems incorporate more advanced deep neural networks due to the black-box nature of deep neural networks and the potentially dire consequences of mispredictions – an automated insulin delivery system’s unchecked “hallucination” (chatgpt_hallucination, ) could have lethal consequences if it delivers an inappropriate insulin dose (cryer2012severe, ; benoit2018trends, ).
-In this section, we discuss the design and implementation of GlucOS’s algorithmic security mechanisms to enable predictive models, including deep neural network ML, to run safely in automated insulin delivery systems.
+While advanced deep neural networks have more predictive power, current systems do not incorporate them due to their black-box nature and the potentially dire consequences of mispredictions.
+Even a well-tested ML model does not guarantee immunity to misprediction (nfl_ml, ; nfl_adversarial, ), especially when predicting human metabolism that is influenced by numerous ever-changing factors such as food, exercise, stress, environment, allergies, puberty, temperature and so on (fourtytwofactors, ).
+An automated insulin delivery system’s unchecked misprediction or “hallucination” (chatgpt_hallucination, ) could have lethal consequences if it delivers an inappropriate insulin dose (cryer2012severe, ; benoit2018trends, ).
+Beyond the use of ML, the complexity of rule-based predictive models (such OpenAPS’ thousands of lines of code (openaps, )), make it difficult to reason about their correctness, even if they are generally effective at regulating glucose for diabetes management.
+Additionally, people requesting enhancements to such models (cob, ; hypo_protect, ) based on their personal struggles further exacerbates model complexity.
+Thus, avoiding predictive models for insulin delivery comes at the cost of effective diabetes management.
+On the other hand, incorporating them introduces safety concerns stemming from challenges in the explainability of complex code and deep neural networks.
 4.2. Mechanism fundamentals
-Our key novel insight underpinning our security mechanism design is that all correct insulin delivery algorithms will deliver the same amount of insulin over a sufficiently long period.
+The key novel insight underpinning our security mechanism design is that all correct insulin dosing algorithms, regardless of whether they are predictive or reactive, will deliver the same amount of insulin over a sufficiently long period (such as a full digestion-absorption cycle).
 People inject insulin to facilitate glucose absorption after eating.
-Digestion converts food to glucose, and insulin facilitates the absorption of glucose to tissue (e.g., the brain) where the body can use it as an energy source.
-Since insulin dosing is driven by food consumption, all correct algorithms should dose the same amount of insulin to cover glucose from digestion.
-However, the timing of this insulin delivery matters for maintaining balanced glucose levels (Section 9).
-Insulin activation is a slow process, with injected insulin taking up to six hours to fully activate.
-To meet this challenge, our algorithmic security mechanism pairs two models.
+Digestion converts food to glucose, and insulin facilitates the absorption of glucose by tissues (e.g., the brain) where the body can use it as an energy source.
+Since insulin dosing is dominated by food consumption, all correct algorithms should dose the same amount of insulin to cover glucose from digestion.
+But the timing of these insulin doses has a large impact on the individual’s glucose regulation.
+Our algorithmic security mechanism pairs two models.
 Our first model is a predictive model that anticipates future metabolic states and doses insulin proactively.
 Our second model is a reactive safe model that measures the current metabolic state and doses insulin based only on what it can measure currently.
-This pairing allows us to balance proactive insulin dosing for faster responses with a more conservative approach for safety.
+This pairing balances proactive insulin dosing for faster responses with a more conservative approach for safety.
+As per our insight, the reactive model would dose enough insulin eventually, so we use it as our baseline but provide predictive models with enough buffer to inject insulin earlier and safely.
 Architecturally, in a novel approach for automated insulin delivery systems, the predictive model runs outside of our trusted computing base while the reactive safe model is part of the trusted computing base.
@@
 It also defends against malicious predictive models, as all insulin delivery decisions are ultimately vetted by the reactive safe model.
-This design also enables people to update their predictive models without violating core safety guarentees.
+This design also enables people to update their predictive models without violating core safety guarantees.
 Conceptually, the predictive model is the primary model in our system and should control the pump most of the time.
@@
 By combining these two models, we strive to get beneficial properties from both.
-Our approach combines the safety of a reactive safe model and the performance from a predictive model.
+Our approach combines the safety of a reactive model and the performance from a predictive model.
 Our security mechanism draws inspiration from previous work on combining simple and complex models for safety (sha2001using, ; mohan2013s3a, ).
@@
 The whole process starts when a new CGM reading arrives.
-Once the BioKernel receives a new CGM reading, it collects a recent history of CGM readings and insulin dosing and sends these as inputs to both the reactive safe model and the predictive ML model.
-These models independently produce pump commands, called a temporary basal command.
-The temporary basal command sets the insulin delivery rate for the insulin pump for a 30-minute duration.
-The BioKernel will collect both temporary basal commands and look back at the differences between the reactive safe model and predictive ML model for the past three hours to see how much insulin delivery it should attribute to the predictive ML model.
-If the predictive ML model is still within its safe insulin delivery limits, it uses the predictive ML model’s temporary basal command to program the pump.
-If it has exhausted its limits, we clamp it to fit within the insulin bounds we set relative to the reactive safe model.
+Once the BioKernel receives a new CGM reading, it sends the recent history of CGM readings and insulin doses to both the reactive safe model and the predictive model.
+These models independently produce pump commands (called a temporary basal command) that set the insulin delivery rate for the pump over the next 30 minutes.
+The BioKernel then collects both commands and looks back at the differences in doses between both models over the past three hours to calculate how much insulin the predictive model has dosed.
+If the predictive model is still within the safe insulin delivery bounds relative to the reactive safe model’s doses, the BioKernel uses the predictive model’s command to program the pump.
+If the predictive model has exhausted its limits, we clamp its calculated insulin to fit within the insulin bounds relative to the reactive safe model.
 To set bounds, we base our calculations on human physiology.
 People produce a background level of glucose for energy.
-People living with T1D need to inject insulin to facilitate the absorption of this background glucose.
 Our method for setting bounds allows the predictive model to use up to three hours worth of background insulin in advance.
 If the prediction is correct and glucose levels rise (e.g., from eating), the proactive insulin will manage it.
-If the prediction is wrong and no extra glucose appears, the body’s ongoing glucose production will help use the extra insulin over the next three hours.
-In the worst case, an incorrect prediction might lead to low glucose.
-However, with our mechanism in place, we limit the impact by our three-hour bound and offset by the body’s continuous glucose production.
+If the prediction is wrong and no extra glucose appears, the body’s ongoing glucose production will use the extra insulin over the next three hours.
 4.3. Model principles
+The mechanism described above can be implemented with various algorithms.
 Our contribution is not in the specific algorithms we use, but rather in the insight around insulin delivery being equal for all correct algorithms and how we use this insight for security.
 In this section, we outline the algorithms we use in our current implementation and define the principles for how to apply security to any predictive insulin dosing algorithms.
-Algorithmically, we use a PID controller (franklin2002feedback, ) from feedback control for our reactive safe model, and we use a deep neural-network from the literature for meal predictions for our predictive ML model (mosquera2023enabling, ).
-Although we use this model in our current implementation, our system supports any predictive model.
-PID controllers are used in automated insulin delivery systems (weinzimer2008fully, ; kushner2018data, ) and research predictive ML models for insulin delivery have been well covered in the literature (Zhu2022, ; Jacobs, ; Zhang2021, ; Josep2020, ; RL, ; Tyler2022, ).
+Algorithmically, we use a PID controller (franklin2002feedback, ) from feedback control for our reactive safe model, and we use a deep neural network from the literature for meal predictions for our predictive ML model (mosquera2023enabling, ).
+Although we use this model in our implementation, our system supports any predictive model.
 The design of the reactive safe model is crucial for algorithmic security, as it serves as the anchor for the security mechanism.
@@
 To ensure the safety of this model, we base its calculations solely on facts observed automatically from the CGM and the insulin pump.
-Given these high-quality sources of data, and a simple insulin absorption physiological model, we can calculate how much insulin an individual needs and how much insulin they have injected but have not yet activated.
+Given these high-quality sources of data, and a simple insulin absorption physiological model, we can calculate how much insulin an individual needs and how much previously injected insulin has not yet taken effect.
 From this calculation, we know how much insulin to inject or withhold.
@@
 The difference in our system is that we run these calculations automatically and every five minutes to adjust to the latest sensor readings.
-This basic formulation is a standard formulation used in most automated insulin delivery systems.
-In contrast, predictive models have no restrictions since their outputs are constrained by the reactive safe model and our security mechanism.
-The ideal predictive model anticipates future metabolic states and adjusts insulin proactively to avoid glycemic imbalance.
-In our evaluation, our predictive model outperforms the reactive safe model both in simulation (Section 9) and for the individual that uses our system for their real-world insulin dosing (Section 10), showing why it is important for automated insulin delivery systems to have a practical way to adopt ML.
-4. Driver security: Withstanding malicious pump drivers
-Pump drivers provide abstractions for insulin delivery applications to communicate with insulin pumps.
-Much like traditional device drivers, they have complex implementations since they help accomplish a wide range of tasks including pump setup/teardown, managing pump UX, handling insulin delivery/suspension, managing pump errors, accurately tracking insulin doses, and so on.
-Given this complexity, there is a strong possibility for bugs or vulnerabilities to be present in existing pump drivers.
-These vulnerabilities can harm or even lead to the death of individuals since pump drivers control insulin pumps.
-For example, consider a scenario where the user decides to inject 3 units (3U) of insulin using the BioKernel.
-The BioKernel would relay this to the pump through abstractions provided by the pump driver.
-If the pump driver inadvertently or intentionally issues this command twice, 6U would be delivered, which could endanger the user by dosing too much insulin.
-Our security goals for GlucOS’s driver security are to protect users from inadvertent bugs in pump drivers and safeguard against potentially malicious pump drivers.
-To achieve these goals, we introduce a shim between the pump driver and the pump, which validates all communications between them.
-We design our driver security mechanism to interpose on existing pump driver-pump communication instead of directly programming the pump.
-This design ensures that our trusted computing base remains small and simple by not including the complexity of pump driver implementations, thereby enabling formal verification.
-This design also provides flexibility with altering or introducing new functionality to pump drivers.
-The most interesting aspect of our driver security mechanism was having to overrule our principle of not programming the pump in certain cases in the interest of protecting the individual.
-These cases involve commands pertaining to canceling ongoing insulin delivery.
-We could not directly adopt the design of other device driver security mechanisms (ref_validation, ; tang2010trust, ; expressos, ) for our driver security mechanism since these cases are specific to insulin pumps.
-We focus our security efforts on pump drivers because they are complex (14.5k LoC), multi-threaded, event-driven, and use two-way communication between the pump and the driver.
-In contrast, CGM drivers are simple (1.9k LoC) and only read sensor values from the device.
-In our future work, we will formally verify our CGM driver using the same techniques we use for the insulin delivery path (Section 7).
-4.1. Driver security mechanism design
-Figure 4. We interpose on communications between the pump driver and the pump to handle untrusted pump drivers.
-Figure 4 provides an overview of our driver security mechanism.
-The BioKernel registers high-level commands, like “deliver 2U of insulin” with our driver security mechanism, and sends them to the pump driver.
-The pump driver converts these commands into I/O messages to program the pump.
-The driver security mechanism then checks to make sure that these I/O messages are consistent with the previously registered high-level pump commands before forwarding them to the pump hardware.
-We only validate commands for insulin delivery since they pose the biggest threats to the user.
-On detecting added, modified or dropped commands from the pump driver, the driver security mechanism rejects further commands and transitions the system to a manual mode where the user takes control.
-Dropped commands pertaining to canceling ongoing insulin delivery requires separate consideration since excess insulin is dangerous and cannot be withdrawn once it has been delivered.
-Users will cancel insulin delivery if they accidentally issued an insulin delivery command or incorrectly set the insulin dosage.
-In cases where a malicious driver drops a cancel command, the driver security mechanism directly issues I/O messages to the pump to cancel insulin delivery before transitioning the pump to a manual mode.
-We make the tradeoff of increasing the size of our trusted computing base to handle these cases in the interest of protecting the user.
-Fortunately, there are only two commands to cancel insulin delivery in our current implementation, so the increase to our trusted computing base is small.
-5. End-to-end security: Biological invariant
-When implementing security systems for humans, even the perfect algorithmic and driver security mechanisms are unable to handle all situations due to changes in human physiology.
-A key novelty in our overall approach is our introduction of an end-to-end check that ensures the BioKernel’s glucose calculations against observations of the user’s metabolic state.
-If there is a large mismatch between our calculations and what we observe, then it means that our calculations are wrong.
-These deviations can happen from drastic changes in the user’s glucose absorption, like during exercise when the body uses glucose without needing insulin (borghouts2000exercise, ), or if the pump insertion site hits a vein and delivers insulin directly into the blood stream (steyn2023accelerated, ).
-Regardless of the cause, the end-to-end check detects these conditions and prevents harm to the user.
-We call this safety check that accounts for changing human physiology a biological invariant.
-In our implementation, we check that changes in glucose levels that we observe from CGM readings and the individual’s insulin delivery history are consistent with our theoretical calculations of how much the person’s glucose should change given these same inputs.
-If their glucose is dropping by more than our theoretical calculations show, then they are more sensitive to insulin then our system believes they should be.
-We set a threshold of 30 mg/dl per hour, for the maximum allowable divergence based on clinical evaluations showing detection of significant glycemic deviations above this level (30mg_1, ; 30mg_2, ).
-We take a three-tiered approach to dealing with deviations in our glucose change calculations.
-First, for small differences of less than 30 mg/dl, we accumulate errors and adjust the target glucose level to compensate, clamping the adjustment to 10 mg/dl.
-For example, if the individual’s insulin sensitivity is elevated and their target glucose level is 90 mg/dl, GlucOS will update their target to 100 mg/dl dynamically to adjust.
-Second, when the differences are 30 mg/dl or greater, we shut off insulin delivery.
-Our rationale is that the individual is more sensitive to insulin than our models believe they should be, thus we want to avoid overdosing.
-If their difference goes back to within the expected range, normal execution resumes.
-Third, if they remain at 30 mg/dl or more for more than two hours, we alert the individual and transition the system into a manual mode that requires the user to dose insulin manually.
+This basic formulation is standard and used in most automated insulin delivery systems.
+Our evaluation shows that a benign predictive ML model outperforms the reactive safe model in terms of glucose regulation, both in simulation (Section 9) and for the individual who uses our system for their real-world insulin dosing (Section 10).
+Our evaluations in simulation also demonstrate the fatal consequences of using an incorrect or malicious ML for insulin dosing, as well as our mechanism’s ability to withstand such models to protect individuals (Section 9).
+5. Holistic security: Biological invariant
+We recognize that human physiology is complex and it is difficult to ensure that the BioKernel’s reactive safe model, which forms the basis of algorithmic security, can handle all changes to human physiology.
+Concretely, since the reactive safe model incorporates parameters that represent the user’s physiological state, any divergence in their values could render the model ineffective in protecting the user.
+With GlucOS, we introduce a biological invariant to detect when such divergence could endanger the user, and shut off automatic insulin dosing to protect them.
+From a high level, the biological invariant runs experiments every 5 minutes to empirically measure physiological parameters.
+These experiments compare the expected drop in glucose from prior insulin doses against the observed change in glucose sensor readings, thereby providing an end-to-end security mechanism.
+The biological invariant effectively serves as a “catch-all” off-switch for insulin delivery when the BioKernel’s view of the user’s physiology becomes inaccurate.
+Although our reactive safe model adapts to small changes in physiological parameters by design, the biological invariant detects large changes that operate outside of the bounds of our design.
+Insulin sensitivity, which is the amount of glucose absorbed per unit of insulin, is a key physiological parameter that can change.
+Changes to insulin sensitivity happen from drastic changes to the user’s glucose absorption, such as during exercise, when the body uses glucose without needing insulin (borghouts2000exercise, ), or if the pump insertion site hits a vein and delivers insulin directly into the blood stream (steyn2023accelerated, ).
+In such cases, the user becomes more sensitive to insulin where the same amount of insulin will facilitate more absorption of glucose than expected, increasing the risk of insulin overdose.
+For example, if the user becomes twice as sensitive to insulin while jogging, the same insulin dose would decrease their glucose by twice the expected amount, pushing them towards dangerously low glucose (hypoglycemia).
+To implement the biological invariant, we compare the actual change in the user’s glucose, against the BioKernel’s calculations of how much their glucose should have changed based on current insulin doses over a fixed time window.
+If the observed drop in glucose deviates beyond a clinically determined threshold of 30 mg/dl (30mg_1, ; 30mg_2, ) from the calculated drop in glucose within a window of 30 minutes, we consider the biological invariant to have been violated and shut off insulin delivery.
+Mathematically, we define the biological invariant as follows:
+(1) $\Delta G^{C}-\Delta G^{A}\leq T$
+Here, $\Delta G^{C}$ denotes the BioKernel’s calculated drop in glucose based on insulin doses over a 30 minute window, $\Delta G^{A}$ denotes the actual drop in the user’s glucose over the same time window, and $T$ is the threshold of 30 mg/dl.
+Crucially, we note that we cannot precisely calculate $\Delta G^{A}$ or $\Delta G^{C}$ since CGMs and pumps are prone to measurement errors.
+To account for these errors under standard use, we enforce error bounds on readings from CGMs (garg2022accuracy, ; dexcom_error, ) and records from pumps (Zisser2011OmniPodAccuracy, ) under standard conditions.
+Based on these bounds, we check the following condition on the BioKernel’s current estimate of insulin sensitivity:
+(2) $S\leq\frac{0.9\cdot G_{t+\Delta t}^{M}-1.1\cdot G_{t}^{M}+T}{B\cdot\Delta t-0.99\cdot\Delta IOB^{M}+0.01\cdot\Sigma I_{i}^{\ M}}$
+Here, $S$ is the BioKernel’s estimate of the user’s insulin sensitivity, $G_{t}^{M}$ and $G_{t+\Delta t}^{M}$ are readings from the CGM, $\Delta IOB^{M}$ and $\Sigma I_{i}^{M}$ are calculated from the insulin doses recorded by the pump, $\Delta t$ denotes the 30 minute window, and $B$ is the user’s basal rate, a physiological parameter that often remains stable for days and is less prone to change than insulin sensitivity (Lindmeyer2021BasalProfiles, ).
+We refer to this condition as the implementation invariant and use formal methods to prove that violations to this condition imply violations to the biological invariant (Section 7).
+Since all quantities in the implementational invariant are observable, we shut off insulin delivery when it is violated.
+We show how to adopt mathematical models from the literature (simglucose, ; physio2019, ) to derive the implementational invariant in Appendix B.
+We resume normal execution if the implementational invariant goes back to being satisfied within two hours, based on clinical guidance for exercise for people living with T1D (ucla_t1d, ; borghouts2000exercise, ).
+If it remains violated for more than two hours, we alert the user and transition to manual insulin delivery.
+Appendix G discusses the less dangerous scenario when the actual change in glucose is higher than the calculated change.
 6. Last line of defense: Humans
 In this section, we describe our design and implementation for incorporating humans as a critical component of our multi-layered security system.
-While our algorithmic, system-level, and biological invariant defenses form the primary barriers against attacks, humans serve as the last line of defense, capable of mitigating attacks that manifest as glycemic imbalance.
-We discuss our most interesting findings around preferences and the architectural iterations we make to our design based on our real-world experiences.
+While our algorithmic security mechanism and biological invariant defenses form the primary barriers against attacks, humans serve as the last line of defense, to take corrective action against hypoglycemia.
+We discuss our most interesting findings around user preferences and the architectural iterations to our design from real-world experiences.
 6.1. Predictive alerts for humans
@@
 To recover from hypoglycemia, people need to eat sugar.
-Given that people with T1D are already active participants in managing glycemic imbalance, we focus on how we can use this fact to our advantage to create a more comprehensive security system.
+Given that people with T1D are already active participants in managing diabetes, we focus on how we can use them to create a more comprehensive security system.
 Figure 4. GlucOS predicts hypoglycemia and notifies the user before it happens so they can proactively prevent it.
-We use predictive yet deterministic mobile notifications to alert people when they are likely to experience hypoglycemia before they reach it (Figure 5), as hypoglycemia presents the most immediate health risk and could be an indicator of an attack.
-Alerts for hypoglycemia are not new.
+We use predictive yet deterministic mobile notifications to alert people about impending hypoglycemia before it happens (Figure 4).
+We focus on hypoglycemia since it presents the most immediate and dangerous risk.
+Alerts for hypoglycemia ae not new.
 All CGM software comes with non-predictive hypoglycemia alerts by default (dexcom_alerts, ) and glucose prediction algorithms have been well covered in the literature (Josep2020, ; Zhang2021, ; arora2023multivariate, ).
@@
 In a user study (Section 10.4), we reduced the amount of time six people spent in hypoglycemia as a result of using GlucOS, despite them already having CGM alerts set up prior to our study.
-These results demonstrate a reduction in hypoglycemia for the non-attack case, but the most dangerous successful attacks will result in hypoglycemia.
-Thus, these results suggest that predictive alerting can be an effective security measure.
+We highlight that we did not conduct our user study in a controlled environment, but had participants use our app for predictive alerts as they managed their diabetes in the real world.
+We also highlight that The number of participants in our study is also roughly consistent with recent human factors in computing research (glucomaker, ; moont1d, ; context_aware_t1d, ; t1d_children, ).
+Our results show that predictive alerts can be effective as a security measure against hypoglycemia.
 Using a simple and deterministic algorithm is important since it runs as a part of our trusted computing base.
-Algorithmically, we predict hypoglycemia by doing a least squares fit for the last 20 minutes of CGM readings and extrapolate forward 15 minutes.
-We trigger alerts if our prediction shows hypoglycemia is likely in the next 15 minutes.
+Algorithmically, we predict hypoglycemia using linear regression over the last 20 minutes of CGM readings and extrapolate forward 15 minutes.
+We trigger alerts if our prediction crossed the hypoglycemia threshold in the next 15 minutes.
 The most interesting finding from our study was that personalization was important.
-We provided people with the ability to set the prediction threshold and specify how often they wanted the alert to repeat while they remained out-of-range.
-Five of the six participants customized their alerts and the participant noted that the default values we picked coincided with their alerting preferences.
-In our first implementation we did not provide the ability to customize their security settings and we tried to make the usability vs security tradeoff algorithmically and automatically.
-But in the end, simplifying the system and allowing users to personalize their security policy was important for our participants.
+We provided people with the ability to set the prediction threshold and specify how often they wanted the alert to repeat while they remained in hypoglycemia.
+Five of the six participants customized their alerts and the remaining participant noted that the default values we picked coincided with their alerting preferences.
+In our first implementation we did not provide the ability to customize their settings and we tried to make the usability vs security tradeoff algorithmically and automatically.
+But in the end, simplifying the system to enable the personalization of alerting policy was important for our participants.
 Our finding has the potential to have broader implications for security in healthcare systems.
@@
 It adds complexity to our trusted computing base, but we carefully manage this tradeoff.
-We minimize the added complexity by using traditional, well-understood algorithms for prediction rather than more complex ML models and reuse existing functionality within the BioKernel when possible to keep the implementation simple.
+We minimize the added complexity by using traditional, well-understood algorithms for prediction rather than complex ML models and reuse existing functionality within the BioKernel when possible to keep the implementation simple.
 This approach allowed us to maintain a high level of confidence in the security and correctness of our alerting system while still providing the necessary functionality.
 7. Formal verification
-To provide additional assurance that our implementation is correct, we port our Swift code to Dafny and use Hoare logic.
-As the system we are verifying is complex, our goal is to locate the pieces and target verification in places where formal methods can have the greatest impact; this is a common reality of applying formal methods at scale, see for example (chong2020code, ).
-Since we are porting the Dafny code to Swift, to reduce the chance of human error, we use a line-by-line port as much as possible and add unit tests for each function to provide confidence that the two implementations are equivalent.
-We port updated code back to Swift manually because Dafny does not include code generation for Swift.
-These efforts paid off: we found and fixed nine bugs in our original implementation; in addition to proving various invariants in the code for our insulin delivery path and security mechanisms.
-However, the most interesting part was where these bugs were: all of our core algorithms were correct, but we found bugs in seemingly less critical parts of the system, like user settings, which had subtle interactions with the core algorithms.
-Our implementation operates on CGM readings, but in our application of formal methods, our goal is to reason about the individuals actual glucose level, taking into account sensor errors.
-This raises an interesting challenge for formal verification as it requires modeling the true biological state of the user – a parameter that is both unknown at runtime and distinct from measured parameters that are part of the software implementation.
-We prove that the biological invariant holds, given error bounds set by the CGM manufacturer (garg2022accuracy, ).
-To reason about actual glucose levels formally, we use a ghost variable in Dafny.
-Ghost variables are variables that are only available to the prover, these do not exist in the real software system.
-By defining the actual glucose levels using a ghost variable, we can include it in preconditions and postconditions to prove properties about our software relative to actual glucose levels.
-We believe that having a parameter for the real biological state of the system is an important modeling decision; and it diverges, for example, from previous work that only models the biological data in simulation (zhou2021data, ; zhou2023knowsafe, ).
-Traditional software verification techniques like model checking (cps_1, ) and deductive verification (cps_3, ) have been explored in the past for hybrid system verification (cps_2, ) that interact with the human body, but they do not completely account for physiological variability between individuals or model inaccuracies during the insulin delivery process.
-Also, researchers have used real-world invariants for security in the context of autonomous vehicles (choi2018detecting, ), but these were based on quad-copter physics not human physiology and did not include formal verification of their implementation.
+We formally verify the most critical components of GlucOS to prevent unexpected and undesired outcomes (Thomas1994TheSO, ).
+Specifically, we verify the implementation for: (1) delivering bolus insulin doses (a single dose delivered immediately), (2) delivering basal insulin doses (multiple doses delivered at a constant rate over a specified time period), (3) clamping insulin doses using the reactive safe model, and (4) enforcing the biological invariant.
+We consider these components to be the most critical since they implement our security mechanisms and directly control insulin delivery.
+Formally verifying our GlucOS’s implementation in Swift presented an interesting challenge since automatic translators do not exist between Swift and languages for verification, such as Dafny.
+While automatic translators exist between Dafny and other languages (such as C) we could not pick those languages over Swift for implementation.
+This is because Swify provides the most support to build iOS applications and our goal was to build a complete automated insulin delivery application for real-world use.
+To overcome this limitation, we manually ported our Swift implementation to Dafny.
+While manual translation is prone to human error, we took important steps to reduce the likelihood of such errors.
+Concretely, we ported our Swift code line-by-line to Dafny, preserved variable and function names across implementations (chong2020code, ), and wrote unit tests for each function in both languages to ensure behavioral consistency.
+We establish invariants among the aforementioned critical components along the insulin delivery path (Thomas1994TheSO, ).
+These invariants enforce critical constraints such as maintaining appropriate units insulin doses, disallowing immediately successive bolus doses, ensuring basal doses match doses supported by the pump etc.
+Collectively these invariants are important to ensure the expected safe operation of the BioKernel.
+We discover 9 bugs in our original implementation by verifying 31 invariants across 8 functions.
+Other than the biological invariant (Section 5), some of our other invariants enforced that there should be a gap of at least 4.2 minutes between bolus doses, the amount of insulin on board should also take into account the insulin from basal doses etc.
+We list all our invariants and bugs in Appendix F.
+From our experience with formal verification, we report two key takeaways about our implementation for secure automated insulin delivery.
+First, we were able to prove the correctness of even those components that involve unobservable quantities (such as the user’s true physiological state in the biological invariant) under standard operation.
+Second, while we expected most bugs to reside within the implementation of core algorithms, they were fully concentrated among peripheral components such as user settings.
+7.1. Capturing true physiological state
+Glucose readings from CGMs and insulin doses recorded from pumps do not reflect the user’s true physiological state.
+This is because CGMs and pumps are subject to measurement errors, making precise glucose concentrations and insulin doses unavailable in Swift.
+To model this discrepancy, we use ghost variables in Dafny (Dafny, ) to represent the true glucose values and insulin doses.
+We then introduce device invariants that bound the maximum measurement error within standard use, i.e., 10% for CGMs (garg2022accuracy, ; dexcom_error, ) and 1% for pumps (Zisser2011OmniPodAccuracy, ).
+We enforce these bounds as preconditions when verifying the implementation of all our core components.
+Most importantly, we proved the correctness of the implementational invariant (Equation $(2)$ in Section 5) to verify the biological invariant (Equation $(1)$ in Section5 as a post condition in Dafny).
+7.2. Discovering bugs outside the core algorithms
+7 out of the 9 bugs we discovered with formal verification were related to issuing insulin delivery commands with invalid parameters.
+With these bugs, users could specify negative amounts of insulin for basal and bolus doses, set 0-second duration for basal doses, allowing the time to stop a basal dose to be preced the start time of the dose etc.
+Other bugs involved insfficient error handling, for example when disallowing certain invalid bolus doses.
+We did not discover any bugs in GlucOS’s core algorithms, such as those used by the reactive safe model.
+We resolved bugs by updating our Swift implementation to match the formally verified Dafny code.
 8. Implementation
@@
 9. Evaluation
-In this section, we evaluate GlucOS’s algorithmic and driver security mechanisms. In the next section (Section 10), we evaluate all components of GlucOS in the real world.
+In this section, we evaluate GlucOS’s algorithmic security mechanism in simulation.
+In the next section (Section 10), we evaluate all components of GlucOS in the real world.
+We do not evaluate the biological invariant in simulation since existing simulators do not support changes to human physiology.
 9.1. How effective is GlucOS’s algorithmic security mechanism?
 We evaluate GlucOS’s algorithmic security mechanism on 21 virtual humans from an FDA-approved simulator (simglucose, ).
-GlucOS’s algorithmic security mechanism should protect users against untrusted ML algorithms that administer incorrect insulin doses (overdose or underdose), while minimally impeding algorithms that are efficient at managing T1D.
-We consider extreme cases of attack in our evaluation.
-While subtle attacks are also possible, we focus on extreme attacks since they have the most adverse consequences.
-Concreteny, we first run simulations with a predictive ML model inspired from prior research (mosquera2023enabling, ), without incorporating our security mechanism to set a baseline of an effctive algorithm in managing T1D.
-Then, in the same simulation scenarios, we introduce an order of magnitude increase to the insulin doses computed by the baseline algorithm, without incorporating our security mechanism to quantify the risks of untrusted algorithms.
-We then repeat these experiments with our security mechanism to quantify the impact of the mechanism on correct and incorrect dosing.
+The mechanism should protect users against untrusted models that administer incorrect insulin doses (overdose or underdose), while minimally impeding algorithms that are efficient at managing T1D.
+While subtle attacks are also possible, we focus on extreme attacks in our evaluation since they have the most adverse consequences.
+We first run simulations with a predictive ML model inspired from prior research (mosquera2023enabling, ), without incorporating our security mechanism to set a baseline of an effctive algorithm in managing T1D.
+Then, in the same simulation scenarios, we introduce an order of magnitude increase to the baseline algorithm’s insulin doses to quantify the risks of untrusted algorithms.
+We then repeat these experiments with our security mechanism to evaluate its impact.
 The American Diabetes Association (ADA) recommends that individuals with T1D spend at least 70% of their time in the range (TIR) of 70-180 mg/dl to remain healthy (nathan2014diabetes, ; endocrine_tir, ; journal_glycemic_targets, ).
 Glucose levels less than 70 mg/dl (hypoglycemia) put individuals at risk for acute complications like impaired consciousness, seizure, and death, while levels greater than 180 mg/dl (hyperglycemia) increase risks for long-term vascular damage.
+We use TIR as our metric in these evaluations.
 9.1.1. Without algorithmic security mechanism
 Predictive ML model baseline With the predictive ML model, we followed all ML best practices in carefully tuning hyperparameters and evaluating the model on simulation scenarios that were different from those used in training.
-Figure 6 shows the average proportion of time spent in the safe target range (TIR) by virtual humans from different age groups when employing the predictive ML model.
+Figure 5 shows the average proportion of time spent in the recommended healthy range (TIR) by virtual humans from different age groups when employing the predictive ML model.
 We see that individuals across all groups spend over 85% of their time in range compared to 77% when using only the reactive safe model.
-These results provide a baseline for an algorithm that is effective in managing T1D.
+These results provide a baseline for a predictive algorithm that is effective in managing T1D.
 Reactive TIR with TIR with
@@
 Average proportion of time spent in range (TIR) across 21 virtual humans with a reactive safe model, a predictive ML model, and clamping the predictive ML with the reactive safe model for security.
-Incorrect ML dosing Since GlucOS is designed to be flexible with any predictive algorithm, it could be subject to attacks from malicious ML algorithms.
-In this section, we discuss an experiment with a malicious algorithm that intentionally doses too much insulin.
+Incorrect ML dosing In this section, we discuss an experiment with a malicious algorithm that intentionally doses too much insulin.
 Concretely, our malicious algorithm doses ten times the amount of insulin computed by the previously described predictive ML model.
-We ran this algorithm within the simulator on the same virtual human on the same scenarios as the previous experiment.
-Figure 7 shows the CGM readings on three randomly chosen virtual humans on a particular scenario when we dose more.
-The CGM readings of all 3 individuals fully go below the hypogycemic bound, which would have killed them in the real-world.
+We ran this algorithm within the simulator on the same virtual humans on the same scenarios as the previous experiment.
+Figure 6 shows the CGM readings on three randomly chosen virtual humans. The readings fully go below the hypogycemic bound, which would have killed them in the real world.
 Figure 6.
@@
 Correspondingly, we observe an average increase in the time spent in hypoglycemia from 0% to 100%.
-We also observed similar results with experiments where we intentionally underdosed insulin.
-In these experiments, all virtual humans spend most of their time above the hyperglycemic bound which wouldn’t have been fatal but would have led to severe health consequences in the real world.
-These results motivate the need for our algorithmic security mechanism.
+We report similar results with intentional underdosing in Appendix C.
+These results demonstrate the need for algorithmic security.
 9.1.2. With algorithmic security mechanism
 Impact on incorrect dosing.
-We repeat experiments from the previous section with the same scenarios.
-We run the same malicous model to inject too much insulin, but also run our reactive safe model (Section 3) to bound insulin injection.
+We repeat experiments from the previous section with the same simulation scenarios.
+We run the same malicous model to overdose insulin, but also run our reactive safe model (Section 4).
 Figure 7 show CGM readings on the same three virtual humans on the same scenarios as the experiment in the previous section.
-We see that all of them spend over 70% of their time in range, showing that GlucOS protects individuals even in presence of malicious overdosing.
-On average, across all virtual humans, we only see a drop in TIR from 88.44% with correct dosing to 78.22% when employing safety while overdosing.
-We also report similar results on experiments with a malicious algorithm that underdoses insulin.
-While the times spent in range are less than those of correct dosing, the impact of incorrect dosing is significantly dampened with individuals spending the recommended amount of time in range, despite facing active attacks.
+We see that all of them spend over 70% of their time in the healthy range, showing that GlucOS protects individuals from malicious overdosing.
+On average, across all virtual humans, we only see a drop in TIR from 88.44% with correct dosing to 78.22% when employing GlucOS while overdosing.
+We also report similar results on experiments with a malicious algorithm that underdoses insulin in Appendix C.
+While the times spent in range are less than those of correct dosing, the impact of incorrect dosing is significantly dampened with individuals spending the recommended amount of time in the healthy range, despite active attack.
 Figure 7.
@@
 When employing the reactive safe model as is to calculate insulin doses, we report an average TIR of 82.48% which is lower than the average TIR of 86.84% observed when running the predictive ML model with GlucOS’s security mechanism.
-This shows that GlucOS’s algorithmic security guarantees do not come at the cost of effective T1D management from predictive ML.
-Figure 9.
-We attached an Omnipod DASH insulin pump to a stuffed animal and modified the OmniBLE pump driver to intentionally add, drop and modify insulin commands.
-Our driver security mechanism detected all manipulations.
-9.2. Does GlucOS ensure pump operation while overcoming malicious pump drivers?
-Our mechanism should ensure that all communications go through with a benign pump driver and should catch all manipulated communication with a malicious pump driver.
-Impact on malicious pump drivers.
-We attached an Omnipod DASH insulin pump to a stuffed animal (Figure 9).
-We then modified the OmniBLE pump driver to intentionally add, drop and modify insulin commands.
-We detected all manipulated commands thereby showing that our driver security mechanism has 0 false negatives in detecting them.
-Impact on benign pump drivers.
-For one week, we logged messages passed between the BioKernel, the unmodified OmniBLE pump driver and the pump on an individual running GlucOS (Section 10).
-We report 100% consistency between the commands sent to OmniBLE and commands issued from OmniBLE showing that GlucOS’s driver security mechanism does not incur any false positives.
-These results show that GlucOS can ensure pump operation while withstanding attacks from malicious drivers.
+This shows that GlucOS’s algorithmic security mechanism does not impede effective T1D management.
 10. Using GlucOS in the real world
@@
 One individual, who we will call Bob, has been using the full GlucOS system since November 2023 for insulin delivery and predictive alerting.
+The number of participants in our study is roughly consistent with existing research in human factors in computing (glucomaker, ; moont1d, ; context_aware_t1d, ; t1d_children, ).
+More importantly, we report results on indiviuals using GlucOS to manage their diabetes in the real world.
 10.1. Ethical considerations
 In this section, we outline the steps we take to ensure that we uphold high ethical standards and ensure that the people using our system, are safe.
-Our user study for the six students was approved by our university’s IRB, and Bob’s use GlucOS was determined by our university’s IRB to be “self evaluation” and exempt from IRB.
+Our user study for the six students was approved by our university’s IRB.
+Bob’s use GlucOS was determined by our university’s IRB to be “self evaluation” and exempt from IRB.
 One of the co-authors of this paper is a board-certified Endocrinologist, who specializes in T1D.
 They designed the safety protocol, which defines the specific criteria we use to stop the study, if needed.
-The Appendix will include a more detailed discussion of our ethics and safety protocols.
+Appendix A includes a more detailed discussion of our ethics and safety protocols.
 10.2. Is an ML-based closed-loop system safe and effective in practice?
@@
 Overall, Bob had the best results when using ML.
-His time spent in range increased from 96.3% to 97.2%, and notably he spent no time in hypoglycemia (below 70 mg/dl), down from 0.67% when not using ML.
+His time spent in range increased from 96.3% to 97.2%, and notably he spent no time in hypoglycemia (below 70 mg/dl), going down from 0.67% when not using ML.
+More importantly, Bob’s A1C, which captures his long-term glucose concentration, dropped to 5.8% as a result of using GlucOS.
+This result demonstrates a significant health outcome since it matches up to non-diabetics who have an A1C of 5.6% or lower.
 Next, we measure how often the predictive ML model was responsible for programming the pump vs our reactive safe model.
 For the one week when Bob was using the predictive ML model, 33.0% of the time both the predictive ML model and reactive safe model issued the same commands, 39.9% of the time the BioKernel used the predictive ML model’s commands, and 27.1% of the time the BioKernel used the reactive safe model’s commands.
+The reactive safe model having to take over to issue commands is not surprising since ML models are not immune to mispredictions, especially when taking the variability of the real world into account.
 We believe that this distribution represents a suitable balance for providing the ML with enough flexibility to improve outcomes, while having tight enough bounds to limit potential damage from bad or malicious ML predictions.
+As of July 2025, Bob has swapped the predictive ML model with a rule-based model similar to OpenAPS’ model described in Section 1.
 10.3. Do biological invariant violations trigger in practice?
@@
 Of these violations there were 49 sequences of 30 minutes or longer with the longest sequence being 65 minutes.
-These results suggest that although natural variations in glucose metabolism do lead to biological invariant violations, none of them were long enough to trigger our most severe action of transitioning to manual mode after two hours.
-We reviewed one weeks worth of violations with Bob manually to determine the cause. During the most recent week of data, Bob had five sequences of 30 minutes or longer and all were caused by exercise.
+These results suggest that although natural variations in glucose metabolism do lead to biological invariant violations, none of them were long enough to trigger our most severe action of transitioning to manual insulin dosing after two hours.
+We manually reviewed one weeks worth of Bob’s data and discovered that all five violations within this period were caused by exercise.
 10.4. Are predictive alerts useful for security?
-To measure predictive alerts as a security mechanism (Section 6), we recruited six students who live with T1D and had them use the GlucOS predictive alerts to complement their default CGM alerts.
-Our primary metric is time spent below 70 mg/dl, where a decrease in the amount of time spent in hypoglycemia suggests that predictive alerts are effective to empower people to mitigate security attacks.
+To measure predictive alerts as a security mechanism to avert hypoglycemia (Section 6), we recruited six students who live with T1D and had them use the GlucOS predictive alerts to complement their default CGM alerts.
+Our primary metric is time spent below range (i.e., below 70 mg/dl), where a decrease in the amount of time spent in hypoglycemia suggests that predictive alerts are effective to empower people to mitigate the most immediate and dangerous security risks.
 For the two month study, our participants decreased their average time in hypoglycemia from 2.7% down to 1.6%, and all individuals experienced a decrease in the amount of time spent in hypoglycemia.
 For context, participants in a state-of-the-art automated insulin delivery clinical trial experienced a decrease of time in hypoglycemia of 0.9% (brown2019six, ), showing that our 1.1% decrease is substantial.
-Overall, our results suggest that predictive alerts can be a useful last line of defense for trustworthy automated insulin delivery systems.
+We also report positive feedback from all participants during monthly interviews.
+Other than customization, participants acknowledged the importance of being able to reason about the alerts once we explained that we used linear regression for prediction.
+They reported that this helped forge trust in our system which in turn reduced their cognitive load with diabetes.
+All participants have continued using our system even after the study’s conclusion.
+Overall, our results show how predictive alerts are useful for automated insulin delivery.
 10.5. How does the BioKernel perform?
-We look at two aspects of system performance when evaluating our real-world use of GlucOS.
+We evaluate two aspects of GlucOS’s system performance in the real world.
 First, we measure how long our closed-loop algorithm takes to run and second, we count how often the closed-loop algorithm runs successfully end-to-end.
-We use Bob’s iPhone 14 to collect all of the data we show here and these numbers are averages across the entire eleven month period.
+We use Bob’s iPhone 14 to collect all of the data we show and these numbers are averages across the entire eleven month period.
 On average, the BioKernel spends 4.3s running the closed loop algorithm, where almost half of that time is dominated by CGM and pump communications.
@@
 11. Related work
-In addition to the work already discussed in this paper, there are several other related works.
 The complexity of automated insulin delivery systems and the fact that they make real-time healthcare decisions, makes their security paramount.
 Current research on security for these systems focuses primarily on evaluating algorithms using simulations (zhou2021data, ; zhou2023knowsafe, ).
-We argue that this approach falls short of addressing the full spectrum of challenges inherent in building trustworthy systems that operate within real-world constraints.
-Real-world systems must cope with runtime platform limitations, changes in human physiology, and the necessity of human interaction with the system.
-Plus, algorithms and simulations model attacks abstractly through their glucose response, missing out on details like which components of the system are likely to be compromised or how the system architecture can prevent damage from these potentially vulnerable components.
-Our experience shows that when faced with these practical challenges, algorithms designed in simulations often prove insufficient or inapplicable.
+Our experience shows that when faced with the practical challenges of real systems, algorithms designed in simulations often prove insufficient or inapplicable.
 These real-world complexities require a fundamentally different approach to designing security for automated insulin delivery systems.
-To address these challenges, GlucOS moves beyond simulation-based designs and tackles the challenges of real-world systems.
+GlucOS moves beyond simulation and tackles the challenges of real-world systems.
 Previous research has looked at the security of implanted medical devices in general (halperin2008security, ; burleson2012design, ; rushanan2014sok, ), in addition to looking at insulin pumps in particular (li2011hijacking, ; paul2011review, ), with more recent work looking at providing improved security (marin2016feasibility, ; ahmad2018securing, ).
 Also, recent work has looked at applying formal methods to insulin pumps for high assurance (panda2021secure, ).
-These works focus on devices and their communication channels.
-In contrast, with GlucOS we assume that these devices are correct and secure and focus our efforts on the software we use to run the automated insulin delivery system.
+They focus on devices and their communication channels.
+With GlucOS, we assume that these devices are secure and focus on automated insulin delivery software.
 12. Conclusion
@@
 First, our approach to algorithmic security used classic security techniques, like reference monitors and the Simplex architecture.
-However, it was our novel insight around all correct algorithms dosing the same amount of insulin that led us to these classic techniques.
-Plus, from our exploration, the most interesting part was not the ML, it was around the principles of reactive safe models and in particular our principle that these safety-critical algorithms should be reactive.
-Second, our pump driver security went beyond traditional specification and checking.
-We needed to actively program the pump to handle dropped cancel commands, which was an unexpected but crucial insight.
-Third, the biological invariant emerged as a key check in our system, addressing a blind spot in typical simulations that ignore changing human dynamics.
+However, our novel insight around all algorithms dosing the same amount of insulin led to the application of these classic techniques.
+Plus, from our exploration, the most interesting part was not the use ML, but the principle of repurposing reactive models for security.
+Second, the biological invariant emerged as a key check in our system, addressing a blind spot in typical simulations that ignore changing human dynamics.
 This forced us to develop novel formal methods to handle scenarios where our glucose metabolism models could be drastically wrong.
-Fourth, our approach to human interaction in the system was nuanced.
-We found that simply alerting humans or suspending insulin delivery were not always appropriate.
-For security alerts, personalization and customization were important.
-Stopping insulin delivery could be dangerous if the user has high glucose and insufficient insulin on board, which required more context-aware responses.
+Third, our approach to human interaction in the system was nuanced.
+We found that simply alerting humans was not sufficient.
+For security alerts, personalization, customization, and simplicity was important.
 These non-obvious design decisions underscore the complexity of creating a truly trustworthy automated insulin delivery system.
 They reflect the value of combining theoretical security principles with practical, real-world testing and domain-specific knowledge.
-This interdisciplinary research drew on deep expertise from computer security, medicine, and ethics to create a practical system.
-Our exploration is an extreme example of personalized, distributed, and automated healthcare using trustworthy computer systems.
+This interdisciplinary research drew on deep expertise from operating systems, computer security, medicine, and ethics to create a practical system.
+Appendix A Ethical and safety considerations
+One individual, who we will call Bob, used GlucOS for 9 months.
+Six university students also used the predictive alerts from GlucOS for 2 months.
+We put in place a protocol where we would stop our study and ask people to stop using our system and revert back to their previous management regime.
+In our protocol, we have formal in-person meetings with Bob every week to check his data and make sure that he is not incurring risk as a result of using GlucOS, and weekly surveys for the six students, plus monthly meetings.
+We would have stopped our study if any of the following conditions occurred:
+  • Excessive hypoglycemia.
+We define excessive hypoglycemia as having spent 8% or more of time during a week with CGM readings below 70 mg/dl or more than 1% below 54 mg/dl.
+These amounts of time spent in hypoglycemia would have represented a significant increase in time spent in hypoglycemia for participants.
+  • Insufficient CGM data. If they had less than 70% of the time during a week with CGM readings, then the lack of CGM data would represent a decrease for the participants.
+  • Insufficient closed-loop runs.
+We designed GlucOS to run closed-loop algorithms every five minutes, or 288 times per day.
+If Bob had a day where our closed-loop algorithm ran successfully less than 200 times, it would represent a fundamental flaw in the system.
+  • Serious complications due to diabetes.
+Serious complications include any hospitalization for diabetes related issues, severe hypoglycemia where a participant was unable to recover from hypoglycemia themselves and had to get help from someone else, diabetic ketoacidosis, or seizure.
+  • New medical diagnosis requiring significant attention. If anyone had received a new medical diagnosis during the study, the study would have had the potential to be a distraction for participants.
+  • New mental health diagnosis. If anyone were struggling with mental health during the study, the study would have the potential to be a distraction.
+In addition to our safety protocol, we followed best practices for ethical research to ensure that we conducted our study ethically.
+  • IRB.
+Our study went through our university’s IRB process, where the study including six university students who used GlucOS’s alerting software was approved, and our case study on Bob was determined to be “self evaluation” and exempt from IRB.
+  • Informed consent. All participants went through a written informed consent process.
+  • Data management.
+GlucOS does not store any personally identifiable information and uses access control on our Google Cloud Platform infrastructure to ensure that only the authors of the paper can access this anonymized data.
+  • Risk/benefit analysis.
+All participants are already living with T1D and use CGMs and inject insulin, so they already take on the risks inherent in managing T1D.
+The benefits include improved control over their T1D.
+There are risks in using experimental software, but all participants continued to rely on their CGM alerts as a last line of defense.
+  • Participant selection. We selected participants who were university students living with T1D.
+  • Adverse event reporting. Had any participants experienced adverse events during the study, we would have reported it to our IRB for guidance and addressed it according to our safety protocol.
+  • Communication. All surveys were done through our university’s medical school infrastructure for private communication with individuals.
+Since Bob used the full GlucOS software, including for insulin delivery, we included several other safety mechanisms:
+  • Bob increased the frequency of his visits to his Endocrinologist while using the GlucOS software. He went from yearly appointments to quarterly appointments.
+  • During the study, twice Bob got lab work based on blood tests to provide an independent measure of his overall metabolic health.
+This lab work helped to ensure that he was in good health and that there were no hidden negative impacts from his use of GlucOS.
+Overall, GlucOS had a large positive impact on the individuals who used it.
+Bob’s lab-based A1C test (a blood test that measures average glucose levels over time), which includes his time using GlucOS, was 5.8%.
+That is nearly non-diabetic levels of control – healthy individuals will have an A1C of 5.6% or lower.
+All of the students decreased the amount of time spent in hypoglycemia.
+These results are profound for the seven people using GlucOS, and it had a positive impact on their health and quality of life.
+Finally, Bob was a willing and enthusiastic user of the GlucOS system.
+To be clear, Bob is not a graduate student who was forced to use GlucOS so that his advisor could publish a paper.
+Rather, Bob struggled with the cognitive load of T1D management and wanted an automated insulin delivery system to offload part of the burden.
+However, he was unwilling to use any of the other automated insulin delivery systems out there due to concerns over their security.
+Bob found that GlucOS, with its focus on security, correctness, and simplicity, met his needs.
+Appendix B Derving the implementational invariant
+In this section, we show how we derive the implementational invariant from the biological invariant.
+We introduce the following equations to capture the error bounds on CGMs and pumps under standard conditions:
+(3) $0.9\cdot G_{t}^{A}\leq G_{t}^{M}\leq 1.1\cdot G_{t}^{A}$
+(4) $0.99\cdot I_{t}^{A}\leq I_{t}^{M}\leq 1.01\cdot I_{t}^{A}$
+Here, $I_{t}^{M}$ and $G_{t}^{M}$ denote the measured insulin dose and glucose concentration respectively at time $t$, while $I_{t}^{A}$ and $G_{t}^{A}$ represent their true values.
+We now recall the biological invariant (Equation $(1)$ in Section 5):
+(5) $\Delta G^{C}-\Delta G^{A}$ $\leq T,\quad\text{or}\quad\Delta G^{C}-T\leq\Delta G^{A}$
+$\Delta G^{A}=G_{t+\Delta t}^{A}-G_{t}^{A}$ is the actual drop in glucose over the same period.
+We note that we do not know the value of $\Delta G^{A}$ since we cannot capture the user’s true glucose concentration.
+Here, $\Delta G^{C}$ is the calculated drop in the user’s glucose over period $\Delta t$ and $T$ is the clinical threshold of 30 mg/dl.
+We adapt mathematical models from prior literature [7, 69] to compute $\Delta G^{C}$:
+(6) $\Delta G^{C}=S\cdot(B\cdot\Delta t-\Delta IOB^{A}+\Sigma I_{i}^{A})$
+where $S$ is the BioKernel’s estimate of the user’s insulin sensitivity and $B$ is the user’s basal rate (a physiological parameter representing the amount of insulin required per hour to absorb the background glucose produced by the body).
+Basal rate does not change as frequently as the insulin sensitivity and often remains stable for days [46].
+$\Delta IOB^{A}$ captures the change in the amount of insulin in the user’s body over the period $\Delta t$ and $\Sigma I_{i}^{A}$ is the sum of insulin doses delivered over the same period.
+$\Delta IOB^{A}$ at time $t$ is calculated as the dot product of insulin doses across the 6-hour window ending at $t$ with coefficients from standard exponential curves.
+In other words, both $\Delta IOB^{A}$ and $\Sigma I_{i}^{A}$ depend on insulin doses which cannot be captured precisely.
+To account for these unknowns, we combine $(3)$, $(4)$, $(5)$ and $(6)$ to get:
+(7) $S\leq\frac{0.9\cdot G_{t+\Delta t}^{M}-1.1\cdot G_{t}^{M}+T}{B\cdot\Delta t-0.99\cdot\Delta IOB^{M}+0.01\cdot\Sigma I_{i}^{\ M}}$
+Here, $\Delta IOB^{M}$ and $\Sigma I_{i}^{M}$ are the measured equivalents of $\Delta IOB^{A}$ and $\Sigma I_{i}^{M}$, recorded from the pump.
+We note that Equation $(7)$ is the implementational invariant(Equation $(2)$ in Section 5), thereby showing how we derive it from the biological invariant.
+Figure 8.
+Variation in CGM readings for three virtual humans when running a predictive ML algorithm that intentionally doses one-tenth the amount of the required insulin.
+The CGM readings staying mostly above the hyperglycemic bound would have had drastic health consequences in all three individuals.
+Figure 9.
+Variation in CGM readings for three virtual humans when clamping an algorithm that intentionally tries to dose one-tenth the amount of the required insulin.
+On average, CGM readings stay in range over 70% of the time.
+Appendix C Evaluating algorithmic security when intentionally underdosing insulin
+We repeat the experiment described in Section 9.1, but intentionally dose one-tenth the required amount of insulin.
+This experiment represents an algorithm that intentionally underdoses insulin as opposed to overdosing insulin.
+Figure 8 shows that the same virtual humans described in Section 9.1 spend most of their time above the hyperglycemic bound which would have led to severe health consequences in all three individuals.
+Figure 9 shows the same virtual humans in the same simulation scenarios with the same malicious algorithm dosing one-tenth the required amount of insulin, but with GlucOS’s algorithmic security mechanism enforced.
+We now see all virtual humans spending over 70% of their time in the recommended healthy range.
+Appendix D Driver security: Withstanding malicious pump drivers
+Pump drivers provide abstractions for insulin delivery applications to communicate with insulin pumps.
+Much like traditional device drivers, they have complex implementations since they help accomplish a wide range of tasks including pump setup/teardown, managing pump UX, handling insulin delivery/suspension, managing pump errors, accurately tracking insulin doses, and so on.
+Given this complexity, there is a strong possibility for bugs or vulnerabilities to be present in existing pump drivers.
+These vulnerabilities can harm or even lead to the death of individuals since pump drivers control insulin pumps.
+For example, the user could be under risk of insulin overdose if the pump driver inadvertently or intentionally issues the same insulin dosing command twice.
+Our goal is to protect users from inadvertent bugs in pump drivers as well as malicious pump drivers.
+We introduce a shim between the pump driver and the pump, which validates all communication between them.
+We design our mechanism to interpose on existing pump driver-pump communication instead of directly programming the pump.
+This designs ensures that our trusted computing base remains small and simple by not including complex pump driver implementations.
+This design also provides flexibility with altering or introducing new functionality to pump drivers.
+Interestingly, to protect the individual, we had to overrule our principle of not programming the pump in cases pertaining to commands for canceling ongoing insulin delivery.
+We could not directly adopt the design of other device driver security mechanisms [72, 65, 49] for our driver security mechanism since these cases are specific to insulin pumps.
+We focus on pump drivers because they are complex (14.5k LoC), multi-threaded, event-driven, and use two-way communication between the pump and the driver.
+In contrast, CGM drivers are simple (1.9k LoC) and only read sensor values from the device.
+In our future work, we will formally verify CGM drivers using the same techniques we use for the insulin delivery path (Section 7).
+D.1. Driver security mechanism design
+Figure 10. We interpose on communications between the pump driver and the pump to handle untrusted pump drivers.
+Figure 10 provides an overview of our driver security mechanism.
+The BioKernel registers high-level commands, like “deliver 2U of insulin” with our driver security mechanism, and sends them to the pump driver.
+The pump driver converts these commands into I/O messages to program the pump.
+The driver security mechanism then checks to make sure that these I/O messages are consistent with the previously registered high-level pump commands before forwarding them to the pump hardware.
+We only validate commands for insulin delivery since they pose the biggest threats to the user.
+On detecting added, modified or dropped commands from the pump driver, the driver security mechanism rejects further commands and transitions the system to manual insulin delivery where the user takes control.
+We list insulin message types pertaining to insulin delivery, pump setup and UX in Figure 12.
+Dropped commands pertaining to canceling ongoing insulin delivery require separate consideration since excess insulin is dangerous and cannot be withdrawn once it has been delivered.
+Users will cancel insulin delivery if they accidentally issued an insulin delivery command or incorrectly set the insulin dosage.
+In cases where a malicious driver drops a cancel command, the driver security mechanism directly issues I/O messages to the pump to cancel insulin delivery before transitioning the pump to a manual mode.
+We make the tradeoff of increasing the size of our trusted computing base to handle these cases in the interest of protecting the user.
+Fortunately, there are only two commands to cancel insulin delivery in our current implementation, so the increase to our trusted computing base is small.
+Figure 11.
+We attached an Omnipod DASH insulin pump to a stuffed animal and modified the OmniB LE pump driver to intentionally add, drop and modify insulin commands.
+Our driver security mec hanism detected all manipulations.
+D.2. Does GlucOS ensure pump operation while overcoming malicious pump drivers?
+We evaluate the driver security mechanism in allowing communications from a benign pump driver to go through while blocking manipulated communications from a malicious pump driver.
+Malicious pump drivers We modified the OmniBLE pump driver [4] to intentionally add, drop and modify insulin commands sent to an Omnipod DASH insulin pump [3] attached to a stuffed animal (Figure 11).
+We report that the driver security mechanism detected all manipulated commands, resulting in 0 false negatives.
+Benign pump drivers For one week, we logged messages passed between the BioKernel, the unmodified OmniBLE pump driver and the pump on an individual running GlucOS (Section 10).
+We report 100% consistency between all commands showing that GlucOS’s driver security mechanism does not incur any false positives.
+These results show that GlucOS can ensure pump operation while withstanding attacks from malicious drivers.
+Message Message Message
+name type description
+Insulin Request insulin delivery
+getStatus delivery status from the pump
+Insulin Insulin dose delivery
+statusResponse delivery information from the pump
+Insulin Insulin dose delivery
+errorResponse delivery error from the pump
+Insulin Set default rate of
+setBasalSchedule delivery insulin delivery
+Insulin Set temporary rate
+setTempBasal delivery of insulin delivery
+Insulin
+bolus delivery Deliver instant insulin dose
+Insulin
+cancelDelivery delivery Cancel insulin delivery
+setupPod Setup Setup a new pump
+assignAddress Setup Pair pump with phone
+deactivatePod Setup Teardown pump
+acknowledgeAlert UX Acknowledge alert from pump
+configureAlerts UX Configure alerts from pump
+beepConfig UX Configure pump beeps
+Figure 12. Types of messages exchanged between a pump driver and pump pertaining to delivery, setup and UX.
+Figure 13. GlucOS’s CGM state machine to handle missing readings from CGMs.
+Appendix E Runtime failure management
+Figure 14.
+GlucOS’s state machine to transition users to the safest manual dosing state when losing trust in insulin delivery (i.e., with a buggy or malicious pump driver) or when the biological invariant is not satisfied.
+In this section, we describe GlucOS’s state machine that combines its security components, i.e., algorithmic security (Section 4), driver security (Section D) and end-to-end security (Section 5) together to protect users in light of their corresponding failures.
+GlucOS’s state machine consists of three operating states: (1) ML with safety operating in closed loop, (2) reactive safe model operating in closed loop and (3) manual dosing in closed loop.
+ML provides the tightest level of control in managing glucose but poses the largest immediate threat from excess insulin since it preemptively injects insulin by anticipating rise in glucose.
+The reactive safe model provides slightly lesser control, but poses milder immediate threat since it only injects insulin to account for measured excess glucose.
+Manual dosing offers the least amount of control but poses no threat since it only injects the minimum insulin needed for sustenance.
+GlucOS predominantly operates in the state running ML (with safety) in closed loop.
+We transition to the safest manual dosing state when we do not trust insulin delivery (i.e., with a buggy or malicious pump driver).
+We also transition to the safest manual operation when the biological invariant is not satisfied to ensure that we do not risk the possibility of removing too much glucose from our insulin doses.
+If we end up in the manual dosing state, we enforce that only manual user intervention can transition back to other states.
+We visualize these state transitions in Figure 14.
+Inability to retrieve readings from a CGM also count as runtime failures. However, these failures are easier to handle.
+Since ML needs a stream of CGM readings as input, we transition from ML to the reactive safe model if we don’t receive an input from the CGM.
+As long as the number of missing CGM readings is low (say, within 30 minutes), we can interpolate to account for the missing readings to operate the reactive safe model.
+Once we have too many missing CGM readings that make us lose confidence in interpolation, we transition to manual dosing.
+From manual dosing, we transition back to the reactive safe model on receiving a CGM reading.
+Lastly, from the reactive safe model, we transition back to ML once we have enough CGM readings that are required by the ML model.
+Figure 13 summarizes these transitions.
+Bug Type Description
+Division by Zero Error Potential division by zero error when calculating insulin dosage if the user entered the same start and end time for a basal rate, resulting in a duration of zero.
+Negative Value Handling The insulin dosing calculation for microbolusing and basal rate erroneously allowed negative values.
+Out-of-Range Value Handling The microbolusing function failed to handle cases where one of its calculation components was out of the expected range or resulted in a negative value.
+Violation of Intended Postcondition The function responsible for setting the insulin delivery duration violated its intended postcondition by allowing the end date to be earlier than the start date.
+Violation of Intended Postcondition The function responsible for calculating the total insulin dose violated its intended postcondition by returning an incorrect upper bound value under certain conditions.
+Violation of Intended Postcondition The function responsible for updating the user’s insulin dosage violated its intended postcondition by failing to validate the input parameters correctly.
+Violation of Intended Postcondition The function responsible for calculation of ‘unitsDelivered‘ needed additional postconditions accurately reflect the amount of insulin delivered over the time gap.
+Lack of Error Handling Additional error handling was added to insulin delivered over a time segment function to gracefully catch undesirable state.
+Lack of Error Handling Additional error handling was added to microbolusing function to gracefully catch undesirable states.
+Figure 15. Report of bugs fixed by formally verifying GlucOS.
+Appendix F Formal verification: invariants and bugs
+In this section, we list the invariants used to formally verify critical functions of GlucOS. Figure 15 lists the bugs we found and fixed during formal verification.
+F.1. Invariants
+The MicroBolusing function ensures that micro boluses are always administered at safe time intervals while making sure that they are actually needed (the glucose level of the person is in fact high).
+  • Ensure no micro bolus in the last 4.2 minutes.
+  • Ensure glucose is at least 20 mg/dl above the target glucose level.
+  • Ensure predicted glucose is greater than the current glucose level minus 2.
+  • Ensure micro bolus amount is within the allowed range.
+  • Ensure bolus amount is rounded correctly if pumpManager is available.
+We assume that the BioKernel is the only software issuing bolus commands.
+The GuardRails function is responsible for ensuring a safe basal rate value, mindful of the user’s glucose levels. It enforces several invariants:
+  • Ensure the result of rounding newBasalRateRaw to the supported basal rate is within the allowable range.
+  • Ensure the new basal rate does not exceed the maximum basal rate in the settings.
+  • Ensure the new basal rate is not negative.
+  • Ensure that if either the current glucose level or the predicted glucose level falls below or equals the shut-off glucose threshold, the new basal rate is set to 0.0.
+The function operates under the assumption that all paths must go through the guard rails, with a specific evaluation comparing with the pump invariant.
+Additionally, it assumes the integrity of settings, trusting the underlying OS and file system and storage stack.
+The insulinDeliveredForSegment function ensures that intersectionStart is not greater than intersectionEnd when computing the intersection between self.startDate and self.endDate with startDate and endDate.
+  • Verifies that intersectionDuration is non-negative and does not exceed doseDuration, ensuring it accurately reflects the overlapping time duration.
+  • Checks that the units of insulin (units) used for calculation are either deliveredUnits or programmedUnits, ensuring consistency in how the insulin amount is derived.
+  • Verifies that the computation (units * intersectionDuration / doseDuration) accurately represents the amount of insulin delivered during the intersection period.
+The function assumes that the input parameters (self.startDate, self.endDate, startDate, endDate, units, and doseDuration) are valid and within the expected ranges.
+It is also assumed that the underlying data structures and calculations are accurate and consistent throughout the system.
+The CreateBasalDose function requires a time gap greater than 1 second to proceed and return a valid DoseEntry.
+  • Ensures that the calculation of basalRatePerSecond from basalRate accurately represents the rate of insulin delivery per second.
+  • Ensures that the calculation of unitsDelivered accurately reflects the amount of insulin delivered over the time gap.
+  • The DoseEntry constructed should have consistent attributes and adhere to the specified type, units (unitsPerHour), and mutability (false).
+The function assumes that the input parameters (basalRate, startDate, and endDate) are valid and within the expected ranges.
+It also assumes that the underlying time calculations and conversions between units are accurate and consistent.
+The inferBasalDoses function ensures that basalDoses contains only DoseEntry instances where type is .tempBasal, .resume, or .suspend, sorted in ascending order by startDate.
+  • InsulinType should be determined correctly based on the last .tempBasal or .bolus type dose in doses, defaulting to .humalog if none are found.
+  • Each inferred basal dose added to inferredBasalDoses must be created using the createBasalDose function with valid parameters and added in the correct chronological order.
+  • If the last dose in basalDoses is not of type .suspend, a valid basal dose must be inferred from its endDate to at.
+The function assumes that the input parameters (doses, basalDoses, at, and pumpRecordsBasalProfileStartEvents) are valid and within the expected ranges.
+It also assumes the correctness of the underlying data structures and the createBasalDose function.
+The insulinOnBoard function ensures that doses contains unique and valid DoseEntry instances filtered up to at.
+  • The total iob should accurately represent the cumulative insulin on board from all valid dose entries in doses.
+  • basalDoses should contain inferred basal doses that accurately reflect insulin on board contributions when pumpRecordsBasalProfileStartEvents is false.
+  • The final return value of insulinOnBoard should be the sum of iob and basalIob, representing the total insulin on board at at.
+The function assumes that the input parameters (doses, at, and pumpRecordsBasalProfileStartEvents) are valid and within the expected ranges.
+It also assumes the correctness of the underlying data structures and the inferBasalDoses function.
+The ActiveBolus function ensures that data is read consistently from disk before proceeding with any operations.
+  • Verifies that DeduplicatedDoses(events: eventLog, at: at) correctly filters out duplicates and retains only relevant dose entries up to at.
+  • Ensures that doses.filter accurately selects bolus entries that were active at the specified at.
+  • Verifies that doses.last correctly returns the last active bolus entry, ensuring it is non-null if a bolus exists within the specified criteria.
+The function assumes that the input parameters (eventLog and at) are valid and within the expected ranges.
+It also assumes the correctness of the underlying data structures and the DeduplicatedDoses function, as well as the consistency and reliability of disk read operations.
+The TempBasal (Safety Clamps) function ensures that the events array includes only events from safetyStates that fall within the time range [start - duration, at).
+  • Verifies that the historicalMlInsulin value represents the total units of insulin delivered by the machine learning system over the specified time horizon.
+  • Ensures that the machine learning and safety temp basal rates are correctly converted to units of insulin based on the given duration.
+  • Ensures that the upper and lower bounds for insulin units are correctly adjusted based on historicalMlInsulin.
+  • Ensures that the difference between machine learning and safety temp basal units is correctly clamped within the calculated bounds.
+  • The clamped delta units should be converted back to a temp basal rate and added to the safety temp basal rate.
+  • The returned SafetyTempBasal object should correctly encapsulate the adjusted temp basal rate and historical machine learning insulin.
+The function assumes that the input parameters (safetyStates, start, duration, at, mlTempBasal, safetyTempBasal, and lastHistoricalMlInsulin) are valid and within the expected ranges.
+It also assumes the correctness of the underlying data structures and the conversion functions between units, as well as the proper configuration and appropriateness of the safety clamp parameters and thresholds for the user.
+Appendix G Scenarios when the user’s insulin sensitivity is lower than the BioKernel’s estimate
+We do not take any action when the actual change in the user’s glucose is higher than the calculated change.
+Higher actual glucose does not necessarily reflect a shift in the user’s physiological state (decreased insulin sensitivity) since it can also result from digestion when the user consumes food.
+Further, even if the user is less sensitive to insulin, their glucose absorption would be slower, which does not pose them with an immediate threat (hyperglycemia).